### v1

# HEST-1k: A Dataset for Spatial Transcriptomics and Histology Image Analysis

Abstract
Spatial transcriptomics (ST) enables interrogating the molecular composition of tissue with ever-increasing resolution, depth, and sensitivity. However, costs, rapidly evolving technology, and lack of standards have constrained computational methods in ST to narrow tasks and small cohorts. In addition, the underlying tissue morphology as reflected by H&E-stained whole slide images (WSIs) encodes rich information often overlooked in ST studies. Here, we introduce HEST-1k, a collection of 1,108 spatial transcriptomic profiles, each linked to a WSI and metadata. HEST-1k was assembled using HEST-Library from 131 public and internal cohorts encompassing 25 organs, two species (Homo Sapiens and Mus Musculus), and 320 cancer samples from 25 cancer types. HEST-1k processing enabled the identification of 1.5 million expression–morphology pairs and 60 million nuclei. HEST-1k is tested on three use cases: (1) benchmarking foundation models for histopathology (HEST-Benchmark), (2) biomarker identification, and (3) multimodal representation learning. HEST-1k, HEST-Library, and HEST-Benchmark can be freely accessed via https://github.com/mahmoodlab/hest.
**footnotetext: Equal contribution
1 Introduction
Advances in molecular profiling enable spatially-resolved gene expression analysis with increasingly large gene panels, enhanced spatial resolution, and greater sensitivity Asp et al. (2020); Rao et al. (2021). From the early days of bulk RNA sequencing constrained by its coarse resolution and limited gene panels, spatially-resolved technologies have progressed to achieve whole-transcriptome sequencing at sub-cellular resolution Li and Wang (2021). In cancer research, spatial transcriptomics (ST) holds particular promise for characterizing the tumor microenvironment, a key element in understanding disease progression and treatment response Ståhl et al. (2016b); Marx (2021); de Visser and Joyce (2023); Xiao and Yu (2021). With the large amount of transcriptomics data generated by a single ST sample (e.g., $>$10 million transcripts are detected in a typical 10x Genomics Xenium assay), computational methods are often used to identify promising targets and biomarkers, such as based on gene expression clustering or gene-set enrichment analysis Rao et al. (2021).
However, high costs and rapidly evolving technology have constrained computational methods to narrow tasks and data cohorts of only few patients Meylan et al. (2022); Janesick et al. (2023); Valdeolivas et al. (2023). Consequently, we observe a lack of standardized resources and unified formats for storing ST, which limits the development of deep learning models on a large scale Moses and Pachter (2022). In addition, the underlying tissue morphology, traditionally visualized in hematoxylin and eosin (H&E)-stained tissue sections (whole-slide images, WSIs), is often overlooked in ST studies, despite encoding valuable information. In particular, pairs of ST and WSI enable analyzing expression changes in their morphological context, which may facilitate the discovery of morphological biomarkers (e.g., changes in nuclear shape) that correspond to gene regulation patterns. In translational medicine, ST could be employed for identifying biomarkers, which are then morphologically validated and characterized by widely deployed clinical modalities such as H&E. Alternatively, pairs of ST and WSI can enable multimodal tissue representation learning for joint modeling of the morphomolecular signature of tissue at a scale and resolution beyond bulk RNA sequencing Chen et al. (2022). Finally, the development of “foundation models” for encoding histopathology image patches Wang et al. (2021); Filiot et al. (2023); Chen et al. (2024); Huang et al. (2023) has increased the need for new, diverse, and challenging benchmarks beyond diagnostic tasks. Using ST, new tasks can be defined for predicting gene expression from histology.
Here, we introduce HEST-1k, a collection of paired ST and H&E-stained WSIs curated from public and internal cohorts (Figure 1.a). HEST-1k comprises 1,108 samples from 131 cohorts encompassing 25 organs, two species (Homo Sapiens and Mus Musculus), and 320 cancer samples from 25 subtypes. Processing all samples in HEST-1k resulted in 1.5 million expression–morphology pairs and 60 million detected nuclei. With new cohorts frequently made public, we also introduce the HEST-Library, a Python package for querying HEST-1k and assembling new legacy data (Figure 1.b). We highlight the potential of HEST-1k through three use cases: (i) benchmarking foundation models for histology using the HEST-Benchmark, a set of ten tasks (nine human cancer types and ten organs) for gene expression prediction from histology and evaluated on ten state-of-the-art models (Figure 1.c), (ii) a proof-of-concept demonstrating the use of HEST-1k for biomarker characterization (Figure 1.d), and (iii) a proof-of-concept for expression-guided fine-tuning of foundation models for histology (Figure 1.e).
Figure 1: Overview of HEST. a. Description of HEST-1k, a dataset of $n$=1,108 paired Spatial Transcriptomics, H&E-stained whole-slide images and metadata. b. Overview of HEST-Library functionalities. c., d., e. Potential applications of HEST-1k include benchmarking foundation models for histology (c.), biomarker discovery (d.) and multimodal representation learning (e.).
2 Related work
Libraries for ST analysis. Libraries to process, visualize, and analyze ST have been built around two core pipelines: Scanpy Wolf et al. (2018) (and the Anndata format) in Python and Seurat Butler et al. (2018) in R. Scanpy has served as the foundation for several subsequent developments such as Squidpy Palla et al. (2022) for spatial data exploration at cellular-, gene-, and morphological-level, SpatialData Marconato et al. (2024) for multi-technology integration and deep learning interfacing, and STlearn Pham et al. (2023) for cell-cell interactions and spatiotemporal trajectory analyses. In R, Seurat Butler et al. (2018) has been consolidated with packages such as BayeSpace Zhao et al. (2021) for clustering and spot super-resolution, and Giotto suites Chen et al. (2023) for preprocessing, data integration, and visualization of multiple ST technologies. 10x genomics also includes software analytics through the Xenium and Visium Explorer pipelines for multimodal visualization, nuclear segmentation, and cell deconvolution. However, none of these pipelines were designed to handle the diversity of legacy data, where datasets can suffer from missing or incorrect data, such as alignment mismatches, incorrect pixel resolution, inconsistent image file formats, etc.
Molecular profile prediction from H&E. Molecular profiling from histology images has been explored both at (1) slide-level to predict bulk molecular status/changes from a WSI, and at (2) patch-level to predict local molecular status/changes from regions-of-interest. (1) Slide-level profiling has been proposed to predict the presence of gene mutations Saldanha et al. (2023); Wagner et al. (2023); Kather et al. (2020); Echle et al. (2021); Fu et al. (2020); Wang et al. (2019); Loeffler et al. (2022), microsatellite instability Wagner et al. (2023); Kather et al. (2019b), and gene expression changes Schmauch et al. (2020); El Nahhas et al. (2024); He et al. (2020); Alsaafin et al. (2023), among others. The motivation is two-fold: First, patient screening to substitute or emulate costly clinical molecular assays, and second, to identify morphological correlates of molecular alterations for discovering novel biomarkers. Such studies can be conducted on large patient cohorts as they mainly rely on data generated by the routine clinical workflow (e.g., using TCGA cohorts with $>$11,000 cases from 33 cancer types). (2) With ST, several works have explored predicting expression changes from regions-of-interest He et al. (2020); Xie et al. (2023); Pang et al. (2021); Zhao et al. (2024); Rahaman et al. (2023); Monjo et al. (2022); Mondol et al. (2023); Chung et al. (2024); Wang et al. (2024). Due to limited cohort sizes (typically one to ten patients), transfer learning has become the norm using pretrained networks based on ConvNets He et al. (2020) or Vision Transformers Xie et al. (2023); Pang et al. (2021). Transcriptomic measurements are inherently noisy, hence the development of context-based methods that integrate global and local information from surrounding ST spots Chung et al. (2024); Zhao et al. (2024); Wang et al. (2024). While recent technologies offer near-single-cell resolution (such as Visium HD and Xenium), legacy assays operate at a more coarse resolution, which can be upsampled using super-resolution techniques Zhao et al. (2021); Bergenstråhle et al. (2022); Zhang et al. (2024). The potential clinical and research implications of such methods are still being explored, with HEST-1k potentially catalyzing their large-scale development.
Foundation models in pathology. A fundamental task in computational pathology is to extract general-purpose embeddings of image patches (typically 256$\times$256-pixel regions) that can then be used for downstream tasks, such as diagnosis or prognosis prediction. To achieve this, self-supervised learning (SSL) has been extensively applied  Koohbanani et al. (2021); Filiot et al. (2023); Vorontsov et al. (2023); Wang et al. (2021); Kang et al. (2023); Ciga et al. (2022); Wang et al. (2022); Azizi et al. (2023); Chen et al. (2024); Lu et al. (2024), such as based on the DINOv2 framework Oquab et al. (2023). Models for general-purpose patch encoding are trained on increasingly large and more diverse patient cohorts with models growing in model size (e.g., Virchow uses a ViT-Huge trained on more than 1.5M WSIs Vorontsov et al. (2023)). As the number of such models rapidly increases, new, diverse, and challenging benchmarks are needed to replace or complement well-established tasks where performance has saturated. HEST-Benchmark aims to address this by offering a set of ten patch-level tasks for gene expression prediction from histology.
Patch-level benchmarks in histopathology. Early task and dataset contributions in computational pathology revolved around classifying small regions of interest. Over the years, a variety of benchmarks have been established: In prostate cancer, Gleason grading at pixel- and patch-level has been widely explored, with public resources such as PANDA Bulten et al. (2022), AGCC Huo et al. (2022), DiagSet Koziarski et al. (2024), and SICAPv2 Silva-Rodríguez et al. (2020). In colorectal cancer, datasets have been proposed for tissue classification, such as HunCRC Pataki et al. (2022), UniToPatho Barbano et al. (2021), MHIST Wei et al. (2021), and CRC-100k Kather et al. (2019a). In breast cancer, morphological subtyping has been studied (e.g., for atypical ductal hyperplasia detection), such as BACH Aresta et al. (2019), BRACS Brancati et al. (2021), and BreakHis Spanhol et al. (2015), and for lymph node metastasis detection with Patch CAMELYON (pCAM) Veeling et al. (2018), respectively. However, the performance on many of these datasets has saturated; for instance, Gleason scoring reaches similar or better performance than pathologists Bulten et al. (2022), which limits objective comparisons of new methods and hinders well-informed model selection for developing better features. Instead, HEST-Benchmark provides a collection of diverse and challenging tasks that enable assessing the predictive capabilities of foundation models for histology.
3 HEST-1k Dataset
We present HEST-1k, a dataset of paired ST, H&E-stained WSIs, and metadata (Figure 1.a). To this end, we extracted all publicly available cohorts that provide ST with H&E-stained whole-slide images. Specifically, we harvested data from 10x Genomics public datasets (TENX) [^1] https://www.10xgenomics.com/datasets, Mendeley (MEND) [^2] https://data.mendeley.com/, Spatial-Research (SPA) [^3] https://www.spatialresearch.org/, Zenodo (ZEN) [^4] https://zenodo.org/, the National Center for Biotechnology Information (NCBI) [^5] https://www.ncbi.nlm.nih.gov/gds/, GitHub [^6] https://github.com/, the Human Cell Atlas [^7] https://data.humancellatlas.org/, and internal data cohorts. A summary of all sources is provided in Appendix Table A1 with specifics in Appendix Table A2, A4, A6,A7,A8,A9, and A10.
3.1 Metadata
As Spatial Transcriptomics were not intended for large-scale computational research, they are provided in various formats (e.g., images can be in JPG or TIFF format, with or without cross-modal alignment files) and resolutions. We unified all data with comprehensive metadata with generic-, histology-, and expression-related descriptors for all samples. Generic: We provide the reference to the original publication, download link, year of publication, license, and sample species. Each sample is then categorized as either healthy, cancer, treated (which refers to a post-compound administration), genetically modified (mostly knock-out mouse samples), or other diseases (with an additional description). All cancer samples were unified using the OncoTree code, a taxonomy of cancer types provided by the Memorial Sloan Kettering Cancer Center [^8] https://oncotree.mskcc.org. Finally, we include the organ using the highest level of the OncoTree taxonomy as a reference. Expression: We include the number of genes and spots per sample, the spot resolution and spacing, the total number of reads, and the mean number of reads per spot. We additionally provide the transcriptomic technology (ST, Visum, Visium HD, or Xenium). Histology: We provide the image resolution (in $\mu$m/pixel) and magnification as 10$\times$ (1.15 to 0.8 $\mu$m/px), 20$\times$ (0.8 to 0.4 $\mu$m/px) and 40$\times$ (0.4 to 0.1 $\mu$m/px). All images with a resolution lower than 1.15 $\mu$m/px were discarded to ensure an acceptable image quality. In addition, we provide the image size at the highest resolution, and tissue preparation protocol (frozen or formalin-fixed paraffin-embedded, FFPE).
3.2 Histology
All tissue sections were normalized and transformed into a generic TIFF object, a pyramidal image that can easily be integrated into computational frameworks using Openslide or viewers such as QuPath Bankhead et al. (2017). In addition, we provide a contour object that delineates all the tissue regions identified in the image. We developed a robust tissue detection method where we fine-tuned a DeepLabV3 Chen et al. (2017) model with an ImageNet-pretrained ResNet50 backbone on a set of annotated segmentation regions. From the tissue segmentation, we extracted 224$\times$224-pixel patches at 20$\times$ magnification around each spot. This yielded 1.5 million valid patches for which a corresponding expression profile was derived. Such patching can readily be used for various downstream tasks, such as employed in the HEST-Benchmark or multimodal fine-tuning of foundation models for histology (Section 5 and 7).
3.3 Nuclear segmentation and classification
In addition to patching, we include nuclear segmentation that delineates each nucleus identified in all slides from HEST-1k. Specifically, we use CellViT Hörst et al. (2024), a state-of-the-art nuclear segmentation model that was trained on the PanNuke dataset Gamper et al. (2019, 2020). CellViT enables joint instance segmentation and classification of each nucleus into five classes: neoplastic epithelial, non-neoplastic epithelial, inflammatory, stromal, and necrotic. On average, we identified 55.5k nuclei per slide (standard deviation 91.3k nuclei), for a total of 59.6 million nuclei identified across all samples. Among those 17.6 million are classified as neoplastic, 21.5 million as stromal, 4.9 million as normal epithelial, 15.4 million as inflammatory, and 76 thousand as necrotic. The resulting nuclear segmentation and classification can easily be visualized using QuPath (using geojson) or loaded as Python/R objects (using json).
3.4 Gene expression
All expression data were unified in a Anndata object that can be loaded with scanpy. Anndata encodes the gene names (as var) and number of spots (as obs). Each entry represents the raw transcript counts of a gene in a given spot. No additional normalization was conducted, and we let users explore various normalization strategies based on needs, e.g., using total count normalization, log-based re-weighting, etc. In addition, we include metadata to specify the number of genes, the gene panel, and the tissue site.
To use the expression in tandem with the WSI, an alignment file describing the mapping between the image and the spots is needed. However, relying on publicly available alignment information brings three challenges: (1) most datasets report alignment with respect to a low-resolution version of the image, (2) they are not standardized, and (3) alignment quality can be low. To address these limitations, we re-aligned all samples under the same unified format between the WSI and the corresponding expression profile. To this end, we developed an automatic alignment pipeline (see Section 4) and embedded the alignment in the scanpy object.
4 HEST-Library
The HEST-Library is built around scanpy and Anndata. At its core, the HEST-Library enables (1) assembling and querying HEST-1k, and (2) running the HEST-Benchmark (see Section 5). We describe its core functionalities, in particular for unifying legacy data.
Conversion to generic TIFF. We integrate functions to convert a WSI from common formats found in public ST datasets (e.g., OME.TIF, JPG, BigTIFF, etc.) to a pyramidal generic TIFF format. Pyramidal formats offer seamless integration with OpenSlide (commonly used in computational pathology pipelines) and QuPath (open-access software for WSI visualization and annotation).
Automatic alignment. Spot alignment is crucial to ensure an accurate match between the ST spots and the WSI. While software such as LoupeBrowser enables manual alignment using the fiducials (in Visium), the process remains time-consuming when processing large batches of samples. Instead, we implemented an automatic fiducial detection algorithm based on Yolov8 Redmon et al. (2016) for processing Visium samples (Appendix Figure 4). Specifically, we manually annotated 119 fiducial regions that we further augmented using tissue and fiducials mixing. We then fine-tuned Yolov8 pretrained on COCO dataset. In samples without fiducials (e.g., early versions of Spatial Transcriptomics), we realigned using the provided spot position files.
Automatic detection of image resolution. From the alignment and the spot resolution, we can infer the exact pixel resolution. We start by deriving the distance in pixel between two neighboring spots and leverage the known inter-spot distance in $\mu$m to estimate the pixel width in $\mu$m/px. For Xenium samples, we directly read the H&E alignment file which provides an affine transformation from the DAPI-stained image (with known pixel size) to the H&E image. We then compared the self-reported image resolution and our re-estimations to manually inspect and correct discrepancies.
Conversion to Anndata. ST data is provided in multiple formats, such as CSVs, MEX, TXT, h5, etc. We provide functions to unify a large set of existing formats into a AnnData object that stores the raw transcript counts as a matrix of genes by the number of spots.
Tissue segmentation and patching. We provide a tissue segmentation pipeline optimized for Visium/Xenium images. The segmented tissue can automatically be tessellated into fixed-size image patches at a predefined resolution around each spot.
Automatic HEST-1k download. To facilitate downloading part or all of the HEST-1k dataset (in total 825 GB), we implemented an easy download option where the user can specify entries of the metadata, for instance, to query all human breast cancer cases.
5 HEST-Benchmark
From HEST-1k, we curated the HEST-Benchmark, a set of ten tasks for gene expression prediction from histology in human cancer samples. The goal is two-fold: (i) benchmarking foundation models for histology under a diverse and challenging benchmark, and (ii) understanding the predictive capabilities of state-of-the-art models in predicting expression from morphology. Compared to existing tasks (e.g., Camelyon16), HEST-Benchmark brings more morphological and data acquisition diversity, and more complex tasks as expression prediction is inherently challenging.
5.1 Task definition
We define ten tasks with data from nine human cancers and ten organs (nine primary and one metastatic dataset), which include invasive ductal carcinoma (breast cancer, IDC, Task 1), prostate adenocarcinoma (prostate cancer, PRAD, Task 2), pancreatic adenocarcinoma (pancreatic cancer, PRAD, Task 3), skin cutaneous melanoma (skin cancer, SKCM, Task 4), colonic adenocarcinoma (colon cancer, COAD, Task 5), rectal adenocarcinoma (rectum cancer, READ, Task 6), clear cell renal cell carcinoma (kidney cancer, ccRCC, Task 7), hepatocellular carcinoma (liver cancer, HCC, Task 8), lung adenocarcinoma (lung cancer, LUAD, Task 9), and axillary lymph nodes in IDC (metastatic, LYM-IDC, Task 10). Additional information is provided in Appendix Table A11.
For each task, we predict the expression of the top 50 genes with the highest normalized variance across all samples from 112$\times$112 $\mu$m H&E regions (equivalent to 224$\times$224-pixel patches at 20$\times$). To avoid train/test patient-level data leakage, we use patient-stratified splits, resulting in a $k$-fold cross-validation, where $k$ is the number of patients. In ccRCC, we use $k/2$-fold cross-validation due to the large number of patients.
5.2 Evaluating foundation models for histology on HEST-Benchmark
We use the HEST-Benchmark to evaluate ten patch encoders for histology. Namely, ResNet50 (IN) Lu et al. (2020) (ImageNet pretrained), KimiaNet Riasatian et al. (2021) (pretrained with supervised subtyping classification), Ciga Ciga et al. (2022) (SimCLR pretrained on public histology data), CTransPath Wang et al. (2022) (adapted MoCov3 pretrained on TCGA and PAIP), Remedis Azizi et al. (2023) (iBOT pretrained on TCGA), Phikon Filiot et al. (2023) (iBOT pretrained on TCGA), PLIP Huang et al. (2023) (visual-language model based on CLIP pretrained on histology data from Twitter), UNI Chen et al. (2024) (DINOv2 ViT-Large pretrained on public and private histology data), CONCH Lu et al. (2024) (visual-language model using CoCa pretrained on captions from publications and educational resources), and GigaPath Xu et al. (2024) (DINOv2 ViT-Giant pretrained on proprietary data). Additional information is provided in Table A12 and Section C.3.
Table 1: HEST-Benchmark results. Random Forest regression evaluation (70 trees) on publicly available state-of-the-art patch encoders. Model performance measured with Pearson correlation. We report the mean $\pm$ standard deviation computed over all folds (or patients). Best is bold, second best is underlined. Models sorted by publication date.
ResNet50 KimiaNet Ciga CTransPath Remedis Phikon PLIP UNI CONCH GigaPath
IDC 0.440 0.420 0.406 0.454 0.491 0.430 0.436 $\underline{0.502}$ $\bm{0.504}$ 0.492
±$\pm 0.03$$\pm 0.04$$\pm 0.02$$\pm 0.04$$\pm 0.06$$\pm 0.06$$\pm 0.04$$\pm 0.06$$\pm 0.05$$\pm 0.05$
PRAD 0.318 0.328 0.332 0.346 0.335 $\bm{0.377}$ 0.362 0.357 $\underline{0.373}$ 0.372
±$\pm 0.02$$\pm 0.01$$\pm 0.00$$\pm 0.01$$\pm 0.00$$\pm 0.00$$\pm 0.02$$\pm 0.00$$\pm 0.03$$\pm 0.01$
PAAD 0.389 0.410 0.397 0.406 $\bm{0.451}$ 0.372 0.392 0.424 $\underline{0.431}$ 0.425
±$\pm 0.07$$\pm 0.08$$\pm 0.07$$\pm 0.08$$\pm 0.06$$\pm 0.06$$\pm 0.09$$\pm 0.06$$\pm 0.05$$\pm 0.07$
SKCM 0.446 0.452 0.484 0.535 0.577 0.516 0.461 $\bm{0.613}$ $\underline{0.582}$ 0.541
±$\pm 0.06$$\pm 0.03$$\pm 0.01$$\pm 0.06$$\pm 0.00$$\pm 0.04$$\pm 0.06$$\pm 0.02$$\pm 0.04$$\pm 0.03$
COAD 0.107 0.080 0.102 0.123 0.125 0.137 0.112 $\bm{0.147}$ 0.124 $\underline{0.139}$
±$\pm 0.06$$\pm 0.03$$\pm 0.04$$\pm 0.05$$\pm 0.06$$\pm 0.05$$\pm 0.04$$\pm 0.05$$\pm 0.03$$\pm 0.03$
READ 0.051 0.038 0.046 0.083 0.099 0.138 0.063 $\bm{0.162}$ 0.132 $\underline{0.156}$
±$\pm 0.08$$\pm 0.05$$\pm 0.09$$\pm 0.09$$\pm 0.11$$\pm 0.10$$\pm 0.10$$\pm 0.08$$\pm 0.08$$\pm 0.08$
CCRCC 0.136 0.136 0.127 0.171 $\bm{0.200}$ 0.178 0.124 $\underline{0.186}$ 0.149 0.182
±$\pm 0.04$$\pm 0.05$$\pm 0.04$$\pm 0.04$$\pm 0.04$$\pm 0.05$$\pm 0.04$$\pm 0.05$$\pm 0.06$$\pm 0.05$
HCC 0.034 0.028 0.045 $\bm{0.060}$ $\underline{0.059}$ 0.041 0.038 0.051 0.040 0.055
±$\pm 0.01$$\pm 0.01$$\pm 0.00$$\pm 0.00$$\pm 0.00$$\pm 0.02$$\pm 0.00$$\pm 0.00$$\pm 0.01$$\pm 0.01$
LUNG 0.497 0.507 0.515 0.531 $\bm{0.573}$ 0.541 0.533 0.511 $\underline{0.569}$ 0.547
±$\pm 0.01$$\pm 0.01$$\pm 0.02$$\pm 0.02$$\pm 0.02$$\pm 0.03$$\pm 0.02$$\pm 0.03$$\pm 0.02$$\pm 0.03$
LYMPH_IDC 0.205 0.206 0.218 0.238 0.243 0.243 0.229 0.234 $\bm{0.249}$ $\underline{0.248}$
±$\pm 0.05$$\pm 0.06$$\pm 0.07$$\pm 0.06$$\pm 0.06$$\pm 0.06$$\pm 0.06$$\pm 0.05$$\pm 0.06$$\pm 0.05$
Average 0.262 0.261 0.267 0.295 0.315 0.297 0.275 $\bm{0.319}$ 0.315 $\underline{0.316}$
We learn a regression model to map model-specific patch embeddings (512 to 2048 dimensions) to the log1p-normalized expression of the top-50 highly variable genes. All tasks are evaluated using the Pearson correlation between the predicted and measured gene expression. We report mean and standard deviation across all folds (or patients). All experiments were run on a single NVIDIA 3090 GPU. Table 1 summarizes HEST-Benchmark when employing a Random Forest regression model (70 trees), and Table A13 when employing a Ridge regression model.
When using Random Forest, UNI Chen et al. (2024), GigaPath Xu et al. (2024), Remedis Azizi et al. (2023), and CONCH Lu et al. (2024) provide similar performance (0.315 to 0.319 average Pearson correlation across all tasks), outperforming all other baselines. When using a Ridge regression, CONCH Lu et al. (2024) significantly outperforms all models with an average of +5.0% absolute performance gain over the second-best model, Phikon, and 5.3% better than the third-best model, CTransPath with an average Pearson correlation of 0.316.
Supervised vs. contrastive vs. student-teacher pretraining: When employing both Random Forest or Ridge regression, supervised models (ResNet50-IN and KimiaNet) exhibit lower performance than other models. This can be attributed to (1) a lack of sufficiently large and diverse pretraining datasets and models or (2) fundamental limitations of supervised models in transferring to challenging tasks. Using Random Forest, the best student-teacher models (UNI, CONCH, GigaPath) provide similar performance to the best contrastive approach (Remedis).
Vision Transformer Base vs. Large vs. Giant: CONCH (ViT-Base, 86M parameters), UNI (ViT-Large, 307M) and GigaPath (ViT-Giant, 1.13B) provide similar performance despite GigaPath using 13$\times$ more parameters than CONCH. UNI and GigaPath have been trained in a unimodal fashion by building “views” using traditional augmentation techniques, such as color jittering and cropping. In contrast, CONCH was trained in two steps: first using unimodal iBOT pretraining, and then by fine-tuning the resulting model on image-caption pairs. Such fine-tuning step could explain its parameter-efficiency and high performance in HEST-Benchmark.
Random Forest vs. Ridge regression: The best-performing model using Random Forest (UNI, 0.319) is similar to the best-performing model using Ridge regression (CONCH, 0.316), showing the predictive capabilities of both regression models. However, Ridge regression leads to poor performance using larger models (GigaPath, Remedis, and UNI), which could be explained by the curse of dimensionality (CONCH embedding dimension is 3$\times$ smaller than GigaPath). We also emphasize that the performance could be optimized using more complex models (such as MLP) or dedicated hyper-parameter search.
Overall, HEST-Benchmark brings new insights into the performance of patch encoders for histology. We observe that (1) Random Forest regression leads to multiple models performing similarly (CONCH, UNI, GigaPath) which may suggest a form of plateauing of patch encoders, (2) CONCH is the only model that shows high performance using both Ridge and Random Forest regression models, which could be attributed to its compact size (ViT-B) and visual-language fine-tuning step, and (3) performance remains low for some tasks (e.g., READ and HCC) which implies that for certain cancer types, the morphology might not be as reflective of gene expression, as well as leaving room for follow-up works.
6 HEST for discovery
Figure 2: HEST for biomarker discovery: Analysis of an invasive ductal carcinoma Xenium sample. a. IDC Xenium samples with neoplastic nuclei overlaid in red ($n_{c}=168,033$ detected nuclei). Gray scale bar represents 2 mm. b. Heatmap of Xenium expression of gene GATA3. Blue and red values indicate above and below the mean (in white), respectively. c. Heatmap of neoplastic nuclear area. d. Four randomly selected regions with CellViT segmentation of the neoplastic nuclei. Black scale bar represents 30 $\mu$m. e., f. Correlation between area and GATA3, and minor axis length and MYBPC1.
HEST-1k also enables the analysis of interactions and correlations between tissue morphology (as seen in H&E) and local gene expression (as provided in ST). Here, we showcase the capabilities of HEST-1k (1) by studying morphological correlates of expression changes in invasive breast cancer and (2) by visualizing tumor heterogeneity both on the morphological and molecular sides. Specifically, we focus on invasive ductal carcinoma (IDC) samples imaged with Xenium. Using CellViT nuclear segmentation and classification, we identified neoplastic nuclei (exemplified in two samples: Figure 2.a with n=168,033 nuclei and Appendix Figure 5.a with n=342,018 nuclei). We then overlay the WSI with the expression of specific genes, such as GATA3, a known prognostic gene in breast cancer Mehra et al. (2005)(Figure 2.b). This qualitatively shows that high GATA3 expression is associated with cancerous regions and reveals heterogeneity within invasive regions (e.g., the right-most region shows higher expression of GATA3 than the rest of the tumor, Figure 2.b). Using the nuclear segmentation, we can compute human-interpretable features related to nuclear size (area, perimeter, major axis length, minor axis length, and equivalent diameter), topology and shape (roundness, ellipticity, eccentricity, extent, and roughness), and cell distribution (cell density and crowdedness). A heatmap of the nuclear area of neoplastic cells also indicates morphological heterogeneity among neoplastic regions (Figure 2.c,d). Regions with a high nuclear area and elevated GATA3 expression notably overlap, suggesting that this tumor exhibits molecular heterogeneity, which to some degree is morphologically expressed.
To investigate this hypothesis, we measured the Pearson correlation between the expression of GATA3 and nuclear area in neoplastic cells (Figure 2.e). We observe a significant correlation (R=0.47), which is also observed in other genes and morphological features (Figure 2.f). Overall, out of the 12 human-interpretable features we analyzed, we found the highest association with gene expression for size-related features, while features involving topology, shape, and cell distribution had a lower correlation (R<0.2). A similar analysis in another IDC sample (Appendix Figure 5.b,c) confirmed these observations. In particular, we found the highest associations between nuclear size and expression for the genes FLNB (R=0.49) and TPD52 (R=0.49), both involved in breast tumor growth and proliferation Bandaru et al. (2014); Ren et al. (2021), and FOXA1 (R=0.47), a known prognostic factor associated with better survival Badve et al. (2007); Wolf et al. (2007).
Such analysis highlights how HEST-1k can be used for identifying fine-grained morphological correlates of expression. Additional analyses of morphological and molecular tumor heterogeneity can be leveraged at larger scale using similar approaches.
7 HEST for multimodal representation learning
Access to spatially-resolved expression–morphology pairs unlocks new directions for multimodal representation learning. Several problem statements can be explored, such as cross-modal alignment and retrieval, multimodal fusion, etc. Here, we fine-tune CONCH (strong performer on HEST-Benchmark) on five Xenium invasive breast cancer cases (four ductal and one lobular case) using contrastive alignment. We hypothesize that the resulting breast cancer-specific patch encoder, termed CONCH-FT, can better encode the underlying molecular landscape associated with disease-specific morphologies. To validate the hypothesis, CONCH-FT is benchmarked on an independent breast cancer cohort for molecular subtyping against its non-finetuned version.
[…]
We evaluate the resulting CONCH-FT model to predict ER, PR, and HER2 expression status (binary) from WSIs in the BCNB dataset Xu et al. (2021) (n=1,058 WSIs). To generate a slide representation for a WSI, we take the average of the patch embedding in the WSI, which is subsequently mapped to the expression status using logistic regression (Table 2). We observe that CONCH-FT outperforms CONCH on most metrics, demonstrating that pan-tissue histology patch encoders can be further fine-tuned to obtain better tissue-specific patch encoders. This is further validated by the larger rank induced by the patch embedding space Garrido et al. (2023) for CONCH-FT, suggesting better expressivity of the patch embeddings. While these results are based on only five WSIs, we anticipate additional benefits when training with larger disease-specific cohorts.
Table 2: CONCH fine-tuning on invasive breast cancer. Logistic regression evaluation for ER/PR/HER2 expression status on BCNB (n=1,058 WSIs). A WSI is represented by the average of the patch embeddings within each WSI. We report the mean ± standard deviation computed over all folds (or patients) for ROC-AUC (AUC) and balanced accuracy (Bal.acc.). Best is bold.
Rank ER PR HER2
AUC Bal.acc. AUC Bal.acc. AUC Bal.acc.
CONCH 144.66 0.881 0.745 0.810 0.698 0.715 $\bm{0.624}$
CONCH-FT $\bm{146.47}$ $\bm{0.884}$ $\bm{0.752}$ $\bm{0.818}$ $\bm{0.714}$ $\bm{0.724}$ 0.615
8 Discussion
Summary. We assembled HEST-1k, a dataset comprising paired spatial transcriptomics, H&E-stained whole-slide images, and comprehensive metadata built from public and internal cohorts. HEST-1k includes 1,108 samples, encompassing 1.5 million spots and over 60 million cells. The scale and comprehensiveness of HEST-1k, supported by the HEST-Library, enable exploring directions such as biomarker discovery and multimodal representation learning. Additionally, motivated by the need for new diverse and challenging patch-level benchmarks, we curated the HEST-Benchmark, a set of ten tasks covering nine cancer types and ten organs for gene expression prediction from histology. HEST-Benchmark revealed new insights into the predictive capabilities of foundation models for histology, such as the added benefits of vision-language alignment Lu et al. (2024).
Limitations. Our study includes limitations. First, research data, such as those generated in spatial transcriptomic, are inherently noisy. While we tried to minimize “label” noise (e.g., by re-estimating image magnification and alignment, and unifying cancer samples using oncotree code taxonomy), staining and compression artifacts, varying acquisition protocols, among others can negatively impact the quality of HEST-1k. Second, batch effects (on both the imaging and transcriptomic sides) can be significant across samples, datasets, and technologies. This study does not quantify these batch effects or explore methods to mitigate them. Lastly, although the HEST-Library was designed for versatility, it cannot cover all existing formats and should rather be viewed as a blueprint for processing ST data in a consistent and unified manner.
Future work. Spatial transcriptomics is rapidly evolving with new datasets frequently published. As they become available, we will keep updating HEST-1k with new resources. This study merely starts to uncover the potential of HEST-1k for advancing translational research and biomarker discovery, and we plan to explore these capabilities further. Additionally, the prospects for multimodal representation learning with HEST-1k are promising and are expected to grow with the addition of more data.
Acknowledgements
We thank Dr. Maxime Meylan for his insights and guidance on accessing data published in Meylan et al. (2022). S.J.W. was supported by the Helmholtz Association under the joint research school “Munich School for Data Science - MUDS” and the Add-on Fellowship of the Joachim Herz Foundation.
Checklist
  1. 1.
    Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] Each claim: HEST-1k, HEST-Library, HEST-Benchmark, HEST for discovery and multimodal fine-tuning are supported by dedicated sections in the main text, in addition to supplementary information provided in the appendix. In addition, all the code to reproduce results is made available.
  2. 2.
    Did you describe the limitations of your work? [Yes] We discuss limitations in the Discussion.
  3. 3.
    Did you discuss any potential negative societal impacts of your work? [Yes] We discuss potential negative societal impacts in the section Ethical considerations, intended usage, and license.
  4. 4.
    Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  5. 5.
    Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] In the abstract, we provide a link to access the HEST website and the HEST-Library on GitHub. HEST-Library includes a link to download all data to run the HEST-Benchmark. Finally, we provide all metadata associated with HEST-1k in a CSV as part of the supplementary material. Processed ST and WSIs will be made available upon acceptance.
  6. 6.
    Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] When relevant, we provide training details, such as in the HEST-Benchmark.
  7. 7.
    Did you report error bars? [Yes] HEST-Benchmark results include standard deviation computed from cross-validation across all patients.
  8. 8.
    Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes]
  9. 9.
    If your work uses existing assets, did you cite the creators? [Yes] All public resources used in this study are cited in Appendix Table A2, A4, A6, A7, A8, A9 and A10.
  10. 10.
    Did you mention the license of the assets? [Yes] Metadata associated with HEST-1k includes the license under which data were originally published. We ensured that the reported license allowed for distributing and creating derivatives of the data.
  11. 11.
    Did you include any new assets either in the supplemental material or as a URL? [Yes] As part of HEST-1k, we include internal datasets (see Appendix Table A9).
  12. 12.
    Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [No] We used public resources for which the license was allowing redistributing the work. Users are welcome to inspect the individual IRBs of each publicly available resource.
  13. 13.
    Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] We manually ensured that none of the published and distributed data includes personally identifiable information or offensive content, such as personal health information.
Appendix A Ethical considerations, intended usage and license
All resources provided as part of this study are strictly for research purposes and must not be utilized to support any diagnostic procedures. Users are hereby notified that the nuclear segmentation and classification components are derived from a publicly available model. Consequently, this model should not be regarded as the definitive standard, and users should exercise particular caution when utilizing this part of the dataset. Despite our efforts to exclude sensitive information, such as patient names, addresses, and social security numbers, users are expressly prohibited from attempting to reverse engineer the data to extract any confidential patient information. On the presumption that users will adhere to the aforementioned restrictions, we have not identified any potential adverse societal impacts that could arise from using HEST-1k.
The dataset is hosted on HuggingFace Dataset webpage. All instructions are provided in the main README of HEST-Library. From there, users can choose to download HEST-1k in its entirety, or a subset (e.g., only breast cancer samples). HEST-1k, HEST-Benchmark, and HEST library are released under the Attribution-NonCommercial-ShareAlike 4.0 International license (CC BY-NC-SA 4.0 Deed) [^9] https://creativecommons.org/licenses/by-nc-sa/4.0/.
Appendix B Background
This study connects two fields: (1) computational pathology, which primarily uses routinely acquired clinical data to determine outcomes such as disease diagnosis from H&E-stained digitized tissue sections, and (2) spatial transcriptomics, which has, so far, been confined to biological research and aims to identify new biomarkers predictive of disease progression or treatment response, among others.
B.1 Computational pathology
[…]
Such frameworks have been shown to achieve better or similar performance than humans for Gleason grading in prostate cancerBulten et al. (2022), metastasis detection in lymph nodesBejnordi et al. (2017), determining the origin of a cancer of unknown primaryLu et al. (2021a), predicting heart transplant rejectionLipkova et al. (2022), among others.
B.2 Spatial Transcriptomics (ST)
ST enables the measurement of gene activity and the mapping to its corresponding location on the tissue. In this study, we gather samples from two ST paradigms: sequencing-based (ST, Visium, Visium HD) and imaging-based (Xenium).
[…]
Xenium: Xenium is an imaging-based spatial-profiling technology offering higher resolution compared to its sequencing counterpart, Visium. Xenium imaging-based paradigm breaks free from the resolution limitations of its sequencing-based counterpart, it directly performs in situ RNA capturing on tissue sections by imaging fluorescent RNA markers derived from padlock probes and rolling circle amplification chemistry, hence offering sub-cellular expression resolution. As of 2024, Xenium cannot perform whole transcriptome measurement and is limited to gene panels of up to 580 genes.
Appendix C HEST
C.1 HEST-1k
We provide a comprehensive description of all publicly available and internal cohorts integrated into HEST-1k.
Table A1: HEST-1k data overview. All included samples include a license that allows sharing and redistributing. National Center for Biotechnology Information: NCBI.
Resource Number of datasets Number of samples Size (GB, raw)
10x Genomics 79 97 225
Mendeley 9 118 181
Spatial-Research 4 139 18
Zenodo 3 17 11
NCBI 42 677 265
Internal 3 28 60
Miscellaneous 2 32 65
Table A2: Datasets gathered from 10x Genomics portal. $n$: number of samples in the cohort.
Collection name Organ Technology $\bm{n}$ Num. genes
Adult Mouse Brain (FFPE) Brain Visium 1 19,465
Adult Mouse Brain Coronal Section (Fresh Frozen) 1 Brain Visium 1 32,285
Adult Mouse Brain Coronal Section (Fresh Frozen) 2 Brain Visium 1 32,285
Adult Mouse Kidney (FFPE) Kidney Visium 1 19,465
Adult Mouse Olfactory Bulb Brain Visium 1 32,285
FFPE Human Breast using the Entire Sample Area Breast Xenium 2 541
FFPE Human Breast with Custom Add-on Panel Breast Xenium 2 541
FFPE Human Breast with Pre-designed Panel Breast Xenium 2 541
FFPE Human Pancreas with Xenium Multimodal Cell Segmentation Pancreas Xenium 1 541
Fresh Frozen Mouse Colon with Xenium Multimodal Cell Segmentation Bowel Xenium 1 541
Fresh Frozen Visium on CytAssist: Human Breast Cancer, Probe-Based Whole Transcriptome Profiling Breast Visium 1 18,085
Fresh Frozen Visium on CytAssist: Mouse Brain, Probe-Based Whole Transcriptome Profiling Brain Visium 1 19,465
Human Bone and Bone Marrow Data with Custom Add-on Panel Bone Xenium 3 541
Human Brain Cancer, 11 mm Capture Area (FFPE) Brain Visium 1 18,085
Human Breast Cancer (Block A Section 1) Breast Visium 1 33,538
Human Breast Cancer (Block A Section 2) Breast Visium 1 33,538
Human Breast Cancer: Ductal Carcinoma In Situ, Invasive Carcinoma (FFPE) Breast Visium 1 17,943
Human Breast Cancer: Targeted, Immunology Panel Breast Visium 1 1,056
Human Breast Cancer: Visium Fresh Frozen, Whole Transcriptome Breast Visium 1 36,601
Human Breast Cancer: Whole Transcriptome Analysis Breast Visium 1 32,285
Human Cerebellum: Targeted, Neuroscience Panel Brain Visium 1 1,186
Human Cerebellum: Whole Transcriptome Analysis Brain Visium 1 1,186
Human Cervical Cancer (FFPE) Cervix Visium 1 17,943
Human Colon Preview Data (Xenium Human Colon Gene Expression Panel) Bowel Xenium 2 541
Human Colorectal Cancer, 11 mm Capture Area (FFPE) Bowel Visium 1 18,085
Human Colorectal Cancer: Targeted, Gene Signature Panel Bowel Visium 1 1,142
Human Colorectal Cancer: Whole Transcriptome Analysis Bowel Visium 1 36,601
Human Glioblastoma: Targeted, Pan-Cancer Panel Brain Visium 1 1,253
Human Glioblastoma: Whole Transcriptome Analysis Brain Visium 1 36,601
Human Heart Heart Visium 1 36,601
Human Heart Data with Xenium Human Multi-Tissue and Cancer Panel Heart Xenium 1 541
Human Intestine Cancer (FPPE) Bowel Visium 1 17,943
Human Kidney Preview Data (Xenium Human Multi-Tissue and Cancer Panel) Kidney Xenium 2 541
Human Kidney, 11 mm Capture Area (FFPE) Kidney Visium 1 18,085
Human Liver Data with Xenium Human Multi-Tissue and Cancer Panel Liver Xenium 2 541
Human Lung Cancer (FFPE) Lung Visium 1 18,085
Human Lung Cancer, 11 mm Capture Area (FFPE) Lung Visium 1 18,085
Human Lymph Node Lymph node Visium 1 36,601
Table A3: Datasets gathered from 10x Genomics portal. Continuation.
Collection name Organ Technology $\bm{n}$ Num. genes
Human Ovarian Cancer (FFPE) Ovary Visium 1 17,943
Human Ovarian Cancer, 11 mm Capture Area (FFPE) Ovary Visium 1 18,085
Human Prostate Cancer, Acinar Cell Carcinoma (FFPE) Prostate Visium 1 17,943
Human Prostate Cancer, Adenocarcinoma with Invasive Carcinoma (FFPE) Prostate Visium 1 17,943
Human Skin Data with Xenium Human Multi-Tissue and Cancer Panel Skin Xenium 2 541
Human Skin Preview Data (Xenium Human Skin Gene Expression Panel with Custom Add-On) Skin Xenium 1 541
Human Skin Preview Data (Xenium Human Skin Gene Expression Panel) Skin Xenium 1 541
Human Tonsil Data with Xenium Human Multi-Tissue and Cancer Panel Lymph node Xenium 2 541
Mouse Bone Data with Custom Add-on Panel Bone Xenium 3 541
Mouse Brain Coronal Section 1 (FFPE) Brain Visium 1 19,465
Mouse Brain Coronal Section 2 (FFPE) Brain Visium 1 19,465
Mouse Brain Section (Coronal) Brain Visium 1 31,053
Mouse Brain Serial Section 1 (Sagittal-Anterior) Brain Visium 1 31,053
Mouse Brain Serial Section 1 (Sagittal-Posterior) Brain Visium 1 31,053
Mouse Brain Serial Section 2 (Sagittal-Anterior) Brain Visium 1 31,053
Mouse Brain Serial Section 2 (Sagittal-Posterior) Brain Visium 1 31,053
Mouse Embryo, 11 mm Capture Area (FFPE) Embryo Visium 1 19,465
Mouse Kidney Section (Coronal) Kidney Visium 1 31,053
Mouse Tissue Microarray in 3x3 Layout with 1 mm Edge to Edge Spacing (FFPE) Lung/Brain Visium 1 19,465
Mouse Tissue Microarray in 3x3 Layout with 2 mm Edge to Edge Spacing (FFPE) Lung/Brain Visium 1 19,465
Mouse Tissue Microarray in 5x5 Layout with 1 mm Edge to Edge Spacing (FFPE) Kidney/Brain Visium 1 19,465
Normal Human Prostate (FFPE) Prostate Visium 1 17943
Pancreatic Cancer with Xenium Human Multi-Tissue and Cancer Panel Pancreas Xenium 1 538
Preservation Method Comparison on CytAssist: FFPE Mouse Brain (Sagittal), 11 mm Capture Area Brain Visium 1 19,465
Preservation Method Comparison on CytAssist: Fixed Frozen Mouse Brain (Sagittal), 11 mm Capture Area Brain Visium 1 19,465
Preservation Method Comparison on CytAssist: Fresh Frozen Mouse Brain (Sagittal), 11 mm Capture Area Brain Visium 1 19,465
Preservation Method Comparison on Visium CytAssist: FFPE Mouse Brain (Sagittal), 11 mm Capture Area Brain Visium 1 19,465
Preservation Method Comparison on Visium CytAssist: Fixed Frozen Mouse Brain (Sagittal), 11 mm Capture Area Brain Visium 1 19,465
Preservation Method Comparison on Visium CytAssist: Fresh Frozen Mouse Brain (Sagittal), 11 mm Capture Area Brain Visium 1 19,465
Preview Data: FFPE Human Lung Cancer with Xenium Multimodal Cell Segmentation Lung Xenium 1 541
Visium CytAssist Gene Expression Libraries of Post-Xenium Human Colon Cancer (FFPE) Bowel Visium 4 18,085
Visium CytAssist Gene Expression Libraries of Post-Xenium Mouse Brain (FF) Brain Visium 4 19,465
Visium CytAssist, Mouse Embryo, 11 mm Capture Area (FFPE) Embryo Visium 1 19,465
Whole Mouse Pup Preview Data (Xenium Mouse Tissue Atlassing Panel) Whole organism Xenium 1 541
Table A4: Datasets gathered from NCBI.
Publication Organ Technology $\bm{n}$ Num. genes
10X Visium Spatial transcriptomics of murine colon at d14 (mucosa healing) in B cell sufficient-deficient mice Parigi et al. (2022) Bowel Visium 2 31,053
10X Visium Spatial transcriptomics of murine colon in steady state and during recovery after DSS colitis Parigi et al. (2022) Bowel Visium 2 31,053
A Spatial Transcriptomic atlas of the human kidney papilla identifies significant immune injury in patients with stone disease Canela et al. (2023) Kidney Visium 7 36,601
A cellular hierarchy in melanoma uncouples growth and metastasis Karras et al. (2022) Skin Visium 3 31,053
A new epithelial cell subpopulation predicts response to surgery, chemotherapy, and immunotherapy in bladder cancer Abdel-Hafiz et al. (2023); Gouin et al. (2021) Bladder Visium 4 33,538
A novel model of binge ethanol exposure reveals enhanced neurodegeneration with advanced age Anton et al. (2024) Brain Visium 4 32,285
A single-cell transcriptomic analysis of endometriosis Fonseca et al. (2023) Uterus Visium 2 36,601
Distinct mesenchymal cell states mediate prostate cancer progression Pakula et al. (2024) Prostate Visium 2 32,589
Epithelial Plasticity and Innate Immune Activation Promote Lung Tissue Remodeling following Respiratory Viral Infection Beppu et al. (2023) Lung Visium 1 32,285
Gene expression within a human choroidal neovascular membrane using spatial transcriptomics Voigt et al. (2023) Eye Visium 5 36,601
Genome-wide Spatial Expression Profiling in Formalin-fixed Tissues Gracia Villacampa et al. (2021) Kidney Visium 14 33,538
High resolution mapping of the tumor microenvironment using integrated single-cell, spatial and in situ analysis Janesick et al. (2023) Breast Xenium 4 541
Identification of TREM1+CD163+ myeloid cells as a deleterious immune subset in HCC Giraud et al. (2022) Liver Visium 2 36,601
Integration of spatial and single cell transcriptomics localizes epithelial-immune cross-talk in kidney injury Ferreira et al. (2021) Kidney Visium 4 33,538
Molecular Atlas of the Adult Mouse Brain Ortiz et al. (2020) Brain Spatial Transcriptomics 75 23,371
Regional differential gene expression analyses of brains from four 24w-old Nf1+- mice Brain Visium 4 32,285
SARS-CoV-2 Niches in Human Placenta Revealed by Spatial Transcriptomics Barrozo et al. (2023) Uterus Visium 16 36,612
Schwann Cells Shape Tumor Cells and Cancer-Associated Fibroblasts in the Pancreatic Ductal Adenocarcinoma Microenvironment Xue et al. (2023) Pancreas Visium 4 20,615
Single Cell and Spatial Analysis of Human Squamous Cell Carcinoma Ji et al. (2020) Skin Spatial Transcriptomics 12 17,138
Single cell profiling of primary and paired metastatic lymph node tumors in breast cancer patients Liu et al. (2022) Lymph node Visium 4 33,931
Table A5: Datasets gathered from NCBI. Continuation.
Publication Organ Technology $\bm{n}$ Num. genes
Single-cell and spatial transcriptomics characterisation of the immunological landscape in the healthy and PSC human liver Andrews et al. (2024) Liver Visium 4 36,601
Single-nucleus Ribonucleic Acid-sequencing and Spatial Transcriptomics Reveal the Cardioprotection of Shexiang Baoxin Pill (MUSKARDIA) in Mice with Myocardial Ischemia-Reperfusion Injury Lin et al. (2023) Heart Visium 2 32,285
Spatial Multimodal Analysis: MALDI-MSI and Spatial Transcriptomics within the same tissue section Vicari et al. (2023) Brain Visium 19 32,285
Spatial RNA sequencing of regenerating mouse hindlimb muscle McKellar et al. (2021) Muscle Visium 3 33,217
Spatial Total RNA-Sequencing of regenerating mouse hindlimb muscle and Type 1-Lang reovirus-infected mouse heart McKellar et al. (2023) Muscle Visium 7 55,414
Spatial localization with Spatial Transcriptomics for an atlas of healthy and injured cell states and niches in the human kidney Lake et al. (2023) Kidney Visium 23 33,538
Spatial sequencing of Foreign body granuloma Krausgruber et al. (2023) None Visium 1 15,524
Spatial transcriptomics landscape of non-communicable inflammatory skin diseases Schäbitz et al. (2022) Skin Visium 59 20,613
Spatial transcriptomics of adenoid cystic carcinoma of the lacrimal gland Moeyersoms et al. (2023) Eye Visium 1 17,943
Spatial transcriptomics of the mouse brain across three age groups Brain Visium 6 32,285
Spatial transcriptomics reveal unnresolved wound repair as potential driver of PFA Ependymoma progression Fu et al. (2023) Brain Visium 14 36,601
Spatiotemporal dynamics of molecular pathology in amyotrophic lateral sclerosis Silas Maniatis et al. (2019) Spinal cord Spatial Transcriptomics 302 12,572
The neurons that restore walking after paralysis Kathe et al. (2022) Spinal cord Visium 16 22,127
Visium spatial transcriptomics analysis of lacrimal gland during chronic inflammation progression Mauduit et al. (2022) Eye Visium 4 32,285
YAP Drives Assembly of a Spatially Colocalized Cellular Triad Required for Heart Renewal Li et al. (2024) Heart Visium 2 32,285
Zika virus co-opts miRNA networks to persist in placental microenvironments detected by spatial transcriptomics Barrozo et al. (2024) Placenta Visium 8 32,298
Mouse model Heptablastoma spatial transcriptomics Pilet et al. (2023) Liver Visium 10 31,053
Table A6: Datasets gathered on Spatial-Research.
Publication Organ Technology $\bm{n}$ Num. genes
A spatiotemporal organ-wide gene expression and cell atlas of the developing human heart Asp et al. (2019) Heart Spatial Transcriptomics 19 39,739
Integrating spatial gene expression and breast tumour morphology via deep learning He et al. (2020) Breast Spatial Transcriptomics 68 16,744
Spatial deconvolution of HER2-positive breast cancer delineates tumor-associated cell type interactions Andersson et al. (2021) Breast Spatial Transcriptomics 36 15,045
Visualization and analysis of gene expression in tissue sections by spatial transcriptomics Ståhl et al. (2016a) Brain Spatial Transcriptomics 16 16,573
Table A7: Datasets gathered on Mendeley.
Publication Organ Technology $\bm{n}$ Num. genes
Ex-ST Andrusivova and Fan (2023) Brain Visium 5 31,053
Genome-wide spatial expression profiling in formalin-fixed tissues Villacampa et al. (2021) Brain Visium 15 31,053
Human ileum, Visium Mirzazadeh et al. (2021) Bowel Visium 4 33,538
Human squamous cell carcinoma Abalo et al. (2021) Skin Visium 4 33,538
Prostate needle biopsies pre- and post-ADT: Count matrices, histological-, and Androgen receptor immunohistochemistry images Marklund (2022) Prostate Spatial Transcriptomics 24 26,437
Spatially resolved clonal copy number alterations in benign and malignant tissue Erickson et al. (2022) Prostate Visium 23 33,538
Spatially resolved transcriptomic profiling of degraded Mirzazadeh et al. (2023) Bowel Visium 35 17,943
spatialRNAseq heart raw suppdata Heart Visium 4 54,848
spatialRNAseq ileum raw suppdata Bowel Visium 4 54,848
Table A8: Datasets gathered on Github and the Human Cell Atlas data explorer.
Publication Organ Technology $\bm{n}$ Num. genes
Transcriptome-scale spatial gene expression in the human dorsolateral prefrontal cortexMaynard et al. (2021) Brain Visium 12 33,538
A spatially resolved atlas of the human lung characterizes a gland-associated immune nicheMadissoon et al. (2023) Lung Visium 20 17,922
Table A9: Internal datasets.
Publication Organ Technology $\bm{n}$ Num. genes
Prostate ST Internal Prostate Visium 4 17,943
Tertiary lymphoid structures generate and propagate anti-tumor antibody-producing plasma cells in renal cell cancer Meylan et al. (2022) Lymph node Visium 24 17,943
Table A10: Datasets gathered on Zenodo.
Publication Organ Technology $\bm{n}$ Num. genes
Charting the Heterogeneity of Colorectal Cancer Consensus Molecular Subtypes using Spatial Transcriptomics: datasets Valdeolivas et al. (2023) Bowel Visium 14 36,601
Demo 10x Visium dataset for STQ zen (2024a) Skin Visium 1 68,886
Spotiphy: generative modeling in single-cell spatial whole transcriptomics zen (2024b) Brain Visium 2 32,285
C.2 HEST-Library
The HEST-Library helps transform unstructured spatial transcriptomics and histology data into a unified format. An overview of the HEST-Library is provided in Suppl. Figure 3. An example of fiducial detection is presented in Suppl. Figure 4.
Figure 3: Overview of HEST-Library functionalities. HEST was designed to transform legacy data scrapped in multiple public repositories, such as NCBI, into unified and actionable objects that can easily be integrated into multimodal computational pipelines.
Figure 4: Fiducial detection and automatic alignment in Visium samples. Corner fiducials on 6.5$\times$6.5mm and 11mm×$\times$11mm Visium slides are automatically detected with a finetuned Yolov8 model, the spots coordinates are then derived if at least 3 of the 4 corner fiducials are detected. This process enables automatically estimating the pixel resolution.
C.3 HEST-Benchmark
Gene selection, Random Forest, and Ridge regression models: We learn a regression model that maps the patch embeddings of each encoder to its corresponding gene expression profile. The Random Forest model is based on 70 trees (sklearn implementation). The Ridge regression uses a fixed $L2$ regularization coefficient $\lambda$ to $100/MC$, where $M$ is the embedding dimension and $C=50$ is the number of targets trained with the Regularized Least-Squares Routine solver (sklearn implementation). We train the regression models to predict a panel of 50 most variable genes tailored to each task. Specifically, for each task, we select the 50 most variable genes across all spots and samples, after excluding the genes that have non-zero counts in less than 10% of the spots.
Benchmark task description: We provide complementary information on each task introduced as part of the HEST-Benchmark.
Table A11: Overview of the HEST-Benchmark. Each task involves predicting the expression levels of the 50 most variable genes from 112$\times$112 $\mu$m H&E-stained image patches centered on each spatial transcriptomics spot. The tasks are formulated as multivariate regression problems. The Oncotree code describes the cancer type diagnosed in samples, e.g., PAAD denotes pancreatic adenocarcinoma. Additional information is provided in the Appendix.
Task ID Oncotree Organ Number of Patients Number of Samples Technology
Task 1 Janesick et al. (2023) IDC Breast 4 4 Xenium
Task 2 PRAD Prostate 2 23 Visium
Task 3 PAAD Pancreas 3 3 Xenium
Task 4 SKCM Skin 2 2 Xenium
Task 5 Valdeolivas et al. (2023) COAD Colon 3 6 Visium
Task 6 Valdeolivas et al. (2023) READ Rectum 2 4 Visium
Task 7 Meylan et al. (2022) ccRCC Kidney 24 24 Visium
Task 8 Giraud et al. (2022) HCC Liver 2 2 Visium
Task 9 LUAD Lung 2 2 Xenium
Task 10 Liu et al. (2022) IDC Axillary lymph nodes 4 4 Visium
Task 1: Expression prediction in invasive ductal carcinoma (breast cancer, IDC). We used all publicly available Xenium samples available on 10x Genomics (“FFPE Human Breast using the Entire Sample Area”, 2 patients) and two samples published in  Janesick et al. (2023) (2 patients). All samples are FFPE sections imaged with the Xenium pipeline v1.
Task 2: Expression prediction in prostate adenocarcinoma (prostate cancer, PRAD). We used all 23 Visium samples (fresh frozen sections) from 2 patients published in Erickson et al. (2022). Both patients were diagnosed with prostatic acinar adenocarcinoma with a (4+3) Gleason score (ISUP group 4).
Task 3: Expression prediction in pancreatic adenocarcinoma (pancreatic cancer, PAAD). We used 3 samples from 3 different patients from 10x Genomics (“FFPE Human Pancreas with Xenium Multimodal Cell Segmentation” and “Pancreatic Cancer with Xenium Human Multi-Tissue and Cancer Panel”). All samples are FFPE sections processed with Xenium pipeline v1.
Task 4: Expression prediction in skin cutaneous melanoma (skin cancer, SKCM). We used 2 samples from 2 different patients from 10x Genomics website (“Human Skin Data with Xenium Human Multi-Tissue and Cancer Panel”). All samples are FFPE sections processed with Xenium pipeline v1.
Task 5: Expression prediction in colon adenocarcinoma (colon cancer, COAD). We used 6 COAD samples from 3 different patients published in Valdeolivas et al. (2023). All samples are fresh frozen sections processed with Visium.
Task 6: Expression prediction in rectal adenocarcinoma (rectum cancer, READ). We used 4 READ samples from different 2 patients published in Valdeolivas et al. (2023). All samples are fresh frozen sections processed with Visium.
Task 7: Expression prediction in clear cell renal cell carcinoma (kidney cancer, ccRCC). We used all 24 ccRCC samples from 24 different patients published in Meylan et al. (2022). All samples are fresh frozen sections processed with Visium.
Task 8: Expression prediction in hepatocellular carcinoma (liver cancer, HCC). We used 2 HCC liver samples from 2 different patients published in Giraud et al. (2022). All samples are fresh frozen sections processed with Visium.
Task 9: Expression prediction in lung adenocarcinoma (lung cancer, LUAD). We used 2 LUAD samples from 2 different patients from 10x genomics (“Preview Data: FFPE Human Lung Cancer with Xenium Multimodal Cell Segmentation”). All samples are fresh frozen sections processed with Xenium pipeline v1.
Task 10: Expression prediction in axillary lymph nodes in IDC patients. We used 4 axillary lymph node samples from 2 IDC patients published in Liu et al. (2022). All samples are fresh frozen sections processed with Visium.
Table A12: State-of-the-art foundation models for histology evaluated on HEST-Benchmark. ResNet-50: 23M parameters, ResNet-18: 11M, ResNet-152: 795M, ViT-B: Base (86M), L: Large (307M), G: Giant (1.13B)
Name Number of Number of Magnification Model Training
slides patches recipe
ResNet50-IN Lu et al. (2021b) N/A 1M N/A ResNet-50 Supervised
KimiaNet Riasatian et al. (2021) 7K 240K 20x ResNet-18 Supervised
Ciga Ciga et al. (2022) UNK UNK 10x, 20x ResNet-50 Supervised
CTransPath Wang et al. (2021) 32k 17M 10x UNK MoCov2
Remedis Azizi et al. (2023) 10 10M 20x ResNet-152 iBOT
Phikon Filiot et al. (2023) 6k 43.3M 20x ViT-B iBOT
PLIP Huang et al. (2023) N/A 800k Many ViT-B CLIP
UNI Chen et al. (2024) 100k 100M 20x ViT-L DINOv2
CONCH Lu et al. (2024) N/A 1.17M Many ViT-B CoCa
GigaPath Xu et al. (2024) 171k 1.3B 20x ViT-G CoCa
We provide a brief description of each patch encoder assessed on the HEST-Benchmark.
ResNet50 (IN) Lu et al. (2020): This model uses a ResNet50 backbone He et al. (2016) trained on ImageNet dataset Deng et al. (2009). Following prior work Lu et al. (2020), the patch embedding corresponds to the deep representation at the penultimate layer before classification.
Ciga Ciga et al. (2022): This model uses a ResNet-50 backbone He et al. (2016) trained on public histology datasets such as TCGA and Cameleyon16 using SimCLR.
CTransPath Wang et al. (2022): This model uses a “Tiny” Swin Transformer backbone Liu et al. (2021) with a window size of 14 (Swin-T/14, 28 million parameters) pretrained on TCGA and PAIP datasets (17 million images) using MoCoV3 Chen et al. (2021).
Remedis Azizi et al. (2023): This model uses a ResNet-152$\times$2 (232 million parameters) initialized with the “Big Transfer”-medium protocol Kolesnikov et al. (2020) on ImageNet-22K and pretrained with SimCLR Chen et al. (2020) on TCGA.
Phikon Filiot et al. (2023): This model uses a Vision Transformer-Base (ViT-B, 86 million parameters) Dosovitskiy et al. (2021) trained on TCGA data using iBOT Zhou et al. (2022).
PLIP Huang et al. (2023): This model uses a ViT-B (86 million parameters) trained on 208 thousand pathology images paired with natural language descriptions extracted from Twitter and trained using CLIP Radford et al. (2021).
UNI Chen et al. (2024): This model uses a ViT-Large (ViT-L, 307 million parameters) Dosovitskiy et al. (2021) trained on 100 million histology images (over 100,000 slides) from proprietary and public data using DINOv2 Oquab et al. (2023).
CONCH Lu et al. (2024): This model uses a ViT-B (86 million parameters) trained on 1.17 million histology image–caption pairs extracted from online educational and research resources using CoCa Yu et al. (2022).
GigaPath Xu et al. (2024): This model uses a ViT-Giant (1.13 billion parameters) trained on 1.3 billion image patches from 171,189 WSIs at 20$\times$ magnification using DINOv2.
Appendix D HEST for multimodal representation learning
CONCH-FT model, a ViT-Base model initialized with CONCH weights, was fine-tuned for 50 epochs using a cosine learning rate scheduler, with a base learning rate of $10^{-4}$ for the image encoder and $10^{-3}$ for the expression encoder. Only the last 3 layers of the model were fine-tuned, with a layer-wise learning decay rate of 0.7. For training with the infoNCE loss, a contrastive temperature of $10^{-2}$ and batch size of 1,024 pairs of patch and transcriptomics were used. A combination of random horizontal/vertical flip and color jittering was employed for image patch augmentation.
The rank of the embedding space, also referred to as smooth rank measure Garrido et al. (2023), measures the quality of the embeddings produced from encoders trained in unsupervised or self-supervised manners. Given the patch embedding matrix $H\in\mathcal{R}^{N\times d}$ and $d<N$, where $N$ is the number of patches and $d$ is the feature dimension, we compute the rank as the entropy of the $d$ L1-normalized singular values of $H$.
Appendix E HEST for discovery
Cells were segmented and classified using CellViT Hörst et al. (2024). To find the gene expression profile of each neoplastic cell, we matched each cell to its corresponding cell index in Xenium by assigning the index for which the distance between the cell centroids was the smallest. After matching all neoplastic cells, only those cells for which the assignment was unique were kept. After this filtering step, an average of 91% of the cells per sample were kept while 9% of the cells were discarded. The discarded cells likely correspond to over-segmentation errors by CellViT.
Figure 5: HEST for biomarker discovery: Analysis of an invasive ductal carcinoma Xenium sample. a. IDC Xenium sample with neoplastic nuclei overlaid in red ($n_{c}$=342,018 detected nuclei). Six randomly selected regions with CellViT segmentation of the neoplastic nuclei. Black scale bar represents 30 $\mu$m. b. Pearson correlation between the major axis length of neoplastic nuclei and the log1p-normalized expression of TPD52. c. Analogous analysis between nuclear area and FLNB expression.
Appendix F Datasheet for HEST-1k
We provide a DataSheet for HEST-1k that summarizes the contributions, analyses, and intended usages presented in the study.
F.1 Motivation for dataset creation
  • Why was the dataset created? HEST-1k was designed with three key applications: (1) multimodal representation learning of histology and transcriptomics, (2) biomarker discovery, and (3) benchmarking of histopathology patch encoder models. Despite many publicly available resources, no existing unified and user-friendly formatting was available to bring ST in the world of deep learning.
  • What (other) tasks could the dataset be used for? Are there obvious tasks for which it should not be used? Users are welcome to propose new creative ways to use the dataset. Users are not allowed to try to retrieve patient information from the existing data. A dedicated section is allocated to discuss ethical considerations and intended usage.
  • Has the dataset been used for any tasks already? If so, where are the results so others can compare (e.g., links to published papers)? All samples that were made public as part of a publication (peer-review or not) are reported in the metadata attached with HEST-1k.
  • Who funded the creation of the dataset? The work was funded by an R35 grant.
F.2 Dataset composition
  • What are the instances? The modalities used in this study are histopathology whole-slide images, gene expression data, and derivatives of these two modalities, such as nuclear segmentation maps.
  • Are relationships between instances made explicit in the data Each whole-slide image maps to a unique gene expression profile in an unequivocal way.
  • What data does each instance consist of? Imaging data consists of Generic TIFF objects stored in pyramidal format, and gene expression data consists of scanpy objects. Derivatives are stored in JSON and Hierarchical Data Format (HDF) files.
  • Is there a label/target associated with instances? If the instances are related to people, are subpopulations identified (e.g., by age, gender, etc.), and what is their distribution? Each sample pair (slide and expression profile) is associated with comprehensive metadata. All metadata information is thoroughly described in the main paper. Age and gender are only reported in a subset of cases.
  • Is everything included or does the data rely on external resources? (e.g., websites, tweets, datasets) If external resources, a) are there guarantees that they will exist, and remain constant, over time; b) is there an official archival version. Are there licenses, fees or rights associated with any of the data? We provide all data as part of HEST-1k release. In addition, a link to the original data is provided in the metadata. Each sample is associated with a license as provided by the original publication, where we ensured that the reported license allowed for distributing and creating derivatives of the data. Should the author of the original publication would like to remove their samples from HEST-1k, we will update the data promptly.
  • Are there recommended data splits or evaluation measures? HEST-1k comes with the HEST-Benchmark, a series of tasks for gene expression prediction from histology images. All patient-stratified splits are specified in the attached comma-separated values (CSV) files.
  • What experiments were initially run on this dataset? Have a summary of those results and, if available, provide the link to a paper with more information here. At the time of submission, all experiments run with HEST-1k are part of this study. The reader can refer to the main text for a thorough description of all experiments (see HEST-Benchmark, HEST for discovery, HEST for multimodal representation learning).
F.3 Data collection process
  • How was the data collected? The data were manually inspected and curated by the authors of the present study.
  • Who was involved in the data collection process? All authors of the present study were involved in the data collection, inspection, and curation. The reader can refer to the original publication to understand how the data were orginally acquired.
  • Over what time-frame was the data collected? Does the collection time-frame match the creation time-frame? The original data comprise publications from 2016 to 2024. As the dataset grows, more recent publications might be included in HEST-1k.
  • Does the dataset contain all possible instances? Or is it, for instance, a sample (not necessarily random) from a larger set of instances? All pairs of gene expression data and whole-slide images of the underlying studies were included and are unique.
  • Is there information missing from the dataset and why? (this does not include intentionally dropped instances; it might include, e.g., redacted text, and withheld documents) Is this data missing because it was unavailable? Original publications include some missing information, such as the alignment file between the slide and the expression profile. We developed computational tools to minimize missing information to reach near-complete metadata.
  • Are there any known errors, sources of noise, or redundancies in the data? All whole-slide images have been manually inspected. The quality from one sample to another varies significantly, for instance, due to poor staining, compression artifact, lower resolution, etc. Gene expression data are inherently noisy. Users can decide to apply post-hoc normalization methods to reduce noise.
F.4 Data preprocessing
  • What preprocessing/cleaning was done? All whole-slide images were converted into a pyramidal TIFF objects with re-estimated pixel resolution. All alignment files have been manually inspected and included if missing. All gene expression data have been transformed into scanpy objects following the same process.
  • Was the “raw” data saved in addition to the preprocessed/cleaned data? (e.g., to support unanticipated future uses) Raw data are downloaded, but not publicly shared. In the case of public samples, users can re-download them using the metadata provided as part of the dataset release.
  • Is the preprocessing software available? Yes, the source code to preprocess HEST-1k is made publicly available as part of the HEST library.
F.5 Dataset distribution
  • How is the dataset distributed? HEST-1k is distributed using internal storage. We have several internal copies to ensure data preservation over the years.
  • When will the dataset be released/first distributed? The dataset will be made public with the camera-ready version of this publication. This study serves as the canonical reference.
  • What license (if any) is it distributed under? Are there any copyrights on the data? The dataset is distributed under the Attribution-NonCommercial-ShareAlike 4.0 International license (CC BY-NC-SA 4.0 Deed).
  • Are there any fees or access/export restrictions? No access/export restrictions unless they violates the terms of the above-mentioned license (CC BY-NC-SA 4.0 Deed).
  • Who is supporting/hosting/maintaining the dataset? How does one contact the owner/curator/manager of the dataset? The dataset is maintained by the authors of the publication.
  • Will the dataset be updated? How often and by whom? How will updates/revisions be documented and communicated (e.g., mailing list, GitHub)? Is there an erratum? The dataset might evolve as additional samples become publicly available. Dataset versioning will be put in place. A website will clearly state what the latest version of the dataset contains and keep track of the changes in each version.
  • If the dataset becomes obsolete how will this be communicated? The corresponding website will state if the dataset becomes obsolete.
  • Is there a repository to link to any/all papers/systems that use this dataset? There is no repository to link papers that use HEST-1k. Users are required to cite HEST-1k if they use it in their own research.
  • If others want to extend/augment/build on this dataset, is there a mechanism for them to do so? If so, is there a process for tracking/assessing the quality of those contributions. What is the process for communicating/distributing these contributions to users? Users are welcome to contact us if they would like to provide additional data that meet the standards of the data contained in HEST-1k. We do not have a dedicated system to communicate these contributions. Newly added data will be tracked in the versioning.
F.6 Legal and ethical considerations
  • If the dataset relates to people (e.g., their attributes) or was generated by people, were they informed about the data collection? (e.g., datasets that collect writing, photos, interactions, transactions, etc.) HEST-1k does not include patient information (such as name, address, etc.).
  • If it relates to other ethically protected subjects, have appropriate obligations been met? (e.g., medical data might include information collected from animals) For animal samples (Mus musculus tissue), we refer to the original publication for an in-depth analysis.
  • If it relates to people, were there any ethical review applications/reviews/approvals? (e.g. Institutional Review Board applications) For human tissue, we refer to the original publication for an in-depth analysis. Internal cohorts were ethically reviewed and collected as part of dedicated IRBs.
  • If it relates to people, could this dataset expose people to harm or legal action? (e.g., financial social or otherwise) What was done to mitigate or reduce the potential for harm? No, patients cannot be linked to the corresponding histology and gene expression profile.
  • If it relates to people, does it unfairly advantage or disadvantage a particular social group? In what ways? How was this mitigated? Most datasets do not include specific demographics. When reported, we include this information in the metadata associated with each sample. To our knowledge, the representation of HEST-1k does not unfairly advantage or disadvantage a particular social group.
  • Does the dataset contain information that might be considered sensitive or confidential? (e.g., personally identifying information) No.
  • Does the dataset contain information that might be considered inappropriate or offensive? No.
Appendix G Author statement
The authors of this paper bear all responsibility in case of violation of rights associated with HEST-1k, HEST-Library, and HEST-Benchmark.
Table A13: HEST-Benchmark results. Ridge regression evaluation on public state-of-the-art patch encoders. Model performance measured with Pearson correlation. We report the mean $\pm$ standard deviation computed over all folds (or patients). Best is bold, second best is underlined.
ResNet50 KimiaNet Ciga CTransPath Remedis Phikon PLIP UNI CONCH Gigapath
IDC 0.446 0.455 0.425 0.499 0.496 0.526 0.450 $\bm{0.544}$ 0.541 $\underline{0.542}$
±$\pm 0.04$$\pm 0.03$$\pm 0.04$$\pm 0.06$$\pm 0.07$$\pm 0.08$$\pm 0.04$$\pm 0.07$$\pm 0.07$$\pm 0.09$
PRAD 0.276 0.323 0.258 $0.290$ 0.264 0.249 0.262 0.257 $\bm{0.353}$ 0.301
±$\pm 0.06$$\pm 0.04$$\pm 0.06$$\pm 0.07$$\pm 0.08$$\pm 0.13$$\pm 0.02$$\pm 0.08$$\pm 0.01$$\pm 0.03$
PAAD 0.343 0.374 0.364 $\underline{0.383}$ 0.288 0.360 0.358 0.377 $\bm{0.424}$ 0.323
±$\pm 0.07$$\pm 0.08$$\pm 0.07$$\pm 0.07$$\pm 0.05$$\pm 0.07$$\pm 0.06$$\pm 0.06$$\pm 0.04$$\pm 0.00$
SKCM $\underline{0.413}$ 0.357 0.381 0.404 0.412 0.369 0.337 0.343 $\bm{0.546}$ 0.190
±$\pm 0.08$$\pm 0.05$$\pm 0.06$$\pm 0.07$$\pm 0.04$$\pm 0.11$$\pm 0.08$$\pm 0.06$$\pm 0.05$$\pm 0.05$
COAD $\underline{0.116}$ 0.103 0.093 0.113 0.075 0.114 0.108 0.110 $\bm{0.135}$ 0.085
±$\pm 0.05$$\pm 0.04$$\pm 0.02$$\pm 0.05$$\pm 0.04$$\pm 0.05$$\pm 0.05$$\pm 0.06$$\pm 0.06$$\pm 0.03$
READ 0.067 0.058 0.081 0.090 0.077 0.114 0.084 0.124 $\bm{0.154}$ $\underline{0.125}$
±$\pm 0.06$ 0.05 ±$\pm 0.05$$\pm 0.08$$\pm 0.07$$\pm 0.07$$\pm 0.06$$\pm 0.04$$\pm 0.04$$\pm 0.05$
ccRCC 0.210 0.190 0.180 0.214 0.219 $\bm{0.253}$ 0.174 $\underline{0.240}$ 0.221 0.226
±$\pm 0.05$$\pm 0.06$$\pm 0.04$$\pm 0.04$$\pm 0.04$$\pm 0.05$$\pm 0.04$$\pm 0.05$$\pm 0.03$$\pm 0.05$
HCC $\underline{0.039}$ 0.031 0.027 0.024 0.016 0.038 0.032 0.017 $\bm{0.046}$ 0.008
±$\pm 0.02$$\pm 0.01$$\pm 0.02$$\pm 0.02$$\pm 8.71$$\pm 0.01$$\pm 0.03$$\pm 0.00$$\pm 0.01$$\pm 0.00$
LUAD 0.400 0.419 0.417 0.403 0.311 $\underline{0.422}$ 0.406 0.371 $\bm{0.493}$ 0.297
±$\pm 0.06$$\pm 0.07$$\pm 0.04$$\pm 0.07$$\pm 0.04$$\pm 0.06$$\pm 0.01$$\pm 0.11$$\pm 0.03$$\pm 0.08$
LYM-IDC 0.203 0.227 0.218 0.209 0.167 0.215 0.215 0.224 $\bm{0.247}$ 0.191
±$\pm 0.05$$\pm 0.04$$\pm 0.04$$\pm 0.04$$\pm 0.04$$\pm 0.04$$\pm 0.05$$\pm 0.03$$\pm 0.04$$\pm 0.02$
Average 0.251 0.253 0.244 0.263 0.232 0.266 0.243 0.260 $\bm{0.316}$ 0.229

### v2

@@
 Abstract
-Spatial transcriptomics (ST) enables interrogating the molecular composition of tissue with ever-increasing resolution, depth, and sensitivity.
+Spatial transcriptomics enables interrogating the molecular composition of tissue with ever-increasing resolution and sensitivity.
 However, costs, rapidly evolving technology, and lack of standards have constrained computational methods in ST to narrow tasks and small cohorts.
-In addition, the underlying tissue morphology as reflected by H&E-stained whole slide images (WSIs) encodes rich information often overlooked in ST studies.
-Here, we introduce HEST-1k, a collection of 1,108 spatial transcriptomic profiles, each linked to a WSI and metadata.
-HEST-1k was assembled using HEST-Library from 131 public and internal cohorts encompassing 25 organs, two species (Homo Sapiens and Mus Musculus), and 320 cancer samples from 25 cancer types.
-HEST-1k processing enabled the identification of 1.5 million expression–morphology pairs and 60 million nuclei.
-HEST-1k is tested on three use cases: (1) benchmarking foundation models for histopathology (HEST-Benchmark), (2) biomarker identification, and (3) multimodal representation learning.
-HEST-1k, HEST-Library, and HEST-Benchmark can be freely accessed via https://github.com/mahmoodlab/hest.
+In addition, the underlying tissue morphology, as reflected by H&E-stained whole slide images (WSIs), encodes rich information often overlooked in ST studies.
+Here, we introduce HEST-1k, a collection of 1,229 spatial transcriptomic profiles, each linked to a WSI and extensive metadata.
+HEST-1k was assembled from 153 public and internal cohorts encompassing 26 organs, two species (Homo Sapiens and Mus Musculus), and 367 cancer samples from 25 cancer types.
+HEST-1k processing enabled the identification of 2.1 million expression–morphology pairs and over 76 million nuclei.
+To support its development, we additionally introduce the HEST-Library, a Python package designed to perform a range of actions with HEST samples.
+We test HEST-1k and Library on three use cases: (1) benchmarking foundation models for pathology (HEST-Benchmark), (2) biomarker exploration, and (3) multimodal representation learning.
+HEST-1k, HEST-Library, and HEST-Benchmark can be freely accessed at https://github.com/mahmoodlab/hest.
 **footnotetext: Equal contribution
 1 Introduction
-Advances in molecular profiling enable spatially-resolved gene expression analysis with increasingly large gene panels, enhanced spatial resolution, and greater sensitivity Asp et al.
-(2020); Rao et al.
-(2021).
-From the early days of bulk RNA sequencing constrained by its coarse resolution and limited gene panels, spatially-resolved technologies have progressed to achieve whole-transcriptome sequencing at sub-cellular resolution Li and Wang (2021).
-In cancer research, spatial transcriptomics (ST) holds particular promise for characterizing the tumor microenvironment, a key element in understanding disease progression and treatment response Ståhl et al.
-(2016b); Marx (2021); de Visser and Joyce (2023); Xiao and Yu (2021).
-With the large amount of transcriptomics data generated by a single ST sample (e.g., $>$10 million transcripts are detected in a typical 10x Genomics Xenium assay), computational methods are often used to identify promising targets and biomarkers, such as based on gene expression clustering or gene-set enrichment analysis Rao et al.
-(2021).
-However, high costs and rapidly evolving technology have constrained computational methods to narrow tasks and data cohorts of only few patients Meylan et al.
-(2022); Janesick et al.
-(2023); Valdeolivas et al.
-(2023).
-Consequently, we observe a lack of standardized resources and unified formats for storing ST, which limits the development of deep learning models on a large scale Moses and Pachter (2022).
+Advances in molecular profiling enable spatially-resolved gene expression analysis with increasingly large gene panels, enhanced spatial resolution, and greater sensitivity [asp2020spatially, rao2021exploring].
+From the early days of bulk RNA sequencing constrained by its coarse resolution and limited gene panels, spatially-resolved technologies have progressed to achieve whole-transcriptome sequencing at sub-cellular resolution [li2021from].
+In cancer research, spatial transcriptomics (ST) holds particular promise for characterizing the tumor microenvironment, a key element in understanding disease progression and treatment response [staahl2016visualization, marx2021method, devisser2023evolving, xiao2021tumor].
+With the large amount of transcriptomics data generated by a single ST sample (e.g., $>$10 million transcripts are detected in a typical 10x Genomics Xenium assay), computational methods are often used to uncover promising biomarkers, such as employing clustering methods for cell phenotyping [rao2021exploring].
+However, high costs and rapidly evolving technology have constrained computational methods to narrow tasks and data cohorts of only a few patients [meylan2022tertiary, janesick2023high, valdeolivas2023charting].
+Consequently, we observe a lack of standardized resources and unified formats for handling ST, which limits the development of deep learning models on a large scale [moses2022museum].
 In addition, the underlying tissue morphology, traditionally visualized in hematoxylin and eosin (H&E)-stained tissue sections (whole-slide images, WSIs), is often overlooked in ST studies, despite encoding valuable information.
-In particular, pairs of ST and WSI enable analyzing expression changes in their morphological context, which may facilitate the discovery of morphological biomarkers (e.g., changes in nuclear shape) that correspond to gene regulation patterns.
-In translational medicine, ST could be employed for identifying biomarkers, which are then morphologically validated and characterized by widely deployed clinical modalities such as H&E.
-Alternatively, pairs of ST and WSI can enable multimodal tissue representation learning for joint modeling of the morphomolecular signature of tissue at a scale and resolution beyond bulk RNA sequencing Chen et al.
-(2022).
-Finally, the development of “foundation models” for encoding histopathology image patches Wang et al.
-(2021); Filiot et al.
-(2023); Chen et al.
-(2024); Huang et al.
-(2023) has increased the need for new, diverse, and challenging benchmarks beyond diagnostic tasks.
-Using ST, new tasks can be defined for predicting gene expression from histology.
+In particular, pairs of ST and WSI enable analyzing expression changes in their morphological context, which may facilitate the identification of morphological biomarkers (e.g., changes in nuclear shape) that correspond to gene regulation patterns.
+Alternatively, pairs of ST and WSI can enable multimodal tissue representation learning for joint modeling of the morphomolecular signature of tissue at a scale and resolution beyond bulk RNA sequencing [chen2022pan].
+Finally, the development of “foundation models” for encoding histopathology images [wang2021transpath, filiot2023scaling, chen2024towards, huang2023visual, lu2024towards] has increased the need for new, diverse, and challenging benchmarks beyond diagnostic tasks.
+Using ST, new tasks can be defined to predict gene expression changes from histology.
 Here, we introduce HEST-1k, a collection of paired ST and H&E-stained WSIs curated from public and internal cohorts (Figure 1.a).
-HEST-1k comprises 1,108 samples from 131 cohorts encompassing 25 organs, two species (Homo Sapiens and Mus Musculus), and 320 cancer samples from 25 subtypes.
-Processing all samples in HEST-1k resulted in 1.5 million expression–morphology pairs and 60 million detected nuclei.
-With new cohorts frequently made public, we also introduce the HEST-Library, a Python package for querying HEST-1k and assembling new legacy data (Figure 1.b).
-We highlight the potential of HEST-1k through three use cases: (i) benchmarking foundation models for histology using the HEST-Benchmark, a set of ten tasks (nine human cancer types and ten organs) for gene expression prediction from histology and evaluated on ten state-of-the-art models (Figure 1.c), (ii) a proof-of-concept demonstrating the use of HEST-1k for biomarker characterization (Figure 1.d), and (iii) a proof-of-concept for expression-guided fine-tuning of foundation models for histology (Figure 1.e).
-Figure 1: Overview of HEST.
+HEST-1k comprises 1,229 samples from 153 cohorts encompassing 26 organs, two species (Homo Sapiens and Mus Musculus), and 367 cancer samples from 25 different subtypes.
+Processing all samples in HEST-1k resulted in 2.1 million expression–morphology pairs and 76 million detected nuclei.
+With new cohorts frequently made public, we also introduce the HEST-Library, a Python package for interacting with HEST-1k data and assembling new samples as they become available (Figure 1.b).
+We highlight the potential of HEST-1k through three use cases: (i) benchmarking foundation models for histology using the HEST-Benchmark, a set of nine tasks (eight human cancer types and nine organs) for gene expression prediction from histology and evaluated on eleven state-of-the-art models (Figure 1.c), (ii) a proof-of-concept demonstrating the use of HEST-1k for biomarker characterization (Figure 1.d), and (iii) a proof-of-concept for expression-guided fine-tuning of foundation models for histology (Figure 1.e).
+Figure 1: The HEST environment.
 a.
-Description of HEST-1k, a dataset of $n$=1,108 paired Spatial Transcriptomics, H&E-stained whole-slide images and metadata.
+Overview of HEST-1k, a dataset of $n$=1,229 paired spatial transcriptomics, H&E-stained whole-slide images and metadata.
+“Pathological” cases refer to non-tumor/non-cancer samples; “Tumor” refers to non-cancer samples.
 b.
@@
 c., d., e.
-Potential applications of HEST-1k include benchmarking foundation models for histology (c.), biomarker discovery (d.) and multimodal representation learning (e.).
+Applications of HEST-1k include benchmarking foundation models for histology (c.), biomarker exploration (d.) and multimodal representation learning (e.).
 2 Related work
 Libraries for ST analysis.
-Libraries to process, visualize, and analyze ST have been built around two core pipelines: Scanpy Wolf et al.
-(2018) (and the Anndata format) in Python and Seurat Butler et al.
-(2018) in R.
-Scanpy has served as the foundation for several subsequent developments such as Squidpy Palla et al.
-(2022) for spatial data exploration at cellular-, gene-, and morphological-level, SpatialData Marconato et al.
-(2024) for multi-technology integration and deep learning interfacing, and STlearn Pham et al.
-(2023) for cell-cell interactions and spatiotemporal trajectory analyses.
-In R, Seurat Butler et al.
-(2018) has been consolidated with packages such as BayeSpace Zhao et al.
-(2021) for clustering and spot super-resolution, and Giotto suites Chen et al.
-(2023) for preprocessing, data integration, and visualization of multiple ST technologies.
-10x genomics also includes software analytics through the Xenium and Visium Explorer pipelines for multimodal visualization, nuclear segmentation, and cell deconvolution.
+Libraries to process, visualize, and analyze ST have been built around two core pipelines: Scanpy [wolf2018scanpy] (and the Anndata format) in Python and Seurat [butler2018integrating] in R.
+Scanpy has served as the foundation for several subsequent developments such as Squidpy [palla2022squidpy] for spatial data exploration at cellular-, gene-, and morphological-level, SpatialData [marconato2024spatialdata] for multi-technology integration and deep learning interfacing, STlearn [pham2023robust] for cell-cell interactions and spatiotemporal trajectory analyses, and SOPA [blampey2024sopa] for designing multistep pipelines.
+In R, Seurat [butler2018integrating] has been consolidated with packages such as BayeSpace [zhao2021spatial] for clustering and spot super-resolution, and Giotto suites [chen2023giotto] for preprocessing, data integration and visualization of multiple ST technologies.
+10x Genomics also includes proprietary software analytics through the Xenium and Visium Explorer pipelines for multimodal visualization, nuclear segmentation, and cell deconvolution.
 However, none of these pipelines were designed to handle the diversity of legacy data, where datasets can suffer from missing or incorrect data, such as alignment mismatches, incorrect pixel resolution, inconsistent image file formats, etc.
 Molecular profile prediction from H&E.
-Molecular profiling from histology images has been explored both at (1) slide-level to predict bulk molecular status/changes from a WSI, and at (2) patch-level to predict local molecular status/changes from regions-of-interest.
-(1) Slide-level profiling has been proposed to predict the presence of gene mutations Saldanha et al.
-(2023); Wagner et al.
-(2023); Kather et al.
-(2020); Echle et al.
-(2021); Fu et al.
-(2020); Wang et al.
-(2019); Loeffler et al.
-(2022), microsatellite instability Wagner et al.
-(2023); Kather et al.
-(2019b), and gene expression changes Schmauch et al.
-(2020); El Nahhas et al.
-(2024); He et al.
-(2020); Alsaafin et al.
-(2023), among others.
-The motivation is two-fold: First, patient screening to substitute or emulate costly clinical molecular assays, and second, to identify morphological correlates of molecular alterations for discovering novel biomarkers.
+Molecular profiling from histology images has been explored both at (1) slide-level to predict bulk molecular status/changes from a WSI and at (2) patch-level to predict local molecular status/changes from regions-of-interest.
+(1) Slide-level profiling has been explored to predict the gene mutations [saldanha2023self, wagner2023fully, kather2020pan, echle2021deep, fu2020pan, wang2019predicting, loeffler2022artificial], microsatellite instability [wagner2023fully, kather2019deep], and gene expression changes [schmauch2020deep, el2024regression, he2020integrating, alsaafin2023learning], among others.
+The motivation is two-fold: First, patient screening to substitute or complement costly clinical molecular assays, and second, to identify morphological correlates of molecular alterations for discovering novel biomarkers.
 Such studies can be conducted on large patient cohorts as they mainly rely on data generated by the routine clinical workflow (e.g., using TCGA cohorts with $>$11,000 cases from 33 cancer types).
-(2) With ST, several works have explored predicting expression changes from regions-of-interest He et al.
-(2020); Xie et al.
-(2023); Pang et al.
-(2021); Zhao et al.
-(2024); Rahaman et al.
-(2023); Monjo et al.
-(2022); Mondol et al.
-(2023); Chung et al.
-(2024); Wang et al.
-(2024).
-Due to limited cohort sizes (typically one to ten patients), transfer learning has become the norm using pretrained networks based on ConvNets He et al.
-(2020) or Vision Transformers Xie et al.
-(2023); Pang et al.
-(2021).
-Transcriptomic measurements are inherently noisy, hence the development of context-based methods that integrate global and local information from surrounding ST spots Chung et al.
-(2024); Zhao et al.
-(2024); Wang et al.
-(2024).
-While recent technologies offer near-single-cell resolution (such as Visium HD and Xenium), legacy assays operate at a more coarse resolution, which can be upsampled using super-resolution techniques Zhao et al.
-(2021); Bergenstråhle et al.
-(2022); Zhang et al.
-(2024).
+(2) With ST, several works have explored predicting expression changes from regions-of-interest [he2020integrating, xie2023spatially, pang2021leveraging, zhao2024hist2cell, rahaman2023breast, monjo2022efficient, mondol2023hist2rna, chung2024accurate, wang2024m2ort].
+Due to limited cohort sizes (typically one to ten patients), transfer learning has become the norm using pretrained models based on ConvNets [he2020integrating] or Vision Transformers [xie2023spatially, pang2021leveraging].
+Due to the inherent noise found in transcriptomic measurements, several methods have been developed for integrating context that can account for global and local information from surrounding ST spots [hu2021spagcn, chung2024accurate, zhao2024hist2cell, wang2024m2ort].
+While recent technologies offer near-single-cell resolution (such as Visium HD and Xenium), legacy assays operate at a more coarse resolution, which can be upsampled using super-resolution techniques [zhao2021spatial, bergenstrahle2022super, zhang2024inferring].
 The potential clinical and research implications of such methods are still being explored, with HEST-1k potentially catalyzing their large-scale development.
 Foundation models in pathology.
-A fundamental task in computational pathology is to extract general-purpose embeddings of image patches (typically 256$\times$256-pixel regions) that can then be used for downstream tasks, such as diagnosis or prognosis prediction.
-To achieve this, self-supervised learning (SSL) has been extensively applied  Koohbanani et al.
-(2021); Filiot et al.
-(2023); Vorontsov et al.
-(2023); Wang et al.
-(2021); Kang et al.
-(2023); Ciga et al.
-(2022); Wang et al.
-(2022); Azizi et al.
-(2023); Chen et al.
-(2024); Lu et al.
-(2024), such as based on the DINOv2 framework Oquab et al.
-(2023).
-Models for general-purpose patch encoding are trained on increasingly large and more diverse patient cohorts with models growing in model size (e.g., Virchow uses a ViT-Huge trained on more than 1.5M WSIs Vorontsov et al.
-(2023)).
+A fundamental task in computational pathology is to extract general-purpose embeddings of image patches (typically 256$\times$256 to 512$\times$512-pixel regions) that can then be used for downstream tasks, such as diagnosis or prognosis prediction.
+To achieve this, self-supervised learning (SSL) has been extensively applied [koohbanani2021self, filiot2023scaling, vorontsov2024foundation, wang2021transpath, kang2023benchmarking, ciga2022self, wang2022transformer, azizi2023robust, chen2024towards, lu2024towards, jaume2024transcriptomics, jaume2024multistain], such as based on the DINOv2 framework [oquab2023dinov2].
+General-purpose patch encoders are trained on increasingly large and diverse patient cohorts (e.g., UNI [chen2024towards] uses a ViT-Large trained on 100k WSIs, Virchow [vorontsov2024foundation] uses a ViT-Huge trained on more than 1.5M WSIs).
+Recently, vision-language encoders designed for pathology have also been proposed [gamper2021multiple, lu2023visual, huang2023visual, lu2024towards, lu2024multimodal] and rely on large-scale paired data scraped from social media, textbooks, or publications.
 As the number of such models rapidly increases, new, diverse, and challenging benchmarks are needed to replace or complement well-established tasks where performance has saturated.
-HEST-Benchmark aims to address this by offering a set of ten patch-level tasks for gene expression prediction from histology.
+HEST-Benchmark aims to address this by offering a set of nine patch-level tasks for gene expression prediction from histology.
 Patch-level benchmarks in histopathology.
 Early task and dataset contributions in computational pathology revolved around classifying small regions of interest.
-Over the years, a variety of benchmarks have been established: In prostate cancer, Gleason grading at pixel- and patch-level has been widely explored, with public resources such as PANDA Bulten et al.
-(2022), AGCC Huo et al.
-(2022), DiagSet Koziarski et al.
-(2024), and SICAPv2 Silva-Rodríguez et al.
-(2020).
-In colorectal cancer, datasets have been proposed for tissue classification, such as HunCRC Pataki et al.
-(2022), UniToPatho Barbano et al.
-(2021), MHIST Wei et al.
-(2021), and CRC-100k Kather et al.
-(2019a).
-In breast cancer, morphological subtyping has been studied (e.g., for atypical ductal hyperplasia detection), such as BACH Aresta et al.
-(2019), BRACS Brancati et al.
-(2021), and BreakHis Spanhol et al.
-(2015), and for lymph node metastasis detection with Patch CAMELYON (pCAM) Veeling et al.
-(2018), respectively.
-However, the performance on many of these datasets has saturated; for instance, Gleason scoring reaches similar or better performance than pathologists Bulten et al.
-(2022), which limits objective comparisons of new methods and hinders well-informed model selection for developing better features.
+Over the years, a variety of benchmarks have been established: In prostate cancer, Gleason grading at pixel- and patch-level has been widely explored, with public resources such as AGGC [huo2022comprehensive], DiagSet [koziarski2024diagset], and SICAPv2 [silvarodriguez2020going].
+In colorectal cancer, datasets have been proposed for tissue classification, such as HunCRC [pataki2022huncrc], UniToPatho [barbano2021unitopatho], MHIST [wei2021petri], and CRC-100k [kather2019predicting].
+In breast cancer, morphological subtyping has been vastly explored (e.g., for atypical ductal hyperplasia detection), such as BACH [aresta2019bach], BRACS [brancati2021bracs], and BreakHis [spanhol2015dataset], and for lymph node metastasis detection with Patch CAMELYON (pCAM) [veeling2018rotation], respectively.
+However, the performance on many of these datasets has saturated; for instance, Gleason scoring reaches similar or better performance than pathologists [bulten2022artificial], which limits objective comparisons of new methods and hinders well-informed model selection for developing better features.
 Instead, HEST-Benchmark provides a collection of diverse and challenging tasks that enable assessing the predictive capabilities of foundation models for histology.
@@
 To this end, we extracted all publicly available cohorts that provide ST with H&E-stained whole-slide images.
-Specifically, we harvested data from 10x Genomics public datasets (TENX) [^1] https://www.10xgenomics.com/datasets, Mendeley (MEND) [^2] https://data.mendeley.com/, Spatial-Research (SPA) [^3] https://www.spatialresearch.org/, Zenodo (ZEN) [^4] https://zenodo.org/, the National Center for Biotechnology Information (NCBI) [^5] https://www.ncbi.nlm.nih.gov/gds/, GitHub [^6] https://github.com/, the Human Cell Atlas [^7] https://data.humancellatlas.org/, and internal data cohorts.
+Specifically, we harvested data from 10x Genomics public datasets (TENX) [^1] https://www.10xgenomics.com/datasets, Mendeley (MEND) [^2] https://data.mendeley.com/, Spatial-Research (SPA) [^3] https://www.spatialresearch.org/, Zenodo (ZEN) [^4] https://zenodo.org/, the National Center for Biotechnology Information (NCBI) [^5] https://www.ncbi.nlm.nih.gov/gds/, GitHub [^6] https://github.com/, the Human Cell Atlas [^7] https://data.humancellatlas.org/, BioStudies [^8] https://www.ebi.ac.uk/biostudies/, HTAN [^9] https://humantumoratlas.org/, and internal data cohorts.
 A summary of all sources is provided in Appendix Table A1 with specifics in Appendix Table A2, A4, A6,A7,A8,A9, and A10.
 3.1 Metadata
-As Spatial Transcriptomics were not intended for large-scale computational research, they are provided in various formats (e.g., images can be in JPG or TIFF format, with or without cross-modal alignment files) and resolutions.
+As spatial transcriptomics experiments were not intended for large-scale computational research, they are provided in various formats (e.g., images can be in JPG or TIFF format, with or without cross-modal alignment files) and resolutions.
 We unified all data with comprehensive metadata with generic-, histology-, and expression-related descriptors for all samples.
 Generic: We provide the reference to the original publication, download link, year of publication, license, and sample species.
-Each sample is then categorized as either healthy, cancer, treated (which refers to a post-compound administration), genetically modified (mostly knock-out mouse samples), or other diseases (with an additional description).
+Each sample is then categorized as either healthy, cancer, tumor (non-cancer), treated (which refers to a post-compound administration), genetically modified (mostly knock-out mouse samples), or pathological (i.e., non-tumorous with extra specification).
 All cancer samples were unified using the OncoTree code, a taxonomy of cancer types provided by the Memorial Sloan Kettering Cancer Center [^10] https://oncotree.mskcc.org.
-Finally, we include the organ using the highest level of the OncoTree taxonomy as a reference.
-Expression: We include the number of genes and spots per sample, the spot resolution and spacing, the total number of reads, and the mean number of reads per spot.
+Finally, we provide the organ using the highest level of the OncoTree taxonomy as a reference.
+Expression: We report the number of genes and spots per sample, the spot resolution and spacing, the total number of reads, and the mean number of reads per spot.
 We additionally provide the transcriptomic technology (ST, Visum, Visium HD, or Xenium).
 Histology: We provide the image resolution (in $\mu$m/pixel) and magnification as 10$\times$ (1.15 to 0.8 $\mu$m/px), 20$\times$ (0.8 to 0.4 $\mu$m/px) and 40$\times$ (0.4 to 0.1 $\mu$m/px).
-All images with a resolution lower than 1.15 $\mu$m/px were discarded to ensure an acceptable image quality.
-In addition, we provide the image size at the highest resolution, and tissue preparation protocol (frozen or formalin-fixed paraffin-embedded, FFPE).
+All images with a pixel size higher than 1.15 $\mu$m/px were discarded to ensure an acceptable image quality.
+In addition, we provide the image size at the highest resolution and the tissue preparation protocol (frozen or formalin-fixed paraffin-embedded, FFPE).
 3.2 Histology
-All tissue sections were normalized and transformed into a generic TIFF object, a pyramidal image that can easily be integrated into computational frameworks using Openslide or viewers such as QuPath Bankhead et al.
-(2017).
+All tissue sections were normalized and transformed into a generic TIFF object, a pyramidal image that can easily be integrated into computational frameworks using Openslide or viewers such as QuPath [bankhead2017qupath].
 In addition, we provide a contour object that delineates all the tissue regions identified in the image.
-We developed a robust tissue detection method where we fine-tuned a DeepLabV3 Chen et al.
-(2017) model with an ImageNet-pretrained ResNet50 backbone on a set of annotated segmentation regions.
+We developed a robust tissue vs.
+background detection method where we fine-tuned a DeepLabV3 [chen2017rethinking] model with an ImageNet-pretrained ResNet50 backbone on a set of annotated segmentation regions (including pen marks, fiducials, multiple stains, artifacts, etc.).
 From the tissue segmentation, we extracted 224$\times$224-pixel patches at 20$\times$ magnification around each spot.
-This yielded 1.5 million valid patches for which a corresponding expression profile was derived.
+For Xenium samples, we generated “pseudo-Visium” spots by pooling transcripts on 55 $\times$ 55-$\mu$m patches without spacing.
+This yielded 2.1 million valid patches for which a corresponding expression profile was derived.
 Such patching can readily be used for various downstream tasks, such as employed in the HEST-Benchmark or multimodal fine-tuning of foundation models for histology (Section 5 and 7).
@@
 In addition to patching, we include nuclear segmentation that delineates each nucleus identified in all slides from HEST-1k.
-Specifically, we use CellViT Hörst et al.
-(2024), a state-of-the-art nuclear segmentation model that was trained on the PanNuke dataset Gamper et al.
-(2019, 2020).
+We used CellViT [horst2024cellvit], a state-of-the-art nuclear segmentation model that was trained on the PanNuke dataset [gamper2019pannuke, gamper2020pannuke].
 CellViT enables joint instance segmentation and classification of each nucleus into five classes: neoplastic epithelial, non-neoplastic epithelial, inflammatory, stromal, and necrotic.
-On average, we identified 55.5k nuclei per slide (standard deviation 91.3k nuclei), for a total of 59.6 million nuclei identified across all samples.
-Among those 17.6 million are classified as neoplastic, 21.5 million as stromal, 4.9 million as normal epithelial, 15.4 million as inflammatory, and 76 thousand as necrotic.
+On average, we identified 62.1k nuclei per slide, for a total of 76.4 million nuclei identified across all samples.
+Among those, 17.6 million are classified as neoplastic, 21.5 million as stromal, 4.9 million as normal epithelial, 15.4 million as inflammatory, and 76 thousand as necrotic.
 The resulting nuclear segmentation and classification can easily be visualized using QuPath (using geojson) or loaded as Python/R objects (using JSON).
+For all Xenium samples, we additionally provide the nuclear and cell segmentation derived from the DAPI staining finely aligned with the H&E slide.
 3.4 Gene expression
@@
 Each entry represents the raw transcript counts of a gene in a given spot.
-No additional normalization was conducted, and we let users explore various normalization strategies based on needs, e.g., using total count normalization, log-based re-weighting, etc.
+No additional normalization was conducted, and we let users explore various normalization strategies based on needs, e.g., using total count normalization, log-normalization, etc.
 In addition, we include metadata to specify the number of genes, the gene panel, and the tissue site.
+For all Xenium samples, we also provide the list of all measured transcripts with their exact 2D position in the tissue (aligned with the H&E slide).
 To use the expression in tandem with the WSI, an alignment file describing the mapping between the image and the spots is needed.
@@
 To address these limitations, we re-aligned all samples under the same unified format between the WSI and the corresponding expression profile.
-To this end, we developed an automatic alignment pipeline (see Section 4) and embedded the alignment in the scanpy object.
+For all Visium samples, we developed an automatic alignment pipeline based on fiducial detection (see Section 4) and embedded the alignment in the scanpy object.
+For all Xenium samples, we used the publicly available VALIS [gatenbee2023valis] pipeline for fine-grained image registration to align the DAPI image (aligned with the transcripts by design) and the H&E slide.
 4 HEST-Library
 The HEST-Library is built around scanpy and Anndata.
-At its core, the HEST-Library enables (1) assembling and querying HEST-1k, and (2) running the HEST-Benchmark (see Section 5).
-We describe its core functionalities, in particular for unifying legacy data.
+At its core, the HEST-Library enables (1) assembling and querying HEST-1k, (2) visualizing and mitigating batch effects, and (3) running the HEST-Benchmark (Section 5).
+We describe its core functionalities, particularly for unifying legacy data.
 Conversion to generic TIFF.
@@
 Pyramidal formats offer seamless integration with OpenSlide (commonly used in computational pathology pipelines) and QuPath (open-access software for WSI visualization and annotation).
-Automatic alignment.
+Automatic alignment in Visium.
 Spot alignment is crucial to ensure an accurate match between the ST spots and the WSI.
-While software such as LoupeBrowser enables manual alignment using the fiducials (in Visium), the process remains time-consuming when processing large batches of samples.
-Instead, we implemented an automatic fiducial detection algorithm based on Yolov8 Redmon et al.
-(2016) for processing Visium samples (Appendix Figure 4).
-Specifically, we manually annotated 119 fiducial regions that we further augmented using tissue and fiducials mixing.
-We then fine-tuned Yolov8 pretrained on COCO dataset.
-In samples without fiducials (e.g., early versions of Spatial Transcriptomics), we realigned using the provided spot position files.
+While software such as LoupeBrowser enables manual alignment using fiducials (i.e., reference markers placed at the corners of the capture area), the process remains time-consuming when processing large batches of samples.
+Instead, we implemented an automatic fiducial detection algorithm based on YOLOv8 [redmon2016you] for processing Visium samples (Appendix Figure 5).
+Specifically, we manually annotated 119 fiducial regions that we further augmented using tissue and fiducial mixing.
+We then fine-tuned YOLOv8 pretrained on the COCO dataset.
+In early versions that do not provide corner fiducials (e.g., STv1), we realigned using the provided spot position files.
+In Xenium, we use VALIS [gatenbee2023valis] to register the DAPI staining (aligned with the transcripts) with the H&E image.
 Automatic detection of image resolution.
-From the alignment and the spot resolution, we can infer the exact pixel resolution.
-We start by deriving the distance in pixel between two neighboring spots and leverage the known inter-spot distance in $\mu$m to estimate the pixel width in $\mu$m/px.
-For Xenium samples, we directly read the H&E alignment file which provides an affine transformation from the DAPI-stained image (with known pixel size) to the H&E image.
+From the alignment and the spot resolution, we can infer the exact pixel size.
+To this end, we compute the distance in pixel between two neighboring spots and leverage the known inter-spot distance in $\mu$m to estimate the pixel width in $\mu$m/px.
+For Xenium samples, we use the H&E alignment file provided as part of the assay, which provides an affine transformation from the DAPI-stained image (with known pixel size) to the H&E image.
 We then compared the self-reported image resolution and our re-estimations to manually inspect and correct discrepancies.
 Conversion to Anndata.
-ST data is provided in multiple formats, such as CSVs, MEX, TXT, h5, etc.
-We provide functions to unify a large set of existing formats into a AnnData object that stores the raw transcript counts as a matrix of genes by the number of spots.
+ST data is provided in multiple formats, such as CSV, MEX, TXT, h5, etc.
+We provide functions to unify a large set of existing formats into a Anndata object that stores the raw transcript counts as a matrix of genes by the number of spots, in addition to metadata about the samples (e.g., the (x,y) coordinates of each spot, the pixel resolution, etc.).
 Tissue segmentation and patching.
 We provide a tissue segmentation pipeline optimized for Visium/Xenium images.
-The segmented tissue can automatically be tessellated into fixed-size image patches at a predefined resolution around each spot.
+The segmentation can then automatically tessellate the tissue into fixed-size image patches at a predefined resolution (expressed in $\mu$m/px) around each spot.
 Automatic HEST-1k download.
-To facilitate downloading part or all of the HEST-1k dataset (in total 825 GB), we implemented an easy download option where the user can specify entries of the metadata, for instance, to query all human breast cancer cases.
+To facilitate downloading part or all of the HEST-1k dataset (over $>$1TB), we implemented an easy download option where the user can specify entries of the metadata, for instance, to query all human invasive breast cancer cases.
+Batch effect visualization and mitigation.
+We provide functions to help visualize batch effects using dimensionality reduction techniques with user-prompted stratification (e.g., tissue site, institution, disease, etc.).
+In addition, we provide a wrapper of well-established batch effect mitigation strategies (namely ComBat [zhang2020combat], Harmony [korsunsky2019harmony] and matching mutual nearest neighbors [haghverdi2018mnn]), which can be applied to a list of HEST samples.
 5 HEST-Benchmark
-From HEST-1k, we curated the HEST-Benchmark, a set of ten tasks for gene expression prediction from histology in human cancer samples.
-The goal is two-fold: (i) benchmarking foundation models for histology under a diverse and challenging benchmark, and (ii) understanding the predictive capabilities of state-of-the-art models in predicting expression from morphology.
-Compared to existing tasks (e.g., Camelyon16), HEST-Benchmark brings more morphological and data acquisition diversity, and more complex tasks as expression prediction is inherently challenging.
+From HEST-1k, we curated the HEST-Benchmark, a set of nine tasks for gene expression prediction from histology in human cancer samples.
+The goal is two-fold: (i) benchmarking foundation models for histology under a diverse and challenging benchmark and (ii) understanding the predictive capabilities of state-of-the-art models in predicting expression from morphology.
+Compared to existing tasks (e.g., Camelyon16 [bejnordi2017diagnostic]), the HEST-Benchmark brings increased morphological diversity and more complex challenges, particularly with the inherent difficulty of expression prediction.
 5.1 Task definition
-We define ten tasks with data from nine human cancers and ten organs (nine primary and one metastatic dataset), which include invasive ductal carcinoma (breast cancer, IDC, Task 1), prostate adenocarcinoma (prostate cancer, PRAD, Task 2), pancreatic adenocarcinoma (pancreatic cancer, PRAD, Task 3), skin cutaneous melanoma (skin cancer, SKCM, Task 4), colonic adenocarcinoma (colon cancer, COAD, Task 5), rectal adenocarcinoma (rectum cancer, READ, Task 6), clear cell renal cell carcinoma (kidney cancer, ccRCC, Task 7), hepatocellular carcinoma (liver cancer, HCC, Task 8), lung adenocarcinoma (lung cancer, LUAD, Task 9), and axillary lymph nodes in IDC (metastatic, LYM-IDC, Task 10).
+We define nine tasks with data from eight human cancers and nine organs (eight primary and one metastatic dataset), which include invasive ductal carcinoma (breast cancer, IDC, Task 1), prostate adenocarcinoma (prostate cancer, PRAD, Task 2), pancreatic adenocarcinoma (pancreatic cancer, PAAD, Task 3), skin cutaneous melanoma (skin cancer, SKCM, Task 4), colonic adenocarcinoma (colon cancer, COAD, Task 5), rectal adenocarcinoma (rectum cancer, READ, Task 6), clear cell renal cell carcinoma (kidney cancer, ccRCC, Task 7), lung adenocarcinoma (lung cancer, LUAD, Task 8), and axillary lymph nodes in IDC (metastatic, LYMPH-IDC, Task 9).
 Additional information is provided in Appendix Table A11.
@@
 In ccRCC, we use $k/2$-fold cross-validation due to the large number of patients.
-5.2 Evaluating foundation models for histology on HEST-Benchmark
-We use the HEST-Benchmark to evaluate ten patch encoders for histology.
-Namely, ResNet50 (IN) Lu et al.
-(2020) (ImageNet pretrained), KimiaNet Riasatian et al.
-(2021) (pretrained with supervised subtyping classification), Ciga Ciga et al.
-(2022) (SimCLR pretrained on public histology data), CTransPath Wang et al.
-(2022) (adapted MoCov3 pretrained on TCGA and PAIP), Remedis Azizi et al.
-(2023) (iBOT pretrained on TCGA), Phikon Filiot et al.
-(2023) (iBOT pretrained on TCGA), PLIP Huang et al.
-(2023) (visual-language model based on CLIP pretrained on histology data from Twitter), UNI Chen et al.
-(2024) (DINOv2 ViT-Large pretrained on public and private histology data), CONCH Lu et al.
-(2024) (visual-language model using CoCa pretrained on captions from publications and educational resources), and GigaPath Xu et al.
-(2024) (DINOv2 ViT-Giant pretrained on proprietary data).
+5.2 Evaluating foundation model for pathology
+We use the HEST-Benchmark to evaluate 11 foundation models for pathology.
+Namely, ResNet50 (IN) [lu2020data] (ImageNet pretrained), CTransPath [wang2022transformer] (adapted MoCov3 pretrained on TCGA and PAIP), Remedis [azizi2023robust] (SimCLR [chen2020simple] pretrained on TCGA), Phikon [filiot2023scaling] (iBOT pretrained on TCGA), UNI [chen2024towards] (DINOv2 ViT-Large pretrained on internal hospital data and GTEx), CONCH [lu2024towards] (visual-language model using CoCa pretrained on captions from publications and educational resources), GigaPath [xu2024wholeslide] (DINOv2 ViT-giant pretrained on proprietary data), Virchow [vorontsov2024foundation] (DINOv2 ViT-Huge pretrained on proprietary data), Virchow 2 [Zimmermann2024Virchow2SS] (DINOv2 ViT-Huge pretrained on proprietary data), H-Optimus-0 (DINOv2 ViT-giant pretrained on proprietary data), and UNIv1.5 (DINOv2 ViT-giant pretrained on public and proprietary data).
 Additional information is provided in Table A12 and Section C.3.
-Table 1: HEST-Benchmark results.
-Random Forest regression evaluation (70 trees) on publicly available state-of-the-art patch encoders.
+Table 1: Comparison of 11 patch encoders evaluated on the HEST-Benchmark.
+Reported results are based on PCA with 256 factors followed by a ridge regression.
 Model performance measured with Pearson correlation.
-We report the mean $\pm$ standard deviation computed over all folds (or patients).
+Standard deviation is reported across all folds (i.e., patients).
 Best is bold, second best is underlined.
-Models sorted by publication date.
-ResNet50 KimiaNet Ciga CTransPath Remedis Phikon PLIP UNI CONCH GigaPath
-IDC 0.440 0.420 0.406 0.454 0.491 0.430 0.436 $\underline{0.502}$ $\bm{0.504}$ 0.492
-±$\pm 0.03$$\pm 0.04$$\pm 0.02$$\pm 0.04$$\pm 0.06$$\pm 0.06$$\pm 0.04$$\pm 0.06$$\pm 0.05$$\pm 0.05$
-PRAD 0.318 0.328 0.332 0.346 0.335 $\bm{0.377}$ 0.362 0.357 $\underline{0.373}$ 0.372
-±$\pm 0.02$$\pm 0.01$$\pm 0.00$$\pm 0.01$$\pm 0.00$$\pm 0.00$$\pm 0.02$$\pm 0.00$$\pm 0.03$$\pm 0.01$
-PAAD 0.389 0.410 0.397 0.406 $\bm{0.451}$ 0.372 0.392 0.424 $\underline{0.431}$ 0.425
-±$\pm 0.07$$\pm 0.08$$\pm 0.07$$\pm 0.08$$\pm 0.06$$\pm 0.06$$\pm 0.09$$\pm 0.06$$\pm 0.05$$\pm 0.07$
-SKCM 0.446 0.452 0.484 0.535 0.577 0.516 0.461 $\bm{0.613}$ $\underline{0.582}$ 0.541
-±$\pm 0.06$$\pm 0.03$$\pm 0.01$$\pm 0.06$$\pm 0.00$$\pm 0.04$$\pm 0.06$$\pm 0.02$$\pm 0.04$$\pm 0.03$
-COAD 0.107 0.080 0.102 0.123 0.125 0.137 0.112 $\bm{0.147}$ 0.124 $\underline{0.139}$
-±$\pm 0.06$$\pm 0.03$$\pm 0.04$$\pm 0.05$$\pm 0.06$$\pm 0.05$$\pm 0.04$$\pm 0.05$$\pm 0.03$$\pm 0.03$
-READ 0.051 0.038 0.046 0.083 0.099 0.138 0.063 $\bm{0.162}$ 0.132 $\underline{0.156}$
-±$\pm 0.08$$\pm 0.05$$\pm 0.09$$\pm 0.09$$\pm 0.11$$\pm 0.10$$\pm 0.10$$\pm 0.08$$\pm 0.08$$\pm 0.08$
-CCRCC 0.136 0.136 0.127 0.171 $\bm{0.200}$ 0.178 0.124 $\underline{0.186}$ 0.149 0.182
-±$\pm 0.04$$\pm 0.05$$\pm 0.04$$\pm 0.04$$\pm 0.04$$\pm 0.05$$\pm 0.04$$\pm 0.05$$\pm 0.06$$\pm 0.05$
-HCC 0.034 0.028 0.045 $\bm{0.060}$ $\underline{0.059}$ 0.041 0.038 0.051 0.040 0.055
-±$\pm 0.01$$\pm 0.01$$\pm 0.00$$\pm 0.00$$\pm 0.00$$\pm 0.02$$\pm 0.00$$\pm 0.00$$\pm 0.01$$\pm 0.01$
-LUNG 0.497 0.507 0.515 0.531 $\bm{0.573}$ 0.541 0.533 0.511 $\underline{0.569}$ 0.547
-±$\pm 0.01$$\pm 0.01$$\pm 0.02$$\pm 0.02$$\pm 0.02$$\pm 0.03$$\pm 0.02$$\pm 0.03$$\pm 0.02$$\pm 0.03$
-LYMPH_IDC 0.205 0.206 0.218 0.238 0.243 0.243 0.229 0.234 $\bm{0.249}$ $\underline{0.248}$
-±$\pm 0.05$$\pm 0.06$$\pm 0.07$$\pm 0.06$$\pm 0.06$$\pm 0.06$$\pm 0.06$$\pm 0.05$$\pm 0.06$$\pm 0.05$
-Average 0.262 0.261 0.267 0.295 0.315 0.297 0.275 $\bm{0.319}$ 0.315 $\underline{0.316}$
-We learn a regression model to map model-specific patch embeddings (512 to 2048 dimensions) to the log1p-normalized expression of the top-50 highly variable genes.
+IDC PRAD PAAD SKCM COAD READ ccRCC LUAD LYMPH IDC Average
+ResNet50 (IN) 0.4741 0.3075 0.3889 0.4822 0.2528 0.0812 0.2231 0.4917 0.2322 0.326
+$\pm$ 0.047 $\pm$ 0.0309 $\pm$ 0.0754 $\pm$ 0.1141 $\pm$ 0.0372 $\pm$ 0.0517 $\pm$ 0.0554 $\pm$ 0.0119 $\pm$ 0.0491
+CTransPath 0.511 0.3427 0.4378 0.5106 0.2285 0.11 0.2279 0.4985 0.2353 0.3447
+$\pm$ 0.0531 $\pm$ 0.0458 $\pm$ 0.0664 $\pm$ 0.0827 $\pm$ 0.0557 $\pm$ 0.0764 $\pm$ 0.0475 $\pm$ 0.0414 $\pm$ 0.0477
+Phikon 0.5327 0.342 0.4432 0.5355 0.2585 0.1517 0.2423 0.5468 0.2373 0.3656
+$\pm$ 0.0914 $\pm$ 0.0767 $\pm$ 0.0684 $\pm$ 0.0549 $\pm$ 0.0056 $\pm$ 0.0822 $\pm$ 0.0263 $\pm$ 0.0045 $\pm$ 0.0457
+CONCH 0.5363 0.3548 0.4475 0.5791 0.2533 0.1674 0.2179 0.5312 0.2507 0.3709
+$\pm$ 0.0842 $\pm$ 0.0099 $\pm$ 0.0729 $\pm$ 0.0542 $\pm$ 0.0075 $\pm$ 0.0476 $\pm$ 0.0353 $\pm$ 0.0107 $\pm$ 0.042
+REMEDIS 0.529 0.3471 0.4644 0.5818 0.2856 0.1145 0.2647 0.5336 0.2473 0.3742
+$\pm$ 0.069 $\pm$ 0.0074 $\pm$ 0.0722 $\pm$ 0.0421 $\pm$ 0.02 $\pm$ 0.0987 $\pm$ 0.0539 $\pm$ 0.0326 $\pm$ 0.0585
+GigaPath 0.5508 0.3708 0.4768 0.5538 0.301 0.186 0.2391 0.5399 0.2493 0.3853
+$\pm$ 0.0726 $\pm$ 0.021 $\pm$ 0.0489 $\pm$ 0.0586 $\pm$ 0.0145 $\pm$ 0.0704 $\pm$ 0.0364 $\pm$ 0.0369 $\pm$ 0.0522
+UNI 0.5702 0.314 0.4764 0.6254 0.263 0.1762 0.2427 0.5511 0.2565 0.3862
+$\pm$ 0.0833 $\pm$ 0.0715 $\pm$ 0.0687 $\pm$ 0.0338 $\pm$ 0.0311 $\pm$ 0.0565 $\pm$ 0.0368 $\pm$ 0.0198 $\pm$ 0.0436
+Virchow 0.5702 0.3309 0.4875 0.6088 0.311 0.2019 0.2637 0.5459 0.2594 0.3977
+$\pm$ 0.0939 $\pm$ 0.0081 $\pm$ 0.0412 $\pm$ 0.0733 $\pm$ 0.0083 $\pm$ 0.0467 $\pm$ 0.039 $\pm$ 0.0262 $\pm$ 0.043
+Virchow2 0.5922 0.3465 0.4661 0.6174 0.2578 0.2084 0.2788 0.5605 0.2582 0.3984
+$\pm$ 0.0814 $\pm$ 0.029 $\pm$ 0.0676 $\pm$ 0.0174 $\pm$ 0.0189 $\pm$ 0.0502 $\pm$ 0.0516 $\pm$ 0.0172 $\pm$ 0.0296
+UNIv1.5 0.5989 0.3645 0.4902 0.6401 0.2925 0.2240 0.2522 0.5586 0.2597 0.4090
+$\pm$ 0.0842 $\pm$ 0.0308 $\pm$ 0.0502 $\pm$ 0.041 $\pm$ 0.0142 $\pm$ 0.0378 $\pm$ 0.04 $\pm$ 0.026 $\pm$ 0.0424
+H-Optimus-0 0.5982 0.385 0.4932 0.6432 0.2991 0.2292 0.2654 0.5582 0.2595 0.4146
+$\pm$ 0.0843 $\pm$ 0.0008 $\pm$ 0.0443 $\pm$ 0.0668 $\pm$ 0.0007 $\pm$ 0.041 $\pm$ 0.0309 $\pm$ 0.0324 $\pm$ 0.04
+We learn a regression model to map model-specific patch embeddings (512 to 2,048 dimensions) to the log1p-normalized expression of the top 50 highly variable genes.
 All tasks are evaluated using the Pearson correlation between the predicted and measured gene expression.
@@
 All experiments were run on a single NVIDIA 3090 GPU.
-Table 1 summarizes HEST-Benchmark when employing a Random Forest regression model (70 trees), and Table A13 when employing a Ridge regression model.
-When using Random Forest, UNI Chen et al.
-(2024), GigaPath Xu et al.
-(2024), Remedis Azizi et al.
-(2023), and CONCH Lu et al.
-(2024) provide similar performance (0.315 to 0.319 average Pearson correlation across all tasks), outperforming all other baselines.
-When using a Ridge regression, CONCH Lu et al.
-(2024) significantly outperforms all models with an average of +5.0% absolute performance gain over the second-best model, Phikon, and 5.3% better than the third-best model, CTransPath with an average Pearson correlation of 0.316.
-Supervised vs.
-contrastive vs.
-student-teacher pretraining: When employing both Random Forest or Ridge regression, supervised models (ResNet50-IN and KimiaNet) exhibit lower performance than other models.
-This can be attributed to (1) a lack of sufficiently large and diverse pretraining datasets and models or (2) fundamental limitations of supervised models in transferring to challenging tasks.
-Using Random Forest, the best student-teacher models (UNI, CONCH, GigaPath) provide similar performance to the best contrastive approach (Remedis).
-Vision Transformer Base vs.
-Large vs.
-Giant: CONCH (ViT-Base, 86M parameters), UNI (ViT-Large, 307M) and GigaPath (ViT-Giant, 1.13B) provide similar performance despite GigaPath using 13$\times$ more parameters than CONCH.
-UNI and GigaPath have been trained in a unimodal fashion by building “views” using traditional augmentation techniques, such as color jittering and cropping.
-In contrast, CONCH was trained in two steps: first using unimodal iBOT pretraining, and then by fine-tuning the resulting model on image-caption pairs.
-Such fine-tuning step could explain its parameter-efficiency and high performance in HEST-Benchmark.
-Random Forest vs.
-Ridge regression: The best-performing model using Random Forest (UNI, 0.319) is similar to the best-performing model using Ridge regression (CONCH, 0.316), showing the predictive capabilities of both regression models.
-However, Ridge regression leads to poor performance using larger models (GigaPath, Remedis, and UNI), which could be explained by the curse of dimensionality (CONCH embedding dimension is 3$\times$ smaller than GigaPath).
-We also emphasize that the performance could be optimized using more complex models (such as MLP) or dedicated hyper-parameter search.
-Overall, HEST-Benchmark brings new insights into the performance of patch encoders for histology.
-We observe that (1) Random Forest regression leads to multiple models performing similarly (CONCH, UNI, GigaPath) which may suggest a form of plateauing of patch encoders, (2) CONCH is the only model that shows high performance using both Ridge and Random Forest regression models, which could be attributed to its compact size (ViT-B) and visual-language fine-tuning step, and (3) performance remains low for some tasks (e.g., READ and HCC) which implies that for certain cancer types, the morphology might not be as reflective of gene expression, as well as leaving room for follow-up works.
-6 HEST for discovery
-Figure 2: HEST for biomarker discovery: Analysis of an invasive ductal carcinoma Xenium sample.
+We report performance using three downstream regression models: (i) PCA-reduced embeddings (with n=256 factors) followed by Ridge regression trained with adaptive regularization as shown in Table 1, (ii) Ridge regression model as shown in Appendix Table A13, and (iii) an XGBoost regression model with 100 estimators and a maximum depth of 3 as shown in Appendix Table A14.
+Our main results are reported using PCA+Ridge (i) and XGBoost (iii).
+Directly applying Ridge regression may unfairly penalize models with larger embedding dimensions.
+To guarantee a fairer and more objective comparison, we chose to utilize PCA reduction.
+5.3 Scaling laws in HEST-Benchmark
+Overall, H-Optimus-0 brings the best average Pearson correlation in both PCA+Ridge and XGBoost evaluation, outperforming the second-best model, UNIv1.5, by 0.56% and 0.69%, respectively.
+ResNet50 (IN), the only model that was not pretrained on histology images, leads to the lowest performance in both PCA+Ridge and XGBoost.
+Legacy domain-specific models, such as CTransPath, are outperformed by all recent models, including UNIv1.5, UNI, GigaPath, Virchow, and H-Optimus-0.
+The disparity between the top and bottom domain-specific models is notable, showing an absolute improvement of 7.0% for PCA+Ridge and 4.8% for XGBoost.
+When inspecting individual performance, we observe large differences across tasks from 0.6432 Pearson correlation in SKCM to 0.2292 in READ for H-Optimus-0 evaluated using PCA+Ridge.
+Model scaling law.
+By inspecting the number of trainable parameters within the vision encoder for each model, we can describe how model size influences performance (measured with average Pearson correlation across all tasks, Figure 2.a).
+Performance increases with model size following a logarithmic scaling law (Pearson correlation of R=0.81, P-value$<$0.01).
+Models considered “parameter-efficient” are represented on top of the log-transformed linear regression line (e.g., CONCH, UNIv1.5, and H-Optimus-0).
+This observation suggests a trade-off between downstream performance and model size.
+Despite the observation of a model scaling law, significant variations in performance among models of identical size persist, such as between H-Optimus-0 and GigaPath, both of which are ViT-giant models with over one billion parameters.
+Data scaling law.
+We further explored how the number of training samples used for pretraining each model (i.e., the number of image patches) affects performance.
+We observe that increasing the number of patches moderately correlates with the average performance (Pearson correlation of R=0.48, P-value=0.13).
+This correlation is weaker than model size, which we hypothesize is due to this analysis overlooking both the absolute number of WSIs used for pretraining (image patches are not independently and identically distributed per WSI) and disparities of morphological variety among WSIs (e.g., staining variation, disease diversity, artifacts, etc.).
+Overall, HEST-Benchmark brings new insights into the performance of foundation models for pathology.
+We observe that (1) Scaling the model size strongly correlates with average performance, but the gains grow logarithmically with the number of trainable parameters.
+(2) Scaling the number of training patches weakly correlates with a higher performance (also on a logarithmic scale).
+(3) Performance remains low for some tasks (e.g., READ and ccRCC), which suggests that (i) the morphology might not be as reflective of gene expression for some cancer types or (ii) some cohorts have more noise than others (e.g., due to batch effects, low sensitivity, dropout events, or spillover between adjacent spots).
+Figure 2: Scaling laws in HEST-Benchmark.
 a.
-IDC Xenium samples with neoplastic nuclei overlaid in red ($n_{c}=168,033$ detected nuclei).
+Model scaling law comparing the number of training parameters in the vision encoder (log-scale) and the average performance on the HEST-Benchmark.
+Pearson correlation between parameters and performance of R=0.81 (P-value < 0.01).
+b.
+Data scaling law comparing the number of image patches used for pretraining (log-scale) and the average performance on the HEST-Benchmark.
+Pearson correlation between number of patches and performance of R=0.48 (P-value=0.13).
+6 HEST for biomarker exploration
+HEST-1k also enables the analysis of interactions and correlations between tissue morphology (as seen in H&E) and local gene expression (as provided in ST).
+Here, we showcase the capabilities of HEST-1k (1) by studying morphological correlates of expression changes in invasive breast cancer and (2) by visualizing tumor heterogeneity both on the morphological and molecular sides.
+Specifically, we focus on invasive ductal carcinoma (IDC) samples imaged with Xenium.
+Using CellViT nuclear segmentation and classification, we identified neoplastic nuclei (exemplified in two samples: Figure 3.a with n=168,033 nuclei and Appendix Figure 6.a with n=342,018 nuclei).
+We then overlay the WSI with the expression of specific genes, such as GATA3, a known prognostic gene in breast cancer [mehra2005identification](Figure 3.b).
+This qualitatively shows that high GATA3 expression is associated with cancerous regions and reveals heterogeneity within invasive regions (e.g., the right-most region shows higher expression of GATA3 than the rest of the tumor, Figure 3.b).
+Using the nuclear segmentation, we can compute human-interpretable features related to nuclear size (area, perimeter, major axis length, minor axis length, and equivalent diameter), topology and shape (roundness, ellipticity, eccentricity, extent, and roughness), and cell distribution (cell density and crowdedness).
+A heatmap of the nuclear area of neoplastic cells also indicates morphological heterogeneity among neoplastic regions (Figure 3.c,d).
+Regions with a high nuclear area and elevated GATA3 expression notably overlap, suggesting that this tumor exhibits molecular heterogeneity, which to some degree is morphologically expressed.
+To investigate this hypothesis, we measured the Pearson correlation between the expression of GATA3 and nuclear area in neoplastic cells (Figure 3.e).
+We observe a moderate correlation (R=0.47, P-value$<10^{-4}$), which is also observed in other genes and morphological features (Figure 3.f).
+Overall, out of the 12 human-interpretable features we analyzed, we found the highest association with gene expression for size-related features, while features involving topology, shape, and cell distribution had a lower correlation (R<0.2).
+A similar analysis in another IDC sample (Appendix Figure 6.b,c) further asserted these observations.
+In particular, we found the highest associations between nuclear size and expression for the genes FLNB (R=0.45, P-value$<10^{-4}$) and TPD52 (R=0.47, P-value$<10^{-4}$), both involved in breast tumor growth and proliferation [bandaru2014targeting, ren2021tumor], and FOXA1 (R=0.47, P-value$<10^{-4}$), a known prognostic factor associated with better survival [badve2007foxa1, wolf2007foxa1].
+Such analysis highlights how HEST-1k can be used to identify fine-grained morphological correlates of expression.
+Similar approaches can be used to characterize morphological and molecular tumor heterogeneity at a larger scale.
+Figure 3: HEST for biomarker exploration: Analysis of an invasive ductal carcinoma sample imaged with Xenium.
+a.
+IDC Xenium sample with neoplastic nuclei overlaid in red ($n_{c}=168,033$ detected nuclei).
 Gray scale bar represents 2 mm.
@@
 e., f.
-Correlation between area and GATA3, and minor axis length and MYBPC1.
-HEST-1k also enables the analysis of interactions and correlations between tissue morphology (as seen in H&E) and local gene expression (as provided in ST).
-Here, we showcase the capabilities of HEST-1k (1) by studying morphological correlates of expression changes in invasive breast cancer and (2) by visualizing tumor heterogeneity both on the morphological and molecular sides.
-Specifically, we focus on invasive ductal carcinoma (IDC) samples imaged with Xenium.
-Using CellViT nuclear segmentation and classification, we identified neoplastic nuclei (exemplified in two samples: Figure 2.a with n=168,033 nuclei and Appendix Figure 5.a with n=342,018 nuclei).
-We then overlay the WSI with the expression of specific genes, such as GATA3, a known prognostic gene in breast cancer Mehra et al.
-(2005)(Figure 2.b).
-This qualitatively shows that high GATA3 expression is associated with cancerous regions and reveals heterogeneity within invasive regions (e.g., the right-most region shows higher expression of GATA3 than the rest of the tumor, Figure 2.b).
-Using the nuclear segmentation, we can compute human-interpretable features related to nuclear size (area, perimeter, major axis length, minor axis length, and equivalent diameter), topology and shape (roundness, ellipticity, eccentricity, extent, and roughness), and cell distribution (cell density and crowdedness).
-A heatmap of the nuclear area of neoplastic cells also indicates morphological heterogeneity among neoplastic regions (Figure 2.c,d).
-Regions with a high nuclear area and elevated GATA3 expression notably overlap, suggesting that this tumor exhibits molecular heterogeneity, which to some degree is morphologically expressed.
-To investigate this hypothesis, we measured the Pearson correlation between the expression of GATA3 and nuclear area in neoplastic cells (Figure 2.e).
-We observe a significant correlation (R=0.47), which is also observed in other genes and morphological features (Figure 2.f).
-Overall, out of the 12 human-interpretable features we analyzed, we found the highest association with gene expression for size-related features, while features involving topology, shape, and cell distribution had a lower correlation (R<0.2).
-A similar analysis in another IDC sample (Appendix Figure 5.b,c) confirmed these observations.
-In particular, we found the highest associations between nuclear size and expression for the genes FLNB (R=0.49) and TPD52 (R=0.49), both involved in breast tumor growth and proliferation Bandaru et al.
-(2014); Ren et al.
-(2021), and FOXA1 (R=0.47), a known prognostic factor associated with better survival Badve et al.
-(2007); Wolf et al.
-(2007).
-Such analysis highlights how HEST-1k can be used for identifying fine-grained morphological correlates of expression.
-Additional analyses of morphological and molecular tumor heterogeneity can be leveraged at larger scale using similar approaches.
+Correlation between nuclear area and GATA3, and minor axis length and MYBPC1.
 7 HEST for multimodal representation learning
@@
 Several problem statements can be explored, such as cross-modal alignment and retrieval, multimodal fusion, etc.
-Here, we fine-tune CONCH (strong performer on HEST-Benchmark) on five Xenium invasive breast cancer cases (four ductal and one lobular case) using contrastive alignment.
+Here, we fine-tune CONCH [lu2024towards] (ViT-Base model) on five Xenium invasive breast cancer cases (four ductal and one lobular case) using multimodal contrastive alignment.
 We hypothesize that the resulting breast cancer-specific patch encoder, termed CONCH-FT, can better encode the underlying molecular landscape associated with disease-specific morphologies.
@@
 […]
-We evaluate the resulting CONCH-FT model to predict ER, PR, and HER2 expression status (binary) from WSIs in the BCNB dataset Xu et al.
-(2021) (n=1,058 WSIs).
-To generate a slide representation for a WSI, we take the average of the patch embedding in the WSI, which is subsequently mapped to the expression status using logistic regression (Table 2).
+We evaluate the resulting CONCH-FT model to predict ER, PR, and HER2 expression status (binary) from WSIs in the BCNB dataset [xu2021predicting] (n=1,058 WSIs).
+To generate a slide representation for a WSI, we take the average of the patch embedding in the WSI (mean pooling), which is subsequently mapped to the expression status using logistic regression (Table 2).
+The simple mean pooling approach to embedding the slide without additional fine-tuning on the downstream tasks highlights the expressivity of the learned latent space.
 We observe that CONCH-FT outperforms CONCH on most metrics, demonstrating that pan-tissue histology patch encoders can be further fine-tuned to obtain better tissue-specific patch encoders.
-This is further validated by the larger rank induced by the patch embedding space Garrido et al.
-(2023) for CONCH-FT, suggesting better expressivity of the patch embeddings.
-While these results are based on only five WSIs, we anticipate additional benefits when training with larger disease-specific cohorts.
+This is further validated by the larger rank induced by the patch embedding space [garrido2023rankme] for CONCH-FT, suggesting better expressivity of the patch embeddings.
+While these results are based on only five paired WSIs, we anticipate additional benefits when training with larger disease-specific cohorts.
 Table 2: CONCH fine-tuning on invasive breast cancer.
-Logistic regression evaluation for ER/PR/HER2 expression status on BCNB (n=1,058 WSIs).
+Logistic regression evaluation for ER/PR/HER2 status on BCNB (binary task, n=1,058 WSIs).
 A WSI is represented by the average of the patch embeddings within each WSI.
@@
 We assembled HEST-1k, a dataset comprising paired spatial transcriptomics, H&E-stained whole-slide images, and comprehensive metadata built from public and internal cohorts.
-HEST-1k includes 1,108 samples, encompassing 1.5 million spots and over 60 million cells.
-The scale and comprehensiveness of HEST-1k, supported by the HEST-Library, enable exploring directions such as biomarker discovery and multimodal representation learning.
-Additionally, motivated by the need for new diverse and challenging patch-level benchmarks, we curated the HEST-Benchmark, a set of ten tasks covering nine cancer types and ten organs for gene expression prediction from histology.
-HEST-Benchmark revealed new insights into the predictive capabilities of foundation models for histology, such as the added benefits of vision-language alignment Lu et al.
-(2024).
+HEST-1k includes 1,229 samples, encompassing 2.1 million spots and over 76 million cells.
+The scale and comprehensiveness of HEST-1k, supported by the HEST-Library, enable exploring directions such as biomarker exploration and multimodal representation learning.
+Additionally, motivated by the need for new, diverse, and challenging patch-level benchmarks, we curated the HEST-Benchmark, a set of nine tasks covering eight cancer types and nine organs for gene expression prediction from histology.
+The HEST-Benchmark revealed data and model scaling laws across 11 foundation models of different dimensions and pretraining scale [lu2024towards].
 Limitations.
-Our study includes limitations.
+Our study includes a few limitations.
 First, research data, such as those generated in spatial transcriptomic, are inherently noisy.
-While we tried to minimize “label” noise (e.g., by re-estimating image magnification and alignment, and unifying cancer samples using oncotree code taxonomy), staining and compression artifacts, varying acquisition protocols, among others can negatively impact the quality of HEST-1k.
+While we tried to minimize “label” noise (e.g., by re-estimating image magnification and alignment, and unifying cancer samples using oncotree code taxonomy), staining and compression artifacts, varying acquisition protocols, among others, can negatively impact the quality of HEST-1k.
 Second, batch effects (on both the imaging and transcriptomic sides) can be significant across samples, datasets, and technologies.
-This study does not quantify these batch effects or explore methods to mitigate them.
+While this study does not explore batch effects quantification or mitigation, we provide a set of helpers in HEST-Library to let users explore this direction.
 Lastly, although the HEST-Library was designed for versatility, it cannot cover all existing formats and should rather be viewed as a blueprint for processing ST data in a consistent and unified manner.
 Future work.
-Spatial transcriptomics is rapidly evolving with new datasets frequently published.
+Spatial transcriptomics is rapidly evolving, with new datasets frequently published.
 As they become available, we will keep updating HEST-1k with new resources.
-This study merely starts to uncover the potential of HEST-1k for advancing translational research and biomarker discovery, and we plan to explore these capabilities further.
+This study merely starts to uncover the potential of HEST-1k for advancing translational research and biomarker exploration, and we plan to explore these capabilities further.
 Additionally, the prospects for multimodal representation learning with HEST-1k are promising and are expected to grow with the addition of more data.
@@
 We thank Dr.
-Maxime Meylan for his insights and guidance on accessing data published in Meylan et al.
-(2022).
+Maxime Meylan for his insights and guidance on accessing data published in [meylan2022tertiary].
+We thank Rushin Gindra for his support in inspecting HEST-1k data, reporting issues, and providing references.
+HEST is supported by the Brigham and Women’s Hospital (BWH) President’s Fund, Mass General Hospital (MGH) Pathology, and the National Institute of Health (NIH) National Institute of General Medical Sciences (NIGMS) through R35GM138216.
 S.J.W.
-was supported by the Helmholtz Association under the joint research school “Munich School for Data Science - MUDS” and the Add-on Fellowship of the Joachim Herz Foundation.
+is supported by the Helmholtz Association under the joint research school “Munich School for Data Science - MUDS” and the Add-on Fellowship of the Joachim Herz Foundation.
 Checklist
   1. 1.
-    Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-[Yes] Each claim: HEST-1k, HEST-Library, HEST-Benchmark, HEST for discovery and multimodal fine-tuning are supported by dedicated sections in the main text, in addition to supplementary information provided in the appendix.
+    Do the main claims made in the abstract, and introduction accurately reflect the paper’s contributions and scope?
+[Yes] Each claim: HEST-1k, HEST-Library, HEST-Benchmark, HEST for biomarker exploration, and multimodal fine-tuning are supported by dedicated sections in the main text, in addition to supplementary information provided in the appendix.
 In addition, all the code to reproduce results is made available.
@@
     Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)?
-[Yes] In the abstract, we provide a link to access the HEST website and the HEST-Library on GitHub.
-HEST-Library includes a link to download all data to run the HEST-Benchmark.
+[Yes] In the abstract, we provide a link to access the HEST page on GitHub.
+HEST-Library includes a link to download all data and to run the HEST-Benchmark.
 Finally, we provide all metadata associated with HEST-1k in a CSV as part of the supplementary material.
-Processed ST and WSIs will be made available upon acceptance.
   6. 6.
@@
 [Yes] Metadata associated with HEST-1k includes the license under which data were originally published.
-We ensured that the reported license allowed for distributing and creating derivatives of the data.
+We ensured that the reported license allowed the distribution and creation of derivatives of the data.
   11. 11.
@@
 Consequently, this model should not be regarded as the definitive standard, and users should exercise particular caution when utilizing this part of the dataset.
-Despite our efforts to exclude sensitive information, such as patient names, addresses, and social security numbers, users are expressly prohibited from attempting to reverse engineer the data to extract any confidential patient information.
-On the presumption that users will adhere to the aforementioned restrictions, we have not identified any potential adverse societal impacts that could arise from using HEST-1k.
-The dataset is hosted on HuggingFace Dataset webpage.
+Despite our efforts to exclude sensitive information, such as patient names, addresses, and social security numbers, users are expressly prohibited from attempting to reverse engineer the data to extract confidential patient information.
+In the presumption that users will adhere to the aforementioned restrictions, we have not identified any potential adverse social impacts that could arise from the use of HEST-1k.
+The dataset is hosted on the HuggingFace Dataset webpage.
 All instructions are provided in the main README of HEST-Library.
-From there, users can choose to download HEST-1k in its entirety, or a subset (e.g., only breast cancer samples).
-HEST-1k, HEST-Benchmark, and HEST library are released under the Attribution-NonCommercial-ShareAlike 4.0 International license (CC BY-NC-SA 4.0 Deed) [^9] https://creativecommons.org/licenses/by-nc-sa/4.0/.
+From there, users can choose to download HEST-1k in its entirety or a subset (e.g., only breast cancer samples).
+The HEST-1k, HEST-Benchmark, and HEST library are released under the Attribution-NonCommercial-ShareAlike 4.0 International license (CC BY-NC-SA 4.0 Deed) [^11] https://creativecommons.org/licenses/by-nc-sa/4.0/.
 Appendix B Background
-This study connects two fields: (1) computational pathology, which primarily uses routinely acquired clinical data to determine outcomes such as disease diagnosis from H&E-stained digitized tissue sections, and (2) spatial transcriptomics, which has, so far, been confined to biological research and aims to identify new biomarkers predictive of disease progression or treatment response, among others.
+This study connects two fields: (1) computational pathology, which primarily uses routinely acquired clinical data to determine outcomes such as disease diagnosis from H&E-stained digitized tissue sections, and (2) spatial transcriptomics, which so far has been confined to biological research and aims to identify new biomarkers predictive of disease progression or response to treatment, among others.
 B.1 Computational pathology
 […]
-Such frameworks have been shown to achieve better or similar performance than humans for Gleason grading in prostate cancerBulten et al.
-(2022), metastasis detection in lymph nodesBejnordi et al.
-(2017), determining the origin of a cancer of unknown primaryLu et al.
-(2021a), predicting heart transplant rejectionLipkova et al.
-(2022), among others.
+Such frameworks have been shown to achieve better or similar performance than humans for Gleason grading in prostate cancer[bulten2022artificial], metastasis detection in lymph nodes[bejnordi2017diagnostic], determining the origin of a cancer of unknown primary[lu2021ai], predicting heart transplant rejection[lipkova2022deep], among others.
 B.2 Spatial transcriptomics (ST)
-ST enables the measurement of gene activity and the mapping to its corresponding location on the tissue.
-In this study, we gather samples from two ST paradigms: sequencing-based (ST, Visium, Visium HD) and imaging-based (Xenium).
+ST enables the measurement of gene activity and the mapping of its corresponding location in the tissue.
+In this study, we collected samples from two ST paradigms: sequencing-based (ST, Visium, Visium HD) and imaging-based (Xenium).
 […]
-Xenium: Xenium is an imaging-based spatial-profiling technology offering higher resolution compared to its sequencing counterpart, Visium.
-Xenium imaging-based paradigm breaks free from the resolution limitations of its sequencing-based counterpart, it directly performs in situ RNA capturing on tissue sections by imaging fluorescent RNA markers derived from padlock probes and rolling circle amplification chemistry, hence offering sub-cellular expression resolution.
-As of 2024, Xenium cannot perform whole transcriptome measurement and is limited to gene panels of up to 580 genes.
+Xenium: Xenium is an imaging-based spatial profiling technology that offers in situ RNA capturing on tissue sections by imaging fluorescent RNA markers derived from padlock probes and rolling circle amplification chemistry.
+This approach provides the exact 2D location of each measured transcript.
+As of 2024, Xenium cannot perform whole transcriptome measurements and is limited to gene panels of up to 5,000 genes.
 Appendix C HEST
@@
 We provide a comprehensive description of all publicly available and internal cohorts integrated into HEST-1k.
-Table A1: HEST-1k data overview. All included samples include a license that allows sharing and redistributing. National Center for Biotechnology Information: NCBI.
+Table A1: HEST-1k data overview. All samples include a license that allows sharing and redistributing. National Center for Biotechnology Information: NCBI.
 Resource Number of datasets Number of samples Size (GB, raw)
-10x Genomics 79 97 225
+10x Genomics 87 112 275
 Mendeley 9 118 181
 Spatial-Research 4 139 18
-Zenodo 3 17 11
-NCBI 42 677 265
+Zenodo 4 21 18
+NCBI 43 696 298
 Internal 3 28 60
-Miscellaneous 2 32 65
+Miscellaneous 4 114 147
 Table A2: Datasets gathered from 10x Genomics portal. $n$: number of samples in the cohort.
@@
 Adult Mouse Olfactory Bulb Brain Visium 1 32,285
+Characterization of immune cell populations in the tumor microenvironment of colorectal cancer using high definition spatial profiling [oliveira2024characterization] Bowel Mixed 8 18,085
 FFPE Human Breast using the Entire Sample Area Breast Xenium 2 541
@@
 FFPE Human Pancreas with Xenium Multimodal Cell Segmentation Pancreas Xenium 1 541
+FFPE Human Prostate Adenocarcinoma with 5K Human Pan Tissue and Pathways Panel Prostate Xenium 1 10,006
+FFPE Human Skin Primary Dermal Melanoma with 5K Human Pan Tissue and Pathways Panel Skin Xenium 1 10,017
 Fresh Frozen Mouse Colon with Xenium Multimodal Cell Segmentation Bowel Xenium 1 541
+Fresh Frozen Mouse Brain Hemisphere with 5K Mouse Pan Tissue and Pathways Panel Brain Xenium 1 13,780
 Fresh Frozen Visium on CytAssist: Human Breast Cancer, Probe-Based Whole Transcriptome Profiling Breast Visium 1 18,085
@@
 Mouse Tissue Microarray in 5x5 Layout with 1 mm Edge to Edge Spacing (FFPE) Kidney/Brain Visium 1 19,465
-Normal Human Prostate (FFPE) Prostate Visium 1 17943
+Normal Human Prostate (FFPE) Prostate Visium 1 17,943
 Pancreatic Cancer with Xenium Human Multi-Tissue and Cancer Panel Pancreas Xenium 1 538
@@
 Preview Data: FFPE Human Lung Cancer with Xenium Multimodal Cell Segmentation Lung Xenium 1 541
+Preview Data: FFPE Human Lymph Node with 5K Pan Tissue and Pathways Panel Lymphoid Xenium 1 11,094
 Visium CytAssist Gene Expression Libraries of Post-Xenium Human Colon Cancer (FFPE) Bowel Visium 4 18,085
@@
 Visium CytAssist, Mouse Embryo, 11 mm Capture Area (FFPE) Embryo Visium 1 19,465
+Visium HD Spatial Gene Expression Library, Mouse Kidney (FFPE) Kidney Visium HD 1 19,059
+Visium HD Spatial Gene Expression Library, Mouse Embryo (FFPE) Embryo Visium HD 1 19,059
+Visium HD Spatial Gene Expression Library, Human Pancreas (FFPE) Pancreas Visium HD 1 18,085
 Whole Mouse Pup Preview Data (Xenium Mouse Tissue Atlassing Panel) Whole organism Xenium 1 541
@@
 Publication Organ Technology $\bm{n}$ Num. genes
-10X Visium Spatial transcriptomics of murine colon at d14 (mucosa healing) in B cell sufficient-deficient mice Parigi et al. (2022) Bowel Visium 2 31,053
-10X Visium Spatial transcriptomics of murine colon in steady state and during recovery after DSS colitis Parigi et al. (2022) Bowel Visium 2 31,053
-A Spatial Transcriptomic atlas of the human kidney papilla identifies significant immune injury in patients with stone disease Canela et al. (2023) Kidney Visium 7 36,601
-A cellular hierarchy in melanoma uncouples growth and metastasis Karras et al. (2022) Skin Visium 3 31,053
-A new epithelial cell subpopulation predicts response to surgery, chemotherapy, and immunotherapy in bladder cancer Abdel-Hafiz et al. (2023); Gouin et al. (2021) Bladder Visium 4 33,538
-A novel model of binge ethanol exposure reveals enhanced neurodegeneration with advanced age Anton et al. (2024) Brain Visium 4 32,285
-A single-cell transcriptomic analysis of endometriosis Fonseca et al. (2023) Uterus Visium 2 36,601
-Distinct mesenchymal cell states mediate prostate cancer progression Pakula et al. (2024) Prostate Visium 2 32,589
-Epithelial Plasticity and Innate Immune Activation Promote Lung Tissue Remodeling following Respiratory Viral Infection Beppu et al. (2023) Lung Visium 1 32,285
-Gene expression within a human choroidal neovascular membrane using spatial transcriptomics Voigt et al. (2023) Eye Visium 5 36,601
-Genome-wide Spatial Expression Profiling in Formalin-fixed Tissues Gracia Villacampa et al. (2021) Kidney Visium 14 33,538
-High resolution mapping of the tumor microenvironment using integrated single-cell, spatial and in situ analysis Janesick et al. (2023) Breast Xenium 4 541
-Identification of TREM1+CD163+ myeloid cells as a deleterious immune subset in HCC Giraud et al. (2022) Liver Visium 2 36,601
-Integration of spatial and single cell transcriptomics localizes epithelial-immune cross-talk in kidney injury Ferreira et al. (2021) Kidney Visium 4 33,538
-Molecular Atlas of the Adult Mouse Brain Ortiz et al. (2020) Brain Spatial Transcriptomics 75 23,371
+10X Visium Spatial transcriptomics of murine colon at d14 (mucosa healing) in B cell sufficient-deficient mice [parigi2022spatial] Bowel Visium 2 31,053
+10X Visium Spatial transcriptomics of murine colon in steady state and during recovery after DSS colitis [parigi2022spatial] Bowel Visium 2 31,053
+A Spatial Transcriptomic atlas of the human kidney papilla identifies significant immune injury in patients with stone disease [canela2023spatially] Kidney Visium 7 36,601
+A cellular hierarchy in melanoma uncouples growth and metastasis [karras2022cellular] Skin Visium 3 31,053
+A new epithelial cell subpopulation predicts response to surgery, chemotherapy, and immunotherapy in bladder cancer [abdelhafiz2023ychromosome, gouin2021ncadherin] Bladder Visium 4 33,538
+A novel model of binge ethanol exposure reveals enhanced neurodegeneration with advanced age [anton2024binge] Brain Visium 4 32,285
+A single-cell transcriptomic analysis of endometriosis [fonseca2023single] Uterus Visium 2 36,601
+Distinct mesenchymal cell states mediate prostate cancer progression [pakula2024distinct] Prostate Visium 2 32,589
+Epithelial Plasticity and Innate Immune Activation Promote Lung Tissue Remodeling following Respiratory Viral Infection [beppu2023epithelial] Lung Visium 1 32,285
+Image-based spatial transcriptomics identifies molecular niche dysregulation associated with distal lung remodeling in pulmonary fibrosis [Vannan2023.12.15.571954] Lung Xenium 20 17,145
+Gene expression within a human choroidal neovascular membrane using spatial transcriptomics [voigt2023gene] Eye Visium 5 36,601
+Genome-wide Spatial Expression Profiling in Formalin-fixed Tissues [graciaVillacampa2021genome] Kidney Visium 14 33,538
+High-resolution mapping of the tumor microenvironment using integrated single-cell, spatial, and in situ analysis [janesick2023high] Breast Xenium 4 541
+Identification of TREM1+CD163+ myeloid cells as a deleterious immune subset in HCC [giraud2022trem] Liver Visium 2 36,601
+Integration of spatial and single cell transcriptomics localizes epithelial-immune cross-talk in kidney injury [ferreira2021integration] Kidney Visium 4 33,538
+Molecular Atlas of the Adult Mouse Brain [ortiz2020molecular] Brain Spatial Transcriptomics 75 23,371
 Regional differential gene expression analyses of brains from four 24w-old Nf1+- mice Brain Visium 4 32,285
-SARS-CoV-2 Niches in Human Placenta Revealed by Spatial Transcriptomics Barrozo et al. (2023) Uterus Visium 16 36,612
-Schwann Cells Shape Tumor Cells and Cancer-Associated Fibroblasts in the Pancreatic Ductal Adenocarcinoma Microenvironment Xue et al. (2023) Pancreas Visium 4 20,615
-Single Cell and Spatial Analysis of Human Squamous Cell Carcinoma Ji et al. (2020) Skin Spatial Transcriptomics 12 17,138
-Single cell profiling of primary and paired metastatic lymph node tumors in breast cancer patients Liu et al. (2022) Lymph node Visium 4 33,931
+SARS-CoV-2 Niches in Human Placenta Revealed by Spatial Transcriptomics [barrozo2023sars] Uterus Visium 16 36,612
+Schwann Cells Shape Tumor Cells and Cancer-Associated Fibroblasts in the Pancreatic Ductal Adenocarcinoma Microenvironment [xue2023schwann] Pancreas Visium 4 20,615
+Single Cell and Spatial Analysis of Human Squamous Cell Carcinoma [ji2020multimodal] Skin Spatial Transcriptomics 12 17,138
+Single-cell profiling of primary and paired metastatic lymph node tumors in breast cancer patients [liu2022single] Lymph node Visium 4 33,931
 Table A5: Datasets gathered from NCBI. Continuation.
 Publication Organ Technology $\bm{n}$ Num. genes
-Single-cell and spatial transcriptomics characterisation of the immunological landscape in the healthy and PSC human liver Andrews et al. (2024) Liver Visium 4 36,601
-Single-nucleus Ribonucleic Acid-sequencing and Spatial Transcriptomics Reveal the Cardioprotection of Shexiang Baoxin Pill (MUSKARDIA) in Mice with Myocardial Ischemia-Reperfusion Injury Lin et al.
-(2023) Heart Visium 2 32,285
-Spatial Multimodal Analysis: MALDI-MSI and Spatial Transcriptomics within the same tissue section Vicari et al. (2023) Brain Visium 19 32,285
-Spatial RNA sequencing of regenerating mouse hindlimb muscle McKellar et al. (2021) Muscle Visium 3 33,217
-Spatial Total RNA-Sequencing of regenerating mouse hindlimb muscle and Type 1-Lang reovirus-infected mouse heart McKellar et al. (2023) Muscle Visium 7 55,414
-Spatial localization with Spatial Transcriptomics for an atlas of healthy and injured cell states and niches in the human kidney Lake et al. (2023) Kidney Visium 23 33,538
-Spatial sequencing of Foreign body granuloma Krausgruber et al. (2023) None Visium 1 15,524
-Spatial transcriptomics landscape of non-communicable inflammatory skin diseases Schäbitz et al. (2022) Skin Visium 59 20,613
-Spatial transcriptomics of adenoid cystic carcinoma of the lacrimal gland Moeyersoms et al. (2023) Eye Visium 1 17,943
+Single-cell and spatial transcriptomics characterization of the immunological landscape in the healthy and PSC human liver [andrews2024single] Liver Visium 4 36,601
+Single-nucleus Ribonucleic Acid-sequencing and Spatial Transcriptomics Reveal the Cardioprotection of Shexiang Baoxin Pill (MUSKARDIA) in Mice with Myocardial Ischemia-Reperfusion Injury [lin2023single] Heart Visium 2 32,285
+Spatial Multimodal Analysis: MALDI-MSI and Spatial Transcriptomics within the same tissue section [vicari2023spatial] Brain Visium 19 32,285
+Spatial RNA sequencing of regenerating mouse hindlimb muscle [mckellar2021large] Muscle Visium 3 33,217
+Spatial Total RNA-Sequencing of regenerating mouse hindlimb muscle and Type 1-Lang reovirus-infected mouse heart [mcKellar2023spatial] Muscle Visium 7 55,414
+Spatial localization with Spatial Transcriptomics for an atlas of healthy and injured cell states and niches in the human kidney [lake2023an] Kidney Visium 23 33,538
+Spatial sequencing of Foreign body granuloma [krausgruber2023single] None Visium 1 15,524
+Spatial transcriptomics landscape of non-communicable inflammatory skin diseases [schabitz2022spatial] Skin Visium 59 20,613
+Spatial transcriptomics of adenoid cystic carcinoma of the lacrimal gland [moeyersoms2023spatial] Eye Visium 1 17,943
 Spatial transcriptomics of the mouse brain across three age groups Brain Visium 6 32,285
-Spatial transcriptomics reveal unnresolved wound repair as potential driver of PFA Ependymoma progression Fu et al. (2023) Brain Visium 14 36,601
-Spatiotemporal dynamics of molecular pathology in amyotrophic lateral sclerosis Silas Maniatis et al. (2019) Spinal cord Spatial Transcriptomics 302 12,572
-The neurons that restore walking after paralysis Kathe et al. (2022) Spinal cord Visium 16 22,127
-Visium spatial transcriptomics analysis of lacrimal gland during chronic inflammation progression Mauduit et al. (2022) Eye Visium 4 32,285
-YAP Drives Assembly of a Spatially Colocalized Cellular Triad Required for Heart Renewal Li et al. (2024) Heart Visium 2 32,285
-Zika virus co-opts miRNA networks to persist in placental microenvironments detected by spatial transcriptomics Barrozo et al. (2024) Placenta Visium 8 32,298
-Mouse model Heptablastoma spatial transcriptomics Pilet et al. (2023) Liver Visium 10 31,053
+Spatial transcriptomics reveal unnresolved wound repair as potential driver of PFA Ependymoma progression [fu2023spatial] Brain Visium 14 36,601
+Spatiotemporal dynamics of molecular pathology in amyotrophic lateral sclerosis [maniatis2019spatiotemporal] Spinal cord Spatial Transcriptomics 302 12,572
+The neurons that restore walking after paralysis [kathe2022neurons] Spinal cord Visium 16 22,127
+Visium spatial transcriptomics analysis of lacrimal gland during chronic inflammation progression [mauduit2022spatial] Eye Visium 4 32,285
+YAP Drives Assembly of a Spatially Colocalized Cellular Triad Required for Heart Renewal [li2024yap] Heart Visium 2 32,285
+Zika virus co-opts miRNA networks to persist in placental microenvironments detected by spatial transcriptomics [barrozo2024zika] Placenta Visium 8 32,298
+Mouse model Heptablastoma spatial transcriptomics [pilet2023preneoplastic] Liver Visium 10 31,053
 Table A6: Datasets gathered on Spatial-Research.
 Publication Organ Technology $\bm{n}$ Num. genes
-A spatiotemporal organ-wide gene expression and cell atlas of the developing human heart Asp et al. (2019) Heart Spatial Transcriptomics 19 39,739
-Integrating spatial gene expression and breast tumour morphology via deep learning He et al. (2020) Breast Spatial Transcriptomics 68 16,744
-Spatial deconvolution of HER2-positive breast cancer delineates tumor-associated cell type interactions Andersson et al. (2021) Breast Spatial Transcriptomics 36 15,045
-Visualization and analysis of gene expression in tissue sections by spatial transcriptomics Ståhl et al. (2016a) Brain Spatial Transcriptomics 16 16,573
+A spatiotemporal organ-wide gene expression and cell atlas of the developing human heart [asp2019spatiotemporal] Heart Spatial Transcriptomics 19 39,739
+Integrating spatial gene expression and breast tumour morphology via deep learning [he2020integrating] Breast Spatial Transcriptomics 68 16,744
+Spatial deconvolution of HER2-positive breast cancer delineates tumor-associated cell type interactions [andersson2021spatial] Breast Spatial Transcriptomics 36 15,045
+Visualization and analysis of gene expression in tissue sections by spatial transcriptomics [stahl2016visualization] Brain Spatial Transcriptomics 16 16,573
 Table A7: Datasets gathered on Mendeley.
 Publication Organ Technology $\bm{n}$ Num. genes
-Ex-ST Andrusivova and Fan (2023) Brain Visium 5 31,053
-Genome-wide spatial expression profiling in formalin-fixed tissues Villacampa et al. (2021) Brain Visium 15 31,053
-Human ileum, Visium Mirzazadeh et al. (2021) Bowel Visium 4 33,538
-Human squamous cell carcinoma Abalo et al. (2021) Skin Visium 4 33,538
-Prostate needle biopsies pre- and post-ADT: Count matrices, histological-, and Androgen receptor immunohistochemistry images Marklund (2022) Prostate Spatial Transcriptomics 24 26,437
-Spatially resolved clonal copy number alterations in benign and malignant tissue Erickson et al. (2022) Prostate Visium 23 33,538
-Spatially resolved transcriptomic profiling of degraded Mirzazadeh et al. (2023) Bowel Visium 35 17,943
+Ex-ST [andrusivova2023exst] Brain Visium 5 31,053
+Genome-wide spatial expression profiling in formalin-fixed tissues [villacampa2021genome] Brain Visium 15 31,053
+Human ileum, Visium [mirzazadeh2021human] Bowel Visium 4 33,538
+Human squamous cell carcinoma [abalo2021human] Skin Visium 4 33,538
+Prostate needle biopsies pre- and post-ADT: Count matrices, histological-, and Androgen receptor immunohistochemistry images [marklund2022prostate] Prostate Spatial Transcriptomics 24 26,437
+Spatially resolved clonal copy number alterations in benign and malignant tissue [erickson2022spatially] Prostate Visium 23 33,538
+Spatially resolved transcriptomic profiling of degraded [mirzazadeh2023spatially] Bowel Visium 35 17,943
 spatialRNAseq heart raw suppdata Heart Visium 4 54,848
@@
 Publication Organ Technology $\bm{n}$ Num. genes
-Transcriptome-scale spatial gene expression in the human dorsolateral prefrontal cortexMaynard et al. (2021) Brain Visium 12 33,538
-A spatially resolved atlas of the human lung characterizes a gland-associated immune nicheMadissoon et al. (2023) Lung Visium 20 17,922
+A spatially resolved atlas of the human lung characterizes a gland-associated immune niche[madissoon2023spatially] Lung Visium 20 17,922
+Molecular cartography uncovers evolutionary and microenvironmental dynamics in sporadic colorectal tumors[heiser_molecular_2023] Colon Visium 41 19,366
+Spatially resolved multiomics of human cardiac niches[kanemaru_spatially_2023] Heart Visium 41 33,538
+Transcriptome-scale spatial gene expression in the human dorsolateral prefrontal cortex[maynard2021transcriptome] Brain Visium 12 33,538
 Table A9: Internal datasets.
@@
 Prostate ST Internal Prostate Visium 4 17,943
-Tertiary lymphoid structures generate and propagate anti-tumor antibody-producing plasma cells in renal cell cancer Meylan et al. (2022) Lymph node Visium 24 17,943
+Tertiary lymphoid structures generate and propagate anti-tumor antibody-producing plasma cells in renal cell cancer [meylan2022tertiary] Lymph node Visium 24 17,943
 Table A10: Datasets gathered on Zenodo.
 Publication Organ Technology $\bm{n}$ Num. genes
-Charting the Heterogeneity of Colorectal Cancer Consensus Molecular Subtypes using Spatial Transcriptomics: datasets Valdeolivas et al. (2023) Bowel Visium 14 36,601
-Demo 10x Visium dataset for STQ zen (2024a) Skin Visium 1 68,886
-Spotiphy: generative modeling in single-cell spatial whole transcriptomics zen (2024b) Brain Visium 2 32,285
+Charting the Heterogeneity of Colorectal Cancer Consensus Molecular Subtypes using Spatial Transcriptomics: datasets [valdeolivas2023charting] Bowel Visium 14 36,601
+Demo 10x Visium dataset for STQ [zenodo2024demo] Skin Visium 1 68,886
+Nextflow Pipeline for Visium and H&E Data from Patient-Derived Xenograft Samples [Domanskyi2023.07.27.550727] Skin Visium 4 68886
+Spotiphy: generative modeling in single-cell spatial whole transcriptomics [zenodo2024spotiphy] Brain Visium 2 32,285
 C.2 HEST-Library
 The HEST-Library helps transform unstructured spatial transcriptomics and histology data into a unified format.
-An overview of the HEST-Library is provided in Suppl.
-Figure 3.
-An example of fiducial detection is presented in Suppl.
-Figure 4.
+An overview of the HEST-Library is provided in Figure 4.
+An example of fiducial detection is presented in Figure 5.
 Figure 4: Overview of HEST-Library functionalities.
-HEST was designed to transform legacy data scrapped in multiple public repositories, such as NCBI, into unified and actionable objects that can easily be integrated into multimodal computational pipelines.
-Figure 4: Fiducial detection and automatic alignment in Visium samples.
-Corner fiducials on 6.5$\times$6.5mm and 11mm×$\times$11mm Visium slides are automatically detected with a finetuned Yolov8 model, the spots coordinates are then derived if at least 3 of the 4 corner fiducials are detected.
+HEST was designed to transform legacy data scrapped in multiple public repositories, such as NCBI, into unified HEST objects that can easily be integrated into computational pipelines.
+Figure 5: Fiducial detection and automatic alignment in Visium.
+Corner fiducials on 6.5$\times$6.5mm and 11mm×$\times$11mm Visium slides are automatically detected with a finetuned Yolov8 model.
+The spot coordinates are then derived if at least 3 of the 4 corner fiducials are detected.
 This process enables automatically estimating the pixel resolution.
 C.3 HEST-Benchmark
-Gene selection, Random Forest, and Ridge regression models: We learn a regression model that maps the patch embeddings of each encoder to its corresponding gene expression profile.
-The Random Forest model is based on 70 trees (sklearn implementation).
-The Ridge regression uses a fixed $L2$ regularization coefficient $\lambda$ to $100/MC$, where $M$ is the embedding dimension and $C=50$ is the number of targets trained with the Regularized Least-Squares Routine solver (sklearn implementation).
-We train the regression models to predict a panel of 50 most variable genes tailored to each task.
-Specifically, for each task, we select the 50 most variable genes across all spots and samples, after excluding the genes that have non-zero counts in less than 10% of the spots.
+Gene selection, XGBoost Forest, and Ridge regression models: We learn a regression model that maps the patch embeddings of each encoder to its corresponding gene expression profile.
+The XGboost model uses 100 estimators, a 0.1 learning rate, a max depth of 3, 0.8 subsampling, gamma of 0.0, regression alpha of 0.0, and regression lambda of 1.0.
+Additional information can be found in the XGBoost API [^12] https://xgboost.readthedocs.io/en/stable/python/python_api.html.
+The Ridge regression uses a fixed $L2$ regularization coefficient $\lambda$ set to $100/MC$, where $M$ is the embedding dimension and $C=50$ is the number of targets trained with the Regularized Least-Squares Routine solver (sklearn implementation).
+Both regression models are trained to predict a panel constituted of the 50 most variable genes of each task.
+Specifically, for each task, we select the 50 most variable genes across all spots and samples after excluding the genes that have non-zero counts in less than 10% of the spots.
 Benchmark task description: We provide complementary information on each task introduced as part of the HEST-Benchmark.
@@
 Task ID Oncotree Organ Number of Patients Number of Samples Technology
-Task 1 Janesick et al. (2023) IDC Breast 4 4 Xenium
+Task 1 [janesick2023high] IDC Breast 4 4 Xenium
 Task 2 PRAD Prostate 2 23 Visium
@@
 Task 4 SKCM Skin 2 2 Xenium
-Task 5 Valdeolivas et al. (2023) COAD Colon 3 6 Visium
-Task 6 Valdeolivas et al. (2023) READ Rectum 2 4 Visium
-Task 7 Meylan et al. (2022) ccRCC Kidney 24 24 Visium
-Task 8 Giraud et al. (2022) HCC Liver 2 2 Visium
-Task 9 LUAD Lung 2 2 Xenium
-Task 10 Liu et al. (2022) IDC Axillary lymph nodes 4 4 Visium
-Task 1: Expression prediction in invasive ductal carcinoma (breast cancer, IDC).
-We used all publicly available Xenium samples available on 10x Genomics (“FFPE Human Breast using the Entire Sample Area”, 2 patients) and two samples published in  Janesick et al.
-(2023) (2 patients).
+Task 5 [valdeolivas2023charting] COAD Colon 2 4 Xenium
+Task 6 [valdeolivas2023charting] READ Rectum 2 4 Visium
+Task 7 [meylan2022tertiary] ccRCC Kidney 24 24 Visium
+Task 8 LUAD Lung 2 2 Xenium
+Task 9 [liu2022single] IDC Axillary lymph nodes 4 4 Visium
+Task 1: Prediction of expression in invasive ductal carcinoma (breast cancer, IDC).
+We used all publicly available Xenium samples available on 10x Genomics (“FFPE Human Breast using the Entire Sample Area”, 2 patients) and two samples published in  [janesick2023high] (TENX95, TENX99, NCBI783, NCBI785).
 All samples are FFPE sections imaged with the Xenium pipeline v1.
-Task 2: Expression prediction in prostate adenocarcinoma (prostate cancer, PRAD).
-We used all 23 Visium samples (fresh frozen sections) from 2 patients published in Erickson et al.
-(2022).
+Task 2: Prediction of expression in prostate adenocarcinoma (prostate cancer, PRAD).
+We used all 23 Visium samples (fresh frozen sections) from 2 patients published in [erickson2022spatially] (MEND139 to MEND162).
 Both patients were diagnosed with prostatic acinar adenocarcinoma with a (4+3) Gleason score (ISUP group 4).
-Task 3: Expression prediction in pancreatic adenocarcinoma (pancreatic cancer, PAAD).
+Task 3: Prediction of expression in pancreatic adenocarcinoma (pancreatic cancer, PAAD).
 We used 3 samples from 3 different patients from 10x Genomics (“FFPE Human Pancreas with Xenium Multimodal Cell Segmentation” and “Pancreatic Cancer with Xenium Human Multi-Tissue and Cancer Panel”).
-All samples are FFPE sections processed with Xenium pipeline v1.
+All samples are FFPE sections processed with Xenium pipeline v1 (TENX116, TENX126, TENX140).
 Task 4: Expression prediction in skin cutaneous melanoma (skin cancer, SKCM).
 We used 2 samples from 2 different patients from 10x Genomics website (“Human Skin Data with Xenium Human Multi-Tissue and Cancer Panel”).
-All samples are FFPE sections processed with Xenium pipeline v1.
-Task 5: Expression prediction in colon adenocarcinoma (colon cancer, COAD).
-We used 6 COAD samples from 3 different patients published in Valdeolivas et al.
-(2023).
+All samples are FFPE sections processed with Xenium pipeline v1 (TENX115, TENX117).
+Task 5: Prediction of expression in colon adenocarcinoma (colon cancer, COAD).
+We used 4 COAD samples from 2 different patients available on 10x Genomics (TENX111, TENX147, TENX148, TENX149).
 All samples are fresh frozen sections processed with Visium.
-Task 6: Expression prediction in rectal adenocarcinoma (rectum cancer, READ).
-We used 4 READ samples from different 2 patients published in Valdeolivas et al.
-(2023).
-All samples are fresh frozen sections processed with Visium.
-Task 7: Expression prediction in clear cell renal cell carcinoma (kidney cancer, ccRCC).
-We used all 24 ccRCC samples from 24 different patients published in Meylan et al.
-(2022).
-All samples are fresh frozen sections processed with Visium.
-Task 8: Expression prediction in hepatocellular carcinoma (liver cancer, HCC).
-We used 2 HCC liver samples from 2 different patients published in Giraud et al.
-(2022).
-All samples are fresh frozen sections processed with Visium.
-Task 9: Expression prediction in lung adenocarcinoma (lung cancer, LUAD).
+Task 6: Prediction of expression in rectal adenocarcinoma (rectum cancer, READ).
+We used 4 READ samples from 2 different patients published in [valdeolivas2023charting].
+All samples are fresh frozen sections processed with Visium (ZEN36, ZEN40, ZEN48, ZEN49).
+Task 7: Prediction of expression in clear cell renal cell carcinoma (kidney cancer, ccRCC).
+We used the 24 ccRCC samples of 24 different patients published in [meylan2022tertiary].
+All samples are fresh frozen sections processed with Visium (INT1 to INT24).
+Task 8: Prediction of expression in lung adenocarcinoma (lung cancer, LUAD).
 We used 2 LUAD samples from 2 different patients from 10x genomics (“Preview Data: FFPE Human Lung Cancer with Xenium Multimodal Cell Segmentation”).
-All samples are fresh frozen sections processed with Xenium pipeline v1.
-Task 10: Expression prediction in axillary lymph nodes in IDC patients.
-We used 4 axillary lymph node samples from 2 IDC patients published in Liu et al.
-(2022).
-All samples are fresh frozen sections processed with Visium.
+All samples are fresh frozen sections processed with Xenium pipeline v1 (TENX118, TENX141).
+Task 9: Prediction of expression in axillary lymph nodes in IDC patients.
+We used 4 axillary lymph node samples from 2 IDC patients published in [liu2022single].
+All samples are fresh frozen sections processed with Visium (NCBI681, NCBI682, NCBI683, NCBI684).
 Table A12: State-of-the-art foundation models for histology evaluated on HEST-Benchmark.
-ResNet-50: 23M parameters, ResNet-18: 11M, ResNet-152: 795M, ViT-B: Base (86M), L: Large (307M), G: Giant (1.13B)
-Name Number of Number of Magnification Model Training
-slides patches recipe
-ResNet50-IN Lu et al. (2021b) N/A 1M N/A ResNet-50 Supervised
-KimiaNet Riasatian et al. (2021) 7K 240K 20x ResNet-18 Supervised
-Ciga Ciga et al. (2022) UNK UNK 10x, 20x ResNet-50 Supervised
-CTransPath Wang et al. (2021) 32k 17M 10x UNK MoCov2
-Remedis Azizi et al. (2023) 10 10M 20x ResNet-152 iBOT
-Phikon Filiot et al. (2023) 6k 43.3M 20x ViT-B iBOT
-PLIP Huang et al. (2023) N/A 800k Many ViT-B CLIP
-UNI Chen et al. (2024) 100k 100M 20x ViT-L DINOv2
-CONCH Lu et al. (2024) N/A 1.17M Many ViT-B CoCa
-GigaPath Xu et al. (2024) 171k 1.3B 20x ViT-G CoCa
-We provide a brief description of each patch encoder assessed on the HEST-Benchmark.
-ResNet50 (IN) Lu et al.
-(2020): This model uses a ResNet50 backbone He et al.
-(2016) trained on ImageNet dataset Deng et al.
-(2009).
-Following prior work Lu et al.
-(2020), the patch embedding corresponds to the deep representation at the penultimate layer before classification.
-Ciga Ciga et al. (2022): This model uses a ResNet-50 backbone He et al. (2016) trained on public histology datasets such as TCGA and Cameleyon16 using SimCLR.
-CTransPath Wang et al.
-(2022): This model uses a “Tiny” Swin Transformer backbone Liu et al.
-(2021) with a window size of 14 (Swin-T/14, 28 million parameters) pretrained on TCGA and PAIP datasets (17 million images) using MoCoV3 Chen et al.
-(2021).
-Remedis Azizi et al.
-(2023): This model uses a ResNet-152$\times$2 (232 million parameters) initialized with the “Big Transfer”-medium protocol Kolesnikov et al.
-(2020) on ImageNet-22K and pretrained with SimCLR Chen et al.
-(2020) on TCGA.
-Phikon Filiot et al. (2023): This model uses a Vision Transformer-Base (ViT-B, 86 million parameters) Dosovitskiy et al. (2021) trained on TCGA data using iBOT Zhou et al. (2022).
-PLIP Huang et al.
-(2023): This model uses a ViT-B (86 million parameters) trained on 208 thousand pathology images paired with natural language descriptions extracted from Twitter and trained using CLIP Radford et al.
-(2021).
-UNI Chen et al.
-(2024): This model uses a ViT-Large (ViT-L, 307 million parameters) Dosovitskiy et al.
-(2021) trained on 100 million histology images (over 100,000 slides) from proprietary and public data using DINOv2 Oquab et al.
-(2023).
-CONCH Lu et al.
-(2024): This model uses a ViT-B (86 million parameters) trained on 1.17 million histology image–caption pairs extracted from online educational and research resources using CoCa Yu et al.
-(2022).
-GigaPath Xu et al. (2024): This model uses a ViT-Giant (1.13 billion parameters) trained on 1.3 billion image patches from 171,189 WSIs at 20$\times$ magnification using DINOv2.
+B: Base, L: Large, H: Huge, G: Giant.
+: number of patches during pretraining, ∗∗: number of patches during fine-tuning.
+Name Number of Number of Magnification Model Training Number of
+slides patches recipe parameters
+ResNet50 (IN) [lu2021data] N/A 1.2M N/A ResNet-50 Supervised 23M
+CTransPath [wang2021transpath] 32k 17M 10$\times$ Swin-T MoCov3 28M
+Remedis [azizi2023robust] 10k 10M 20$\times$ ResNet-152 iBOT 232M
+Phikon [filiot2023scaling] 6k 43.3M 20$\times$ ViT-B iBOT 86M
+UNI [chen2024towards] 100k 100M 20$\times$ ViT-L DINOv2 307M
+CONCH [lu2024towards] N/A 16M + 1.17M∗∗ Many ViT-B CoCa 86M
+GigaPath [xu2024wholeslide] 171k 1.3B 20$\times$ ViT-g DINOv2 1.13B
+Virchow [vorontsov2024foundation] 1.5M 2B 20$\times$ ViT-H DINOv2 632M
+Virchow 2 [Zimmermann2024Virchow2SS] 3.1M 1.9B Many ViT-H DINOv2 632M
+H-Optimus-0 500K 273M 20x ViT-g DINOv2 1.13B
+UNIv1.5 350K 432M 20x ViT-g DINOv2 1.13B
+We provide a brief description of each patch encoder assessed with the HEST-Benchmark.
+ResNet50 (IN) [lu2020data]: This model uses a ResNet50 backbone [he2016deep] trained on ImageNet [deng2009imagenet] (1.2 million natural images).
+Following prior work [lu2020data], the patch embeddings are extracted by taking the representation at the penultimate layer before final classification.
+CTransPath [wang2022transformer]: This model uses a “Tiny” Swin Transformer backbone [liu2021swin] with a window size of 14 (Swin-T/14, 28 million parameters) pretrained on TCGA and PAIP datasets (17 million images) using MoCoV3 [chen2021an].
+Remedis [azizi2023robust]: This model uses a ResNet-152$\times$2 (232 million parameters) initialized with the “Big Transfer”-medium protocol [kolesnikov2020big] on ImageNet-22K and pretrained with SimCLR [chen2020simple] on TCGA.
+Phikon [filiot2023scaling]: This model uses a Vision Transformer-Base (ViT-B, 86 million parameters) [dosovitskiy2021image] trained on TCGA data using iBOT [zhou2021ibot].
+UNI [chen2024towards]: This model uses a ViT-Large (ViT-L, 307 million parameters) [dosovitskiy2021image] trained on 100 million histology images (over 100,000 slides) from proprietary and public data using DINOv2 [oquab2023dinov2].
+CONCH [lu2024towards]: This model uses a ViT-B (86 million parameters) trained on a smaller version of UNI using iBOT, and then fine-tuned on 1.17 million histology image–caption pairs extracted from online educational and research resources using CoCa [yu2022coca].
+GigaPath [xu2024wholeslide]: This model uses a ViT-giant (1.13 billion parameters) trained on 1.3 billion image patches from 171,189 WSIs at 20$\times$ magnification using DINOv2.
+Virchow [vorontsov2024foundation]: This model uses a ViT-Huge (632M parameters) trained on 2 billion image patches and 1.5M WSIs at 20$\times$ magnification using DINOv2.
+Virchow 2 [Zimmermann2024Virchow2SS]: This model uses a ViT-Huge (632M parameters) trained on 1.9B patches and 3.1M WSIs using DINOv2
+H-Optimus-0: This model uses a ViT-giant (1.13B parameters) trained 273 million image patches from 500,000 WSIs at 20$\times$ magnification using DINOv2.
+UNIv1.5: This model uses a ViT-giant (1.13B parameters) trained on 432 million image patches from 350,000 WSIs using DINOv2.
 Appendix D HEST for multimodal representation learning
+We provide additional information regarding CONCH fine-tuning using multimodal alignment.
 CONCH-FT model, a ViT-Base model initialized with CONCH weights, was fine-tuned for 50 epochs using a cosine learning rate scheduler, with a base learning rate of $10^{-4}$ for the image encoder and $10^{-3}$ for the expression encoder.
@@
 A combination of random horizontal/vertical flip and color jittering was employed for image patch augmentation.
-The rank of the embedding space, also referred to as smooth rank measure Garrido et al.
-(2023), measures the quality of the embeddings produced from encoders trained in unsupervised or self-supervised manners.
+The rank of the embedding space (also referred to as smooth rank measure [garrido2023rankme]) measures the quality of the embeddings produced from encoders trained in unsupervised or self-supervised manners.
 Given the patch embedding matrix $H\in\mathcal{R}^{N\times d}$ and $d<N$, where $N$ is the number of patches and $d$ is the feature dimension, we compute the rank as the entropy of the $d$ L1-normalized singular values of $H$.
 Appendix E HEST for discovery
-Cells were segmented and classified using CellViT Hörst et al.
-(2024).
+Cells were segmented and classified using CellViT [horst2024cellvit].
 To find the gene expression profile of each neoplastic cell, we matched each cell to its corresponding cell index in Xenium by assigning the index for which the distance between the cell centroids was the smallest.
@@
 After this filtering step, an average of 91% of the cells per sample were kept while 9% of the cells were discarded.
-The discarded cells likely correspond to over-segmentation errors by CellViT.
 Figure 6: HEST for biomarker discovery: Analysis of an invasive ductal carcinoma Xenium sample.
@@
   • Why was the dataset created?
-HEST-1k was designed with three key applications: (1) multimodal representation learning of histology and transcriptomics, (2) biomarker discovery, and (3) benchmarking of histopathology patch encoder models.
-Despite many publicly available resources, no existing unified and user-friendly formatting was available to bring ST in the world of deep learning.
+HEST-1k was designed with three key applications: (1) multimodal representation learning of histology and transcriptomics, (2) biomarker exploration and characterization, and (3) benchmarking foundation models for pathology.
+Despite many publicly available resources, no existing unified and user-friendly formatting was available to bring ST into the world of deep learning.
   • What (other) tasks could the dataset be used for?
 Are there obvious tasks for which it should not be used?
-Users are welcome to propose new creative ways to use the dataset.
-Users are not allowed to try to retrieve patient information from the existing data.
-A dedicated section is allocated to discuss ethical considerations and intended usage.
+Users are welcome to introduce new, creative ways to use the dataset.
+However, users are not allowed to try to retrieve patient information from the existing data.
+A dedicated section is provided to discuss ethical considerations and intended usage.
   • Has the dataset been used for any tasks already?
 If so, where are the results so others can compare (e.g., links to published papers)?
-All samples that were made public as part of a publication (peer-review or not) are reported in the metadata attached with HEST-1k.
-  • Who funded the creation of the dataset? The work was funded by an R35 grant.
+The metadata attached to HEST-1k reports all samples that were made public as part of a publication (peer-reviewed or not).
+  • Who funded the creation of the dataset?
+HEST is supported by the Brigham and Women’s Hospital (BWH) President’s Fund, Mass General Hospital (MGH) Pathology, and the National Institute of Health (NIH) National Institute of General Medical Sciences (NIGMS) through R35GM138216.
 F.2 Dataset composition
-  • What are the instances? The modalities used in this study are histopathology whole-slide images, gene expression data, and derivatives of these two modalities, such as nuclear segmentation maps.
+  • What are the instances?
+The modalities used in this study are histopathology whole-slide images, gene expression data, and derivatives of these two modalities, such as nuclear segmentation and classification maps.
   • Are relationships between instances made explicit in the data Each whole-slide image maps to a unique gene expression profile in an unequivocal way.
   • What data does each instance consist of?
-Imaging data consists of Generic TIFF objects stored in pyramidal format, and gene expression data consists of scanpy objects.
-Derivatives are stored in JSON and Hierarchical Data Format (HDF) files.
+Imaging data consists of Generic TIFF objects stored in a pyramidal format, and gene expression data consists of scanpy objects.
+Derivatives are stored in JSON files, parquet files, and Hierarchical Data Format (HDF) files.
   • Is there a label/target associated with instances?
@@
 Are there licenses, fees or rights associated with any of the data?
-We provide all data as part of HEST-1k release.
+We provide all data as part of the HEST-1k release.
 In addition, a link to the original data is provided in the metadata.
 Each sample is associated with a license as provided by the original publication, where we ensured that the reported license allowed for distributing and creating derivatives of the data.
-Should the author of the original publication would like to remove their samples from HEST-1k, we will update the data promptly.
   • Are there recommended data splits or evaluation measures?
@@
 Have a summary of those results and, if available, provide the link to a paper with more information here.
-At the time of submission, all experiments run with HEST-1k are part of this study.
-The reader can refer to the main text for a thorough description of all experiments (see HEST-Benchmark, HEST for discovery, HEST for multimodal representation learning).
+All experiments run with HEST-1k are described in this study.
+The reader can refer to the main text for a thorough description of all experiments (see HEST-Benchmark, HEST for biomarker exploration, HEST for multimodal representation learning).
 F.3 Data collection process
@@
 All authors of the present study were involved in the data collection, inspection, and curation.
-The reader can refer to the original publication to understand how the data were orginally acquired.
-  • Over what time-frame was the data collected?
-Does the collection time-frame match the creation time-frame?
+The reader can refer to the original publication to understand how the data were originally acquired.
+  • Over what time frame was the data collected?
+Does the collection time frame match the creation time frame?
 The original data comprise publications from 2016 to 2024.
@@
 (this does not include intentionally dropped instances; it might include, e.g., redacted text, and withheld documents) Is this data missing because it was unavailable?
-Original publications include some missing information, such as the alignment file between the slide and the expression profile.
-We developed computational tools to minimize missing information to reach near-complete metadata.
+Original publications may include some missing information, such as the alignment file between the slide and the expression profile.
+We developed computational tools to minimize missing information and reach near-complete metadata.
   • Are there any known errors, sources of noise, or redundancies in the data?
@@
 Gene expression data are inherently noisy.
-Users can decide to apply post-hoc normalization methods to reduce noise.
+Users can decide to apply post-hoc normalization methods to reduce noise, e.g., stain normalization on the imaging side or batch effect mitigation on the transcriptomics side.
 F.4 Data preprocessing
   • What preprocessing/cleaning was done?
-All whole-slide images were converted into a pyramidal TIFF objects with re-estimated pixel resolution.
+All whole-slide images were converted into pyramidal TIFF objects with re-estimated pixel resolution.
 All alignment files have been manually inspected and included if missing.
@@
   • Was the “raw” data saved in addition to the preprocessed/cleaned data?
-(e.g., to support unanticipated future uses) Raw data are downloaded, but not publicly shared.
+(e.g., to support unanticipated future uses) Raw data are downloaded but not publicly shared.
 In the case of public samples, users can re-download them using the metadata provided as part of the dataset release.
@@
 F.5 Dataset distribution
-  • How is the dataset distributed? HEST-1k is distributed using internal storage. We have several internal copies to ensure data preservation over the years.
-  • When will the dataset be released/first distributed? The dataset will be made public with the camera-ready version of this publication. This study serves as the canonical reference.
+  • How is the dataset distributed? HEST-1k is distributed using HuggingFace Datasets.
+  • When will the dataset be released/first distributed? The dataset is public and can be accessed through the HuggingFace Datasets interface.
   • What license (if any) is it distributed under?
@@
 The dataset is distributed under the Attribution-NonCommercial-ShareAlike 4.0 International license (CC BY-NC-SA 4.0 Deed).
-  • Are there any fees or access/export restrictions? No access/export restrictions unless they violates the terms of the above-mentioned license (CC BY-NC-SA 4.0 Deed).
+  • Are there any fees or access/export restrictions? No access/export restrictions unless they violate the terms of the above-mentioned license (CC BY-NC-SA 4.0 Deed).
   • Who is supporting/hosting/maintaining the dataset? How does one contact the owner/curator/manager of the dataset? The dataset is maintained by the authors of the publication.
@@
 Dataset versioning will be put in place.
-A website will clearly state what the latest version of the dataset contains and keep track of the changes in each version.
-  • If the dataset becomes obsolete how will this be communicated? The corresponding website will state if the dataset becomes obsolete.
+  • If the dataset becomes obsolete how will this be communicated? The GitHub README will be updated.
   • Is there a repository to link to any/all papers/systems that use this dataset?
@@
 What is the process for communicating/distributing these contributions to users?
-Users are welcome to contact us if they would like to provide additional data that meet the standards of the data contained in HEST-1k.
+Users are welcome to contact us if they would like to provide additional data that meets our standards.
 We do not have a dedicated system to communicate these contributions.
@@
   • If it relates to people, could this dataset expose people to harm or legal action?
-(e.g., financial social or otherwise) What was done to mitigate or reduce the potential for harm?
+(e.g., financial, social or otherwise) What was done to mitigate or reduce the potential for harm?
 No, patients cannot be linked to the corresponding histology and gene expression profile.
@@
 The authors of this paper bear all responsibility in case of violation of rights associated with HEST-1k, HEST-Library, and HEST-Benchmark.
-Table A13: HEST-Benchmark results.
-Ridge regression evaluation on public state-of-the-art patch encoders.
-Model performance measured with Pearson correlation.
-We report the mean $\pm$ standard deviation computed over all folds (or patients).
-Best is bold, second best is underlined.
-ResNet50 KimiaNet Ciga CTransPath Remedis Phikon PLIP UNI CONCH Gigapath
-IDC 0.446 0.455 0.425 0.499 0.496 0.526 0.450 $\bm{0.544}$ 0.541 $\underline{0.542}$
-±$\pm 0.04$$\pm 0.03$$\pm 0.04$$\pm 0.06$$\pm 0.07$$\pm 0.08$$\pm 0.04$$\pm 0.07$$\pm 0.07$$\pm 0.09$
-PRAD 0.276 0.323 0.258 $0.290$ 0.264 0.249 0.262 0.257 $\bm{0.353}$ 0.301
-±$\pm 0.06$$\pm 0.04$$\pm 0.06$$\pm 0.07$$\pm 0.08$$\pm 0.13$$\pm 0.02$$\pm 0.08$$\pm 0.01$$\pm 0.03$
-PAAD 0.343 0.374 0.364 $\underline{0.383}$ 0.288 0.360 0.358 0.377 $\bm{0.424}$ 0.323
-±$\pm 0.07$$\pm 0.08$$\pm 0.07$$\pm 0.07$$\pm 0.05$$\pm 0.07$$\pm 0.06$$\pm 0.06$$\pm 0.04$$\pm 0.00$
-SKCM $\underline{0.413}$ 0.357 0.381 0.404 0.412 0.369 0.337 0.343 $\bm{0.546}$ 0.190
-±$\pm 0.08$$\pm 0.05$$\pm 0.06$$\pm 0.07$$\pm 0.04$$\pm 0.11$$\pm 0.08$$\pm 0.06$$\pm 0.05$$\pm 0.05$
-COAD $\underline{0.116}$ 0.103 0.093 0.113 0.075 0.114 0.108 0.110 $\bm{0.135}$ 0.085
-±$\pm 0.05$$\pm 0.04$$\pm 0.02$$\pm 0.05$$\pm 0.04$$\pm 0.05$$\pm 0.05$$\pm 0.06$$\pm 0.06$$\pm 0.03$
-READ 0.067 0.058 0.081 0.090 0.077 0.114 0.084 0.124 $\bm{0.154}$ $\underline{0.125}$
-±$\pm 0.06$ 0.05 ±$\pm 0.05$$\pm 0.08$$\pm 0.07$$\pm 0.07$$\pm 0.06$$\pm 0.04$$\pm 0.04$$\pm 0.05$
-ccRCC 0.210 0.190 0.180 0.214 0.219 $\bm{0.253}$ 0.174 $\underline{0.240}$ 0.221 0.226
-±$\pm 0.05$$\pm 0.06$$\pm 0.04$$\pm 0.04$$\pm 0.04$$\pm 0.05$$\pm 0.04$$\pm 0.05$$\pm 0.03$$\pm 0.05$
-HCC $\underline{0.039}$ 0.031 0.027 0.024 0.016 0.038 0.032 0.017 $\bm{0.046}$ 0.008
-±$\pm 0.02$$\pm 0.01$$\pm 0.02$$\pm 0.02$$\pm 8.71$$\pm 0.01$$\pm 0.03$$\pm 0.00$$\pm 0.01$$\pm 0.00$
-LUAD 0.400 0.419 0.417 0.403 0.311 $\underline{0.422}$ 0.406 0.371 $\bm{0.493}$ 0.297
-±$\pm 0.06$$\pm 0.07$$\pm 0.04$$\pm 0.07$$\pm 0.04$$\pm 0.06$$\pm 0.01$$\pm 0.11$$\pm 0.03$$\pm 0.08$
-LYM-IDC 0.203 0.227 0.218 0.209 0.167 0.215 0.215 0.224 $\bm{0.247}$ 0.191
-±$\pm 0.05$$\pm 0.04$$\pm 0.04$$\pm 0.04$$\pm 0.04$$\pm 0.04$$\pm 0.05$$\pm 0.03$$\pm 0.04$$\pm 0.02$
-Average 0.251 0.253 0.244 0.263 0.232 0.266 0.243 0.260 $\bm{0.316}$ 0.229
+Table A13: HEST-Benchmark evaluated using Ridge regression. Model performance measured with Pearson correlation. Best is bold, second best is underlined.
+IDC PRAD PAAD SKCM COAD READ ccRCC LUAD LYMPH IDC Average
+REMEDIS 0.4936 0.2632 0.2881 0.4117 0.151 0.0776 0.2201 0.3114 0.1694 0.2651
+$\pm$ 0.0725 $\pm$ 0.0821 $\pm$ 0.0544 $\pm$ 0.0384 $\pm$ 0.0147 $\pm$ 0.0684 $\pm$ 0.0418 $\pm$ 0.0432 $\pm$ 0.0365
+GigaPath 0.532 0.3035 0.3172 0.2231 0.163 0.1236 0.2172 0.3144 0.1925 0.2652
+$\pm$ 0.0812 $\pm$ 0.0279 $\pm$ 0.0165 $\pm$ 0.0071 $\pm$ 0.041 $\pm$ 0.0379 $\pm$ 0.0479 $\pm$ 0.0871 $\pm$ 0.0304
+UNIv1.5 0.5657 0.3065 0.3004 0.258 0.1982 0.1077 0.2023 0.3084 0.1998 0.2719
+$\pm$ 0.0866 $\pm$ 0.02 $\pm$ 0.0199 $\pm$ 0.0514 $\pm$ 0.0262 $\pm$ 0.0222 $\pm$ 0.0595 $\pm$ 0.0747 $\pm$ 0.0129
+H-Optimus-0 0.5789 0.2561 0.3367 0.2778 0.1605 0.1228 0.2342 0.3143 0.1976 0.2754
+$\pm$ 0.0899 $\pm$ 0.0003 $\pm$ 0.0428 $\pm$ 0.0048 $\pm$ 0.0522 $\pm$ 0.0309 $\pm$ 0.0373 $\pm$ 0.083 $\pm$ 0.0253
+Virchow2 0.5666 0.2972 0.2718 0.303 0.1814 0.1208 0.2257 0.3017 0.2172 0.2762
+$\pm$ 0.0848 $\pm$ 0.037 $\pm$ 0.0387 $\pm$ 0.0184 $\pm$ 0.0326 $\pm$ 0.0526 $\pm$ 0.0433 $\pm$ 0.1199 $\pm$ 0.019
+Virchow 0.5583 0.2744 0.3361 0.3389 0.1825 0.0955 0.2375 0.2897 0.2081 0.2801
+$\pm$ 0.0876 $\pm$ 0.0019 $\pm$ 0.037 $\pm$ 0.0063 $\pm$ 0.0369 $\pm$ 0.0527 $\pm$ 0.0371 $\pm$ 0.0785 $\pm$ 0.0234
+ResNet50 0.4453 0.2753 0.3432 0.413 0.2009 0.0669 0.2103 0.4001 0.203 0.2842
+$\pm$ 0.0377 $\pm$ 0.0622 $\pm$ 0.0654 $\pm$ 0.0814 $\pm$ 0.061 $\pm$ 0.0646 $\pm$ 0.0548 $\pm$ 0.0637 $\pm$ 0.0536
+CTransPath 0.4996 0.2895 0.3826 0.4038 0.1751 0.0909 0.2139 0.4026 0.2089 0.2963
+$\pm$ 0.0594 $\pm$ 0.0724 $\pm$ 0.066 $\pm$ 0.065 $\pm$ 0.0423 $\pm$ 0.0808 $\pm$ 0.0438 $\pm$ 0.07 $\pm$ 0.0367
+Phikon 0.5259 0.2493 0.3594 0.3684 0.1697 0.1136 0.253 0.4224 0.2151 0.2974
+$\pm$ 0.0791 $\pm$ 0.1264 $\pm$ 0.0707 $\pm$ 0.1061 $\pm$ 0.0562 $\pm$ 0.0749 $\pm$ 0.0483 $\pm$ 0.0579 $\pm$ 0.0416
+UNI 0.563 0.257 0.3768 0.3433 0.1839 0.1239 0.2395 0.3714 0.2236 0.2981
+$\pm$ 0.0771 $\pm$ 0.0819 $\pm$ 0.0555 $\pm$ 0.0556 $\pm$ 0.0509 $\pm$ 0.0434 $\pm$ 0.0557 $\pm$ 0.1098 $\pm$ 0.0289
+CONCH 0.528 0.3604 0.4224 0.5079 0.2467 0.1443 0.2356 0.4957 0.2462 0.3541
+$\pm$ 0.0794 $\pm$ 0.0135 $\pm$ 0.0773 $\pm$ 0.0281 $\pm$ 0.0045 $\pm$ 0.0455 $\pm$ 0.0387 $\pm$ 0.0203 $\pm$ 0.0349
+Table A14: HEST-Benchmark evaluated using XGBoost regression. Model performance measured with Pearson correlation. Best is bold, second best is underlined.
+IDC PRAD PAAD SKCM COAD READ ccRCC LUAD LYMPH IDC Average
+ResNet50 (IN) 0.4646 0.3433 0.4017 0.4707 0.2892 0.0586 0.181 0.4967 0.2284 0.326
+$\pm$ 0.0353 $\pm$ 0.0168 $\pm$ 0.0648 $\pm$ 0.0834 $\pm$ 0.0115 $\pm$ 0.069 $\pm$ 0.0502 $\pm$ 0.01 $\pm$ 0.0511
+CTransPath 0.4738 0.3514 0.4257 0.5304 0.2921 0.0996 0.2026 0.5225 0.234 0.348
+$\pm$ 0.0394 $\pm$ 0.0032 $\pm$ 0.0701 $\pm$ 0.073 $\pm$ 0.0018 $\pm$ 0.0766 $\pm$ 0.0387 $\pm$ 0.0063 $\pm$ 0.0613
+Phikon 0.4704 0.3943 0.3988 0.5323 0.277 0.1451 0.213 0.542 0.2443 0.3575
+$\pm$ 0.0672 $\pm$ 0.0123 $\pm$ 0.0598 $\pm$ 0.0607 $\pm$ 0.0098 $\pm$ 0.0851 $\pm$ 0.0362 $\pm$ 0.017 $\pm$ 0.0632
+GigaPath 0.5222 0.3749 0.4415 0.5297 0.2876 0.1609 0.2207 0.5506 0.2464 0.3705
+$\pm$ 0.0641 $\pm$ 0.0103 $\pm$ 0.058 $\pm$ 0.0376 $\pm$ 0.0039 $\pm$ 0.0777 $\pm$ 0.0402 $\pm$ 0.0108 $\pm$ 0.0526
+CONCH 0.5175 0.3784 0.4428 0.5766 0.3215 0.1431 0.1738 0.5581 0.2554 0.3742
+$\pm$ 0.0602 $\pm$ 0.0124 $\pm$ 0.0657 $\pm$ 0.0519 $\pm$ 0.0062 $\pm$ 0.0665 $\pm$ 0.0544 $\pm$ 0.0081 $\pm$ 0.0605
+REMEDIS 0.5116 0.3526 0.4621 0.5885 0.319 0.1129 0.2303 0.562 0.2521 0.3768
+$\pm$ 0.0594 $\pm$ 0.0073 $\pm$ 0.0555 $\pm$ 0.0253 $\pm$ 0.0101 $\pm$ 0.0846 $\pm$ 0.0393 $\pm$ 0.0057 $\pm$ 0.0601
+Virchow2 0.5378 0.3772 0.4237 0.5565 0.281 0.1779 0.2428 0.5641 0.2582 0.3799
+$\pm$ 0.0685 $\pm$ 0.007 $\pm$ 0.0525 $\pm$ 0.0152 $\pm$ 0.0162 $\pm$ 0.077 $\pm$ 0.0361 $\pm$ 0.0069 $\pm$ 0.0504
+UNI 0.538 0.3513 0.451 0.6089 0.2921 0.1679 0.235 0.5357 0.2456 0.3806
+$\pm$ 0.0603 $\pm$ 0.0162 $\pm$ 0.0587 $\pm$ 0.0294 $\pm$ 0.0191 $\pm$ 0.0641 $\pm$ 0.0381 $\pm$ 0.0057 $\pm$ 0.05
+Virchow 0.5309 0.3447 0.4448 0.6089 0.3275 0.1419 0.2307 0.5643 0.2617 0.3839
+$\pm$ 0.0764 $\pm$ 0.0117 $\pm$ 0.0501 $\pm$ 0.0165 $\pm$ 0.0254 $\pm$ 0.0669 $\pm$ 0.0336 $\pm$ 0.0091 $\pm$ 0.0537
+UNIv1.5 0.555 0.3654 0.434 0.6025 0.336 0.1742 0.2166 0.5634 0.2515 0.3887
+$\pm$ 0.0763 $\pm$ 0.0098 $\pm$ 0.0568 $\pm$ 0.0385 $\pm$ 0.0179 $\pm$ 0.0568 $\pm$ 0.0337 $\pm$ 0.0054 $\pm$ 0.0434
+H-Optimus-0 0.5564 0.3829 0.4445 0.6502 0.2922 0.1731 0.2402 0.5654 0.2555 0.3956
+$\pm$ 0.0777 $\pm$ 0.0049 $\pm$ 0.0563 $\pm$ 0.0326 $\pm$ 0.0063 $\pm$ 0.0777 $\pm$ 0.0348 $\pm$ 0.0084 $\pm$ 0.0522
+[…]